### 03 Mar 2026 v1

# Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model

Abstract.
In this paper, we study an intertemporal utility maximization problem in which an investor chooses consumption and portfolio strategies in the presence of a stochastic factor and a no-borrowing constraint. In the spirit of the Kim–Omberg model, the stochastic factor represents the excess return of the risky asset and follows an Ornstein–Uhlenbeck process, capturing the mean reversion of expected excess returns—a feature well supported by empirical evidence in financial markets. The investor seeks to maximize expected utility from consumption, subject to the constraint that wealth remains nonnegative at all times. To address the dynamic no-borrowing constraint, we use Lagrange duality to transform the primal problem into a singular control problem in the dual space. We then characterize the solution to the dual singular control problem via an auxiliary two-dimensional optimal stopping problem featuring stochastic volatility, and subsequently retrieve the primal value function as well as the optimal portfolio and consumption plans. Finally, a numerical study is conducted to derive economic and financial implications.
a Bielefeld University, Center for Mathematical Economics (IMW), Bielefeld (Germany).
Keywords: optimal consumption and portfolio choice, Kim-Omberg model, no-borrowing constraint, singular stochastic control, optimal stopping, stochastic volatility.
AMS subject classification: 91G15, 91G30, 49N15, 90C39, 60G40, 93E20
1. Introduction
We study an infinite-horizon optimal consumption and investment problem in which an agent chooses how much to consume and how to allocate wealth in a financial market. The agent receives a constant stream of labor income and faces a no-borrowing constraint, meaning that wealth must remain nonnegative at all times. This rules out borrowing either in financial markets or against future labor income, so that all decisions must be financed by current resources.
Investment opportunities vary over time because the expected excess return of the risky asset is driven by a stochastic factor $(\beta_{t})_{t}$. Empirical evidence (see, e.g., [14] and [37]) suggests that expected excess returns are predictable and mean-reverting. Motivated by this finding, we model $(\beta_{t})_{t}$ as a mean-reverting Ornstein–Uhlenbeck process, following the so-called Kim–Omberg model introduced in [27]. Furthermore, we allow the Brownian motion driving this stochastic factor to be correlated with the Brownian motion driving the risky asset’s returns.
We derive optimal consumption and portfolio policies, as well as regularity results for the value function. As discussed in more detail below, this is achieved by means of a Lagrange duality approach, which allows us to connect the (primal) dynamic optimization problem with a no-borrowing constraint to a singular stochastic control problem. The latter is then analyzed through an auxiliary genuinely two-dimensional optimal stopping problem, in which one of the state variables exhibits stochastic volatility. In particular, we establish continuous differentiability of the optimal stopping problem’s value function and characterize the optimal stopping time in terms of an excess-return-dependent free boundary. These properties are then instrumental in obtaining the complete solution to the original optimal consumption and portfolio choice problem.
Methodology and Results. The no-borrowing constraint requires that the agent’s wealth $(X_{t})_{t}$ remain nonnegative at all times, almost surely. Such a restriction affects the set of admissible consumption and portfolio choices and, when the agent’s optimization problem is addressed via dynamic programming, it complements the Hamilton–Jacobi–Bellman equation with an appropriate boundary condition.
Inspired by [12], [22], and the more recent [24], we instead adopt a duality-based approach to handle the no-borrowing constraint. Instead of enforcing $X_{t}\geq 0$ dynamically at every point in time, we reformulate it as a single static budget constraint. This transformation is achieved by introducing a non-increasing, càdlàg process $(D_{t})_{t}$ that acts as an endogenous dynamic Lagrange multiplier designed to ensure $X_{t}\geq 0$ at all times $t\geq 0$ almost surely. This dual formulation allows us to express the problem in terms of an auxiliary process $(Z_{t})_{t}$ rather than the wealth process itself. Consequently, the dual problem is recast as a two-dimensional singular control problem in the state variables $(Z_{t},\beta_{t})_{t}$, where $(Z_{t})_{t}=(Z_{t}^{D})_{t}$ is controlled through the monotone control $(D_{t})_{t}$. Crucially, because the stochastic factor $(\beta_{t})_{t}$ directly drives the diffusion of the dual state, the system inherently features stochastic volatility. It is worth noticing that directly approaching the dynamic programming principle (Hamilton-Jacobi-Bellmann, HJB) equation related to the primal problem is particularly challenging in our setting, because of a non-geometric wealth dynamics due to the presence of labor income. As a matter of fact, in many classical settings, the wealth process is geometric, and the positivity of wealth is automatically guaranteed. In presence of power utility functions, this geometric structure also allows the value function to be scaled with wealth, reducing the HJB equation to a one-dimensional nonlinear ODE (see, e.g., [32] and [31] for earlier works or the recent [18] and [19]). In our setting, however, the usual scaling arguments break down, the no-borrowing constraint must be enforced explicitly, and a duality approach is needed (see also Remark 2.3 below).
A key methodological step in our analysis is relating the dual singular stochastic control problem to an equivalent two-dimensional optimal stopping problem–which inherits state variables featuring stochastic volatility–using probabilistic arguments. Through this approach, our main results provide a complete characterization of the investor’s optimal behavior under stochastic investment opportunities, labor income, and a binding no-borrowing constraint.
First, we establish that the auxiliary two-dimensional optimal stopping problem features a lower-semicontinuous free boundary $\beta\mapsto z^{*}(\beta)$ that strictly separates the continuation and stopping regions, thereby defining the optimal stopping time. We then prove that the stopping value function $v$ is locally Lipschitz continuous across the entire state space and infinitely differentiable within both the continuation region and the interior of the stopping region. This latter regularity result is obtained by showing that – despite the degeneracy of the second-order differential operator $\mathcal{L}$ associated with the optimal stopping problem’s state process and the stochastic volatility of one state variable – Hörmander’s condition holds. Consequently, $\mathcal{L}$ is hypoelliptic, and $v$ is a classical solution to the associated partial differential equation within the continuation region. Hörmander’s condition also has the important implication that the optimal stopping problem’s state process admits a smooth transition density. This, in turn, guarantees, via an application of a result in [23], that $v$ is continuously differentiable across the entire state space.
By connecting the optimal stopping problem back to the dual singular control problem, we show that the dual value function is precisely the integral of the stopping value function with respect to the first component of the state process. This critical relationship allows us to uniquely characterize the optimal singular control $(D^{*}_{t})_{t}$, which is the minimal process (à la Skorokhod) that keeps the dual state process $(Z_{t},\beta_{t})_{t}$ within the region $\{(z,\beta):z<z^{*}(\beta)\}$. Finally, we prove that strong duality holds, demonstrating that the primal value function can be recovered from the dual value function, and vice versa. Using these results, we are then able to retrieve the optimal consumption plan $(c^{*}_{t})_{t}$, the optimal portfolio strategy $(\pi^{*}_{t})_{t}$, and the optimal wealth process $(X^{*}_{t})_{t}$, which, as required, adheres to the dynamic no-borrowing constraint.
Economically, these mathematical results offer a clear interpretation of the system’s dynamics. The process $(D^{*}_{t})_{t}$ represents a shadow price, reflecting the marginal cost of relaxing the wealth constraint. When consumption increases significantly, it exerts pressure on the borrowing constraint as wealth depletes more quickly. To maintain feasibility, $(D^{*}_{t})_{t}$ adjusts downward, reflecting the reduced capacity to fund future consumption or investments. Concurrently, strong duality will imply that $(Z^{D^{*}}_{t})_{t}$ represents the marginal value of an additional unit of wealth. As wealth becomes scarce and approaches zero, its marginal value naturally rises, causing $(Z^{D^{*}}_{t})_{t}$ to increase. Specifically, as the wealth process approaches zero, the marginal value process $(Z^{D^{*}}_{t})_{t}$ freely evolves upward until it hits the state-dependent free boundary $z^{*}$. Exactly at the point where $(Z^{D^{*}}_{t})_{t}$ touches this boundary, the singular control $(D^{*}_{t})_{t}$ activates and decreases, pushing $(Z^{D^{*}}_{t})_{t}$ downward. By the established strong duality, this reflection of the dual state corresponds precisely to the wealth process being reflected upward, preventing bankruptcy and ensuring that the no-borrowing constraint is strictly satisfied.
Related Literature. The study of optimal consumption and investment problems with labor income and no-borrowing constraints dates back to 1993, when [22] developed a duality approach to study an individual’s optimal consumption and portfolio policy when borrowing against future labor income is limited. Similar to [22], [12] addresses an optimal consumption problem for an agent facing stochastic labor income and a strict no-borrowing constraint. They also utilize a duality approach to transform the constrained problem into a solvable, unconstrained dual problem. The resulting optimal strategy is characterized as a singular control problem, where the agent’s actions near the wealth boundary ($X_{t}=0$) are described by a local time component. Crucially, the constrained optimal wealth is shown to be equivalent to the unconstrained wealth minus the value of an American put option, establishing a direct link between the portfolio problem and optimal stopping theory. However, while these earlier works include labor income and borrowing constraints, they do not feature a stochastic factor. Recent studies such as [24] study optimal consumption, investment, and early retirement decisions for an agent under a finite-time horizon and a strict no-borrowing constraint against future labor income. Using the dual-martingale method, the problem is uniquely formulated as a two-person zero-sum game between a singular controller (managing the borrowing constraint) and a discretionary stopper (choosing the retirement time). The solution is governed by a min-max parabolic variational inequality that results in two time-varying free boundaries: one for optimal retirement and one for the active wealth binding constraint. While [24] considers a related singular control structure, it again operates without a stochastic factor.
Optimal consumption and investment problems in the presence of a stochastic factor have also been widely studied. For example, [33] consider an optimal consumption problem involving stochastic interest rates, while [20] study a Merton consumption and portfolio problem with stochastic asset returns and volatilities (see also [21]). Closely related to our work is the Kim-Omberg model [27], which studies the dynamic non-myopic portfolio behavior of an investor trading a risk-free and a risky asset, with expected excess returns following a mean-reverting Ornstein-Uhlenbeck process. Extending this framework to include intermediate consumption, [40] derives exact optimal portfolio and consumption rules. However, unlike our model, [40] assumes complete markets by driving both the asset return and the stochastic factor with the same Brownian motion, does not incorporate labor income, and enforces the non-negativity constraint on wealth only at the terminal time $T$. Furthermore, [18] derive optimal consumption and investment policies in a complete market featuring a stochastic factor, modeled as a general scalar diffusion that drives investment opportunities. In their framework, the absence of labor income ensures that the wealth dynamics remain purely geometric. This in turn allows the value function to be scaled by wealth, thereby reducing the dimension of the HJB equation. Lastly, [19] consider an optimal consumption and investment model with general stochastic factors. They address optimal investment and consumption for a power utility investor using an incomplete stochastic factor model on an infinite horizon and provide a complete characterization for a finite state space. When the factor follows a diffusion process, they develop a new theoretical framework to prove existence and bound the HJB solution, verifying models like the Heston model rigorously for the first time. Conversely to our approach, [19] allows for general stochastic factors, including the Kim-Omberg setting, but assumes geometric wealth without labor income, which automatically enforces the non-negativity of wealth and crucially permits dimension reduction of the associated HJB equation.
In our analysis, the study of a stationary two-dimensional optimal stopping problem plays a crucial role. For optimal stopping problems involving multi-dimensional processes, the standard guess-and-verify approach is generally no longer applicable. This is because the free boundary separating the continuation and stopping regions is no longer a simple scalar threshold, but rather a complex curve or surface, making it practically infeasible to postulate a parameterized closed-form solution for both the value function and the boundary simultaneously. Consequently, a direct study of the problem’s value function and the corresponding variational inequality must be performed on a case-by-case basis via probabilistic methods and/or techniques from the theory of partial differential equations. Notable contributions in this direction, with applications ranging from optimal dividend distribution to public debt reduction and quickest detection, include [4], [5], [7], [9], [16], and [25], among others.
A major mathematical hurdle in our analysis arises from the presence of the stochastic factor, which explicitly introduces stochastic volatility into the optimal stopping problem. Beyond our specific model, optimal stopping under stochastic volatility is known to be highly challenging in general. In standard problems driven by uniformly elliptic diffusions, the value function typically enjoys strong smoothing properties across the entire state space. However, stochastic volatility breaks uniform ellipticity, rendering the associated infinitesimal generator degenerate. This degeneracy makes it particularly challenging to establish the global regularity of the value function and to rigorously characterize the behavior of the free boundary. For example, comparison theorems for solutions to SDEs, which are usually employed in optimal stopping theory to show monotonicity of the optimal stopping boundary, are generally not helpful in settings with stochastic volatility. As a consequence, it is particularly difficult to provide fine smoothness properties of the optimal stopping problem’s value function.
Because of these severe analytical challenges, the overall body of literature addressing optimal stopping under stochastic volatility remains relatively sparse, with a few notable contributions. [1] proves, via purely probabilistic techniques, the monotonicity and continuity of the value function for an optimal stopping problem featuring stochastic volatility; [17] discusses the superreplication of derivatives in a stochastic volatility model under the additional assumption that the volatility follows a bounded process; [39] considers a stochastic volatility model for the asset price underlying an American option, extending regularity results for the American put option price function and proving that the optimal exercise boundary is a decreasing function of the current volatility process realization. Finally, [29] provides an analytical characterization of the price function of an American option in Heston-type models using an approach based on variational inequalities, while [28] studies important properties of the American option price in the stochastic volatility Heston model, including monotonicity and smoothness of the value function, as well as its early-exercise-premium representation.
Structure of the Paper. The rest of the paper is organized as follows. In Section 2, we introduce the primal problem and show how the no-borrowing constraint can be transformed into a static budget constraint. Section 3 derives the associated dual problem using duality principles and obtains the corresponding singular control formulation. We also establish a probabilistic link between this singular control problem and an auxiliary optimal stopping problem, which is fully analyzed in Sections 3.2, 3.3 and 3.4. Section 4 uses the dual formulation to recover the optimal consumption and investment strategies in the primal problem. Finally, we provide numerical illustrations in Section 5.
2. The Primal Problem
2.1. The Financial Market and the Agent’s Problem
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space and denote by $\mathbb{E}[\cdot]$ the expectation under $\mathbb{P}$. Consider an agent whose goal is to maximize the intertemporal expected utility functional given by
$\mathbb{E}\left[\int_{0}^{\infty}e^{-\delta t}u(c_{t})\,dt\right],$
where $\delta>0$ denotes the discount rate, and the utility function $u$ is of power type; that is,
$u(c)=\frac{c^{1-\gamma}}{1-\gamma},$
where $\gamma>1$ represents the agent’s risk aversion. This value of $\gamma$ is supported by empirical evidence on individual time preferences. Numerous studies, such as those by [6] and [30], document risk aversion rates well above 1 based on experimental and survey data. This behavior reflects strong present-biased preferences, which aligns with the assumption $\gamma>1$ in our model.
The market is described by the so-called Kim-Omberg model (see [27]), as it follows. The agent can invest in a risk-free asset with a risk-free rate $r>0$ as well as a risky asset $(S_{t})_{t}$ whose dynamics are given by
$dS_{t}=(r+\beta_{t})S_{t}\,dt+\sigma S_{t}\,dW_{t},\quad t>0,\quad S_{0}=s>0,$
where $(W_{t})_{t}$ is a standard Brownian motion on $(\Omega,\mathcal{F},\mathbb{P})$ generating the filtration (completed by -$\mathbb{P}\text{-null sets of}\;\mathcal{F}$) $\mathbb{F}^{W}:=(\mathcal{F}_{t}^{W})_{t}$, and $\sigma>0$ denotes the volatility of the risky asset. The expected return of the risky asset is $\mu(\beta_{t}):=r+\beta_{t}$, where the process $(\beta_{t})_{t}$ represents the expected excess return of $(S_{t})_{t}$. Asset pricing studies suggest that expected excess returns are predictable and tend to revert to their long-run mean (see [14] and [37], among others). To capture this mean-reversion, we define $(\beta_{t})_{t}$ as
$d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
where $\kappa>0$ is the speed of mean reversion, $\overline{\beta}>0$ is the equilibrium level, and $\sigma_{\beta}>0$ is the volatility of $(\beta_{t})_{t}$. The standard Brownian motion $(W^{\beta}_{t})_{t}$ on $(\Omega,\mathcal{F},\mathbb{P})$ is correlated with $(W_{t})_{t}$ with correlation parameter $\rho\in[-1,1]$, and generates the filtration (completed by -$\mathbb{P}\text{-null sets of}\;\mathcal{F}$) $\mathbb{F}^{W^{\beta}}:=(\mathcal{F}_{t}^{W^{\beta}})_{t}$. Due to correlation, we may write, for $t\geq 0$,
$W_{t}=\rho W^{\beta}_{t}+\sqrt{1-\rho^{2}}W^{1,\perp}_{t},$
and
$W^{\beta}_{t}=\rho W_{t}+\sqrt{1-\rho^{2}}W^{2,\perp}_{t},$
where the independent standard Brownian motions $(W^{1,\perp}_{t})_{t}$ and $(W^{2,\perp}_{t})_{t}$ are such that $W$ is independent of $W^{2,\perp}$, and $W^{\beta}$ is independent of $W^{1,\perp}$. By definition, the excess return $(\beta_{t})_{t}$ reverts to its long-run average $\overline{\beta}$.
The agent chooses a consumption plan $(c_{t})_{t}$, with $c_{t}\geq 0$, and an investment strategy for the risky asset, denoted by $(\pi_{t})_{t}$. Additionally, the agent receives a flow of constant labor income $\ell>0$. The agent’s wealth $(X_{t})_{t}$ thus follows the dynamics
(2.1) $dX_{t}=\left(rX_{t}+\beta_{t}\pi_{t}-c_{t}+\ell\right)\,dt+\sigma\pi_{t}\,dW_{t},\quad t>0,$
with initial wealth $X_{0}=x>0$.
As usual, we define the market price of risk as
$\theta(\beta_{t}):=\frac{\mu(\beta_{t})-r}{\sigma}=\frac{\beta_{t}}{\sigma},\quad t\geq 0,$
and we also introduce the process $(\mathcal{H}_{t})_{t}$, which acts as a stochastic discount factor, as follows
$\mathcal{H}_{t}:=\exp\left(-\int_{0}^{t}\left[r+\frac{1}{2}\frac{\beta_{s}^{2}}{\sigma^{2}}\right]ds-\int_{0}^{t}\frac{\beta_{s}}{\sigma}\,dW_{s}\right),\quad t\geq 0;$
equivalently,
(2.2) $d\mathcal{H}_{t}=-r\mathcal{H}_{t}\,dt-\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}\,dW_{t},\quad t>0,\quad\mathcal{H}_{0}=1.$
We denote by $\bar{\mathbb{F}}:=(\bar{\mathcal{F}}_{t})_{t}$ the completed filtration jointly generated by $W$ and $W^{\beta}$, that is, $\bar{\mathbb{F}}:=\mathbb{F}^{W}\vee\mathbb{F}^{\beta}$, and make the following standing assumption.
Assumption 2.1.
We assume $\kappa\sigma>\sigma_{\beta}$.
The requirement $\kappa\sigma>\sigma_{\beta}$ in Assumption 2.1 implies (by the Novikov condition; see, e.g., Corollary 5.13 on p. 199 in [26]) that
(2.3) $M_{t}:=\exp\bigg(-\int_{0}^{t}\frac{1}{2}\frac{\beta_{s}^{2}}{\sigma^{2}}ds-\int_{0}^{t}\frac{\beta_{s}}{\sigma}\,dW_{s}\bigg)=e^{rt}\mathcal{H}_{t}$
is an $\bar{\mathbb{F}}$- martingale under $\mathbb{P}$, with $(\mathcal{H}_{t})_{t}$ as in (2.2).
The agent faces a no-borrowing constraint; that is, $X_{t}\geq 0$ $\mathbb{P}$-a.s. for all $t\geq 0$. This implies that the agent cannot borrow against future labor income and motivates the following definition of admissible controls.
Definition 2.2.
We call the pair of controls $(\pi,c)$ admissible if:
  1. (1)
    $(c_{t})_{t}$ and $(\pi_{t})_{t}$ are $\bar{\mathbb{F}}$-progressively measurable, and are such that $c_{t}\geq 0$ $\mathbb{P}$-a.s. for all $t\geq 0$, $\int_{0}^{T}c_{s}\,ds<\infty$ and $\int_{0}^{T}\pi_{s}^{2}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$.
  2. (2)
    $X_{t}\geq 0$ for all $t\geq 0$ $\mathbb{P}$-a.s.
We denote by $\mathcal{A}(x)$ the set of admissible controls.
The agent’s optimization problem then reads as
(2.4) $\sup_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}\left[\int_{0}^{\infty}e^{-\delta t}u(c_{t})\,dt\right].$
Remark 2.3.
The recent [19] treats a general stochastic-factor model (including the Kim–Omberg setting) but works in a geometric-wealth framework without labor income. This framework automatically ensures the nonnegativity of wealth and allows for a dimension reduction of the associated HJB equation. By contrast, since we consider total investment and consumption and explicitly include labor income, these simplifications are no longer available. Consequently, we adopt the duality–singular control–optimal stopping approach developed in the following sections.
2.2. From a Dynamic to a Static Budget Constraint
Following [24] (see also [12] and [22] for earlier studies), we now transform the dynamic budget constraint $X_{t}\geq 0$ $\mathbb{P}$-a.s. $\forall$ $t\geq 0$ into a single, static budget constraint. To that end, we define
(2.5) $\mathcal{D}:=\Big\{(D_{t})_{t\geq 0}:D\text{ is }\bar{\mathbb{F}}\text{-adapted, nonnegative, nonincreasing, càdlàg, and }D_{0^{-}}=1\;\mathbb{P}\text{-a.s.}\Big\},$
and we then have the following result.
Proposition 2.4.
  1. (1)
    Let $(c_{t})_{t}$ be a consumption plan such that $(\pi,c)\in\mathcal{A}(x)$. Then it also satisfies the constraint
    (2.6) $\sup_{D\in\mathcal{D}}\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right]\leq x.$
    Moreover, we have $\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}|c_{t}-\ell|\,dt\right]<\infty$.
  2. (2)
    For any nonnegative $\bar{\mathbb{F}}$-progressively measurable $(c_{t})_{t}$ with $\int_{0}^{T}c_{s}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$, and such that
    (2.7) $\sup_{D\in\mathcal{D}}\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right]=x,$
    there exists a process $(\pi_{t})_{t}$ such that $(\pi,c)\in\mathcal{A}(x)$.
Proof.
The proof borrows arguments from [24].
Proof of (1). Let $(c_{t})_{t}$ be a consumption plan such that $(\pi,c)\in\mathcal{A}(x)$ and let $D\in\mathcal{D}$. An application of Itô’s formula for semimartingales (see Theorem 32 on p. 78 in [38]) yields
(2.8) $d(\mathcal{H}_{t}X_{t}D_{t-})=\mathcal{H}_{t}D_{t}(\ell-c_{t})dt+\mathcal{H}_{t}D_{t}(\sigma\pi_{t}-\frac{\beta_{t}}{\sigma}X_{t})dW_{t}+\mathcal{H}_{t}X_{t}dD_{t-}.$
Next, we define the localizing sequence of stopping times
$\tau_{n}:=\inf\bigg\{s\geq 0:\int_{0}^{s}\mathcal{H}^{2}_{t}D^{2}_{t}\bigg|\sigma\pi_{t}-\frac{\beta_{t}}{\sigma}X_{t}\bigg|^{2}dt\geq n\bigg\},\quad n\in\mathbb{N},$
and note that $\tau_{n}\rightarrow\infty$ a.s., as $n\rightarrow\infty$. Integrating (2.8) and rearranging terms we find
$\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(c_{s}-\ell)ds-x$ $=\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(\sigma\pi_{s}-\frac{\beta_{s}}{\sigma}X_{s})dW_{s}$
$\quad+\int_{0}^{\tau_{n}}\mathcal{H}_{s}X_{s}dD_{s-}-\mathcal{H}_{\tau_{n}}X_{\tau_{n}}D_{\tau_{n}-},$
from which, by taking expectations,
$\mathbb{E}\bigg[\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(c_{s}-\ell)ds\bigg]-x$
(2.9) $\qquad=\mathbb{E}\bigg[\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(\sigma\pi_{s}-\frac{\beta_{s}}{\sigma}X_{s})dW_{s}+\int_{0}^{\tau_{n}}\mathcal{H}_{s}X_{s}dD_{s-}-\mathcal{H}_{\tau_{n}}X_{\tau_{n}}D_{\tau_{n}-}\bigg]$
$\qquad=\underbrace{\mathbb{E}\bigg[\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(\sigma\pi_{s}-\frac{\beta_{s}}{\sigma}X_{s})dW_{s}\bigg]}_{=0}+\underbrace{\mathbb{E}\bigg[\int_{0}^{\tau_{n}}\mathcal{H}_{s}X_{s}dD_{s-}\bigg]}_{\leq 0}-\underbrace{\mathbb{E}\bigg[\mathcal{H}_{\tau_{n}}X_{\tau_{n}}D_{\tau_{n}-}\bigg]}_{\geq 0}$
$\qquad\leq 0.$
Here we used the facts that $\mathcal{H}_{t}\geq 0$, $X_{t}\geq 0$ since $(\pi,c)\in\mathcal{A}(x)$, and that $D$ is nonincreasing and nonnegative. Applying Fatou’s lemma for lower bounded functions, together with (2.2), yields
$\mathbb{E}\bigg[\int_{0}^{\infty}\mathcal{H}_{s}D_{s}(c_{s}-\ell)ds\bigg]-x$ $=\mathbb{E}\bigg[\liminf_{n\rightarrow\infty}\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(c_{s}-\ell)ds\bigg]-x$
$\leq\liminf_{n\rightarrow\infty}\mathbb{E}\bigg[\int_{0}^{\tau_{n}}\mathcal{H}_{s}D_{s}(c_{s}-\ell)ds\bigg]-x\leq 0;$
hence,
$\mathbb{E}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)dt\bigg]\leq x,$
and by arbitrariness of $D\in\mathcal{D}$, we obtain (2.6); that is,
$\sup_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)dt\bigg]\leq x.$
In particular, since $1\in\mathcal{D}$, we have
$\mathbb{E}\!\left[\int_{0}^{\infty}\mathcal{H}_{t}(c_{t}-\ell)\,dt\right]\leq x.$
Using the fact that $\mathcal{H}_{t}=e^{-rt}M_{t}$, where $(M_{t})_{t}$ is the martingale given by (2.3), we also find
$\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}c_{t}\,dt\right]\leq x+\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}\ell M_{t}\,dt\right]=x+\frac{\ell}{r},$
and therefore,
(2.10) $\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}|c_{t}-\ell|\,dt\right]\leq x+\frac{2\ell}{r}<\infty.$
Proof of (2). For $T>0$ arbitrary but fixed, consider now the reflected backward stochastic differential equation (RBSDE):
(2.11) $\begin{cases}\tilde{X}_{s}=\int_{s}^{T}\mathcal{H}_{t}(c_{t}-\ell)\,dt-\int_{s}^{T}\tilde{Z}_{t}\,dW_{t}+\int_{s}^{T}d\tilde{K}_{t},\quad\forall s\geq 0,\\ \tilde{X}_{t}\geq 0,\quad\int_{0}^{T}\tilde{X}_{t}\,d\tilde{K}_{t}=0.\end{cases}$
[…]
$\mathbb{E}\left[\sup_{t\in[0,T]}|\tilde{X}_{t}|^{\nu}\right]<\infty\quad\text{and}\quad\mathbb{E}\left[\left(\int_{0}^{T}|\tilde{Z}_{t}|^{2}\,dt\right)^{\nu/2}\right]<\infty.$
Note that the driver of the RBSDE (2.11), namely $\mathcal{H}_{t}(c_{t}-\ell)$, does not depend on $\tilde{Z}$ and $\tilde{X}$. Furthermore, due to (2.10), we have for all $T>0$
$\mathbb{E}\left[\int_{0}^{T}\mathcal{H}_{t}|c_{t}-\ell|\,dt\right]<\infty.$
[…]
Let then $D\in\mathcal{D}$ and define the sequence of localizing stopping times
$\vartheta_{n}:=\inf\left\{s\geq 0:\int_{0}^{s}\tilde{Z}_{u}^{2}D_{u}^{2}\,du\geq n\right\},\quad n\in\mathbb{N}.$
Note that $\vartheta_{n}\to\infty$ $\mathbb{P}$-a.s. as $n\to\infty$. Using Itô’s formula on the interval $0,T\wedge\vartheta_{n}]$ and taking expectations, we obtain
(2.12) $\tilde{X}_{0}$ $=\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}D_{T\wedge\vartheta_{n}}+\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T\wedge\vartheta_{n}}D_{t-}\,d\tilde{K}_{t}-\int_{0}^{T\wedge\vartheta_{n}}\tilde{X}_{t}\,dD_{t-}\right].$
Since $D$ is decreasing and $\tilde{K}$ is increasing, we have from (2.12)
$\tilde{X}_{0}\geq\mathbb{E}\left[\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right],$
and letting $n\to\infty$ and $T\to\infty$, (2.10) allows us to invoke the Dominated Convergence Theorem to obtain
$\tilde{X}_{0}\geq\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right].$
Since $D\in\mathcal{D}$ was arbitrary, it follows from (2.7) that
(2.13) $\tilde{X}_{0}\geq x.$
On the other hand, let
$\tilde{D}_{s}:=\mathbf{1}_{\{s<\tilde{\eta}\}},\;s\geq 0,\quad\tilde{\eta}:=\inf\left\{s\geq 0:\tilde{X}_{s}\leq 0\right\}\wedge(T\wedge\vartheta_{n}).$
Then,
$\tilde{X}_{\tilde{\eta}}\mathbf{1}_{\{\tilde{\eta}<T\wedge\vartheta_{n}\}}=0,\quad\tilde{K}_{t}=0\quad\text{on}\quad 0\leq t<\tilde{\eta}$,
and
$\int_{0}^{T\wedge\vartheta_{n}}\tilde{D}_{t-}\,d\tilde{K}_{t}=\int_{0}^{T\wedge\vartheta_{n}}\tilde{X}_{t}\,d\tilde{D}_{t-}=0,$
so that (2.12) implies
(2.14) $\tilde{X}_{0}=\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}\tilde{D}_{T\wedge\vartheta_{n}}+\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}\tilde{D}_{t}(c_{t}-\ell)\,dt\right].$
Since $\tilde{X}$ belongs to the Class D, we have $\lim_{n\to\infty}\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}\right]=\mathbb{E}\left[\tilde{X}_{T}\right]=0$. Hence, using the Dominated Convergence Theorem again to take limits as $n\rightarrow\infty$ and $T\rightarrow\infty$ in (2.14), we obtain
(2.15) $\tilde{X}_{0}\leq\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}\tilde{D}_{t}(c_{t}-\ell)\,dt\right]\leq x,$
where we used the fact that $\tilde{D}\in\mathcal{D}$. Therefore, combining (2.13) and (2.15), we conclude that $\tilde{X}_{0}=x$.
From the RBSDE (2.11), we know that $\tilde{X}$ satisfies
$0\leq\tilde{X}_{s}+\int_{0}^{s}d\tilde{K}_{t}=x+\int_{0}^{s}\mathcal{H}_{t}(\ell-c_{t})\,dt+\int_{0}^{s}\tilde{Z}_{t}\,dW_{t},\quad\forall s\geq 0.$
If we now choose $\pi$ defined as
(2.16) $\pi_{t}:=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}X^{\pi,c}_{t}\mathcal{H}_{t}\right),\quad t\geq 0,$
where $(X^{\pi,c}_{t})_{t}$ is the wealth process controlled through $(\pi,c)$, we obtain for any $s\geq 0$
$0\leq\tilde{X}_{s}+\int_{0}^{s}d\tilde{K}_{t}=x+\int_{0}^{s}\mathcal{H}_{t}(\ell-c_{t})\,dt+\int_{0}^{s}\mathcal{H}_{t}\left(\sigma\pi_{t}-\frac{\beta_{t}}{\sigma}X^{\pi,c}_{t}\right)\,dW_{t}.$
We need to check that such a $\pi$ is indeed such that $(\pi,c)\in\mathcal{A}(x)$. Observe that the expression on the right-hand side of the last display equation is exactly the discounted wealth $X^{\pi,c}_{s}\mathcal{H}_{s}$. Since $\mathcal{H}_{s}$ is strictly positive for all $s\geq 0$, we thus get $X^{\pi,c}_{s}\geq 0$ for all $s\geq 0$.
Let us now check that
$\int_{0}^{T}\pi_{t}^{2}\,dt<\infty\quad\mathbb{P}-a.s.\ \text{for all}\;T>0.$
Since $X_{t}^{\pi,c}=\mathcal{H}_{t}^{-1}(\tilde{X}_{t}+\tilde{K}_{t})$, by substituting this expression into (2.16), we find
$\pi_{t}=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}(\tilde{X}_{t}+\tilde{K}_{t})\right).$
As $(\mathcal{H}_{t})_{t}$, $(\beta_{t})_{t}$, $(\tilde{X}_{t})_{t}$ and $(\tilde{K}_{t})_{t}$ are continuous, these processes are $\mathbb{P}$-a.s. locally bounded. Furthermore, $\int_{0}^{T}\tilde{Z}_{t}^{2}dt<\infty$ $\mathbb{P}$-a.s. for all $T>0$, by definition of the $L^{1}$-solution to (2.11). Thus, $\int_{0}^{T}\pi_{t}^{2}dt<\infty$ $\mathbb{P}$-a.s. for all $T>0$, which implies $(\pi,c)\in\mathcal{A}(x)$. ∎
Remark 2.5.
Note that since $1\in\mathcal{D}$, we in particular have from Proposition 2.4-$(1)$
$\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}(c_{t}-\ell)\,dt\right]\leq x.$
This means that the present value of discounted consumption, net of labor income, cannot exceed the agent’s initial wealth. However, this condition alone does not guarantee that the wealth process remains nonnegative with probability one at all times. Therefore, we need a stronger requirement, which is achieved by using the processes $D$ as in Proposition 2.4 above.
3. The Dual Problem as a Singular Control Problem
3.1. Derivation of the Dual Problem
In this section, we derive the dual problem expected to be associated to (3.11). Because (cf. (2.6))
$\sup_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)dt\bigg]\leq x,$
with $(\mathcal{H}_{t})_{t}$ as in (2.2), for a given $z>0$ we can write
$\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]$ $\leq\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]+z\bigg(x-\sup_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)dt\bigg]\bigg)$
(3.1) $=\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]+zx+\inf_{D\in\mathcal{D}}\mathbb{E}\bigg[-\int_{0}^{\infty}z\mathcal{H}_{t}D_{t}(c_{t}-\ell)dt\bigg]$
$=\inf_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(u(c_{t})-e^{\delta t}z\mathcal{H}_{t}D_{t}(c_{t}-\ell)\big)dt\bigg]+zx.$
To simplify notation, in the following we set
(3.2) $Z^{D}_{t}:=e^{\delta t}z\mathcal{H}_{t}D_{t},\quad Z_{0^{-}}=z>0.$
Given that the Legendre–Fenchel transform of the utility function $u$ is such that
(3.3) $\tilde{u}(z):=\sup_{c\geq 0}\left(u(c)-zc\right)=\sup_{c\geq 0}\bigg(\frac{c^{1-\gamma}}{1-\gamma}-zc\bigg)=\frac{\gamma}{1-\gamma}z^{-\frac{1-\gamma}{\gamma}},$
using $u(c_{t})-Z^{D}_{t}c_{t}\leq\tilde{u}(Z^{D}_{t})$ in (3.1), we obtain
(3.4) $\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]\leq\inf_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)dt\bigg]+zx,$
which, by arbitrariness of $z>0$, in turn yields
$\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]\leq\inf_{\begin{subarray}{c}z>0\\ D\in\mathcal{D}\end{subarray}}\bigg(\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)dt\bigg]+zx\bigg).$
Hence, we have the weak duality
(3.5) $\sup_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]\leq\inf_{z>0}\bigg(\inf_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)dt\bigg]+zx\bigg).$
In the subsequent analysis, we shall focus on the problem
(3.6) $\inf_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)dt\bigg],$
with the aim of proving that actually the strong duality
(3.7) $\sup_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})dt\bigg]=\inf_{z>0}\bigg(\inf_{D\in\mathcal{D}}\mathbb{E}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)dt\bigg]+zx\bigg)$
holds true. Problem (3.6) is a singular stochastic control problem for the two-dimensional state process
(3.8) $dZ^{D}_{t}=(\delta-r)Z^{D}_{t}\,dt-\frac{\beta_{t}}{\sigma}Z^{D}_{t}\,dW_{t}+Z^{D}_{t}\frac{dD_{t}}{D_{t}},\quad t>0,\quad Z_{0^{-}}^{D}=z>0,$
(3.9) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
with $D\in\mathcal{D}$. Notice that the dynamics of $(Z_{t}^{D})_{t}$ is easily obtained from (3.2) via Itô’s formula.
Given the Markovian structure, from now on we stress the dependency of the value of (3.6) with respect to the problem’s initial data and write
(3.10) $\tilde{V}(z,\beta):=\inf_{D\in\mathcal{D}}\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)\,dt\bigg],$
where $\mathbb{E}_{z,\beta}[\;\cdot\;]=\mathbb{E}[\;\cdot\;|Z_{0^{-}}^{D}=z,\beta_{0}=\beta]$ denotes the expectation under $\mathbb{P}_{z,\beta}(\cdot):=\mathbb{P}(\;\cdot\;|\;Z^{D}_{0-}=z,\;\beta_{0}=\beta)$. For later use, we notice that $\lim_{z\downarrow 0}\tilde{V}(z,\beta)=0$. Analagously, in the following, we shall write (cf. (2.4))
(3.11) $V(x,\beta):=\max_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}_{x,\beta}\left[\int_{0}^{\infty}e^{-\delta t}u(c_{t})\,dt\right],$
where $\mathbb{E}_{x,\beta}[\cdot]$ denotes the expectation under $\mathbb{P}_{x,\beta}(\cdot):=\mathbb{P}(\;\cdot\;|\;X_{0}=x,\;\beta_{0}=\beta)$.
In Proposition 4.4 below we will show that
$V(x,\beta)=\inf_{z>0}\bigg(\tilde{V}(z,\beta)+zx\bigg),$
so that (3.7) indeed holds true. This will be achieved through a series of intermediate results aimed at characterizing the optimal policy of problem (3.6). A major ingredient towards this characterization is the identification of an optimal stopping problem whose value coincides with $\tilde{V}_{z}$. Such an optimal stopping problem is introduced and studied in the following sections.
3.2. Derivation of the Auxiliary Optimal Stopping Problem
Denote by $(Z^{1}_{t})_{t}$ the uncontrolled state process (i.e., $Z^{D}$ as in (3.8) with $D\equiv 1$) satisfying
(3.12) $dZ^{1}_{t}=(\delta-r)Z^{1}_{t}\,dt-\frac{\beta_{t}}{\sigma}Z^{1}_{t}\,dW_{t},\quad t>0,\quad Z_{0}^{1}=z>0.$
Furthermore, to simplify notation, set $\mathcal{O}:=(0,\infty)\times\mathbb{R}$.
Inspired by [2], [5] and [16] we introduce the optimal stopping problem
(3.13) $v(z,\beta):=\inf_{\tau}\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}M_{t}\,(\tilde{u}^{\prime}(Z_{t}^{1})+\ell)\,dt\bigg],$
where $\tilde{u}^{\prime}(z)=-z^{1/\gamma}$, and where we take the infimum over $\bar{\mathbb{F}}$-stopping times $\tau\geq 0$, $Z^{1}$ evolves as in (3.12), and $\beta$ as in (3.9). We expect $v$ to be such that $\tilde{V}_{z}=v$ on $\mathcal{O}$. Theorem 4.1 below will indeed prove that such a relation holds true and that an optimizer for $v$ is in one-to-one correspondence to an optimizer for $\tilde{V}$. In the following, we shall study (3.13) and characterize its optimal policy. In order to achieve this, it is convenient to perform a change of measure to remove the martingale $(M_{t})_{t}$ (cf. (2.3)) from the stopping functional. This leads to the next proposition whose proof is postponed to the appendix.
Proposition 3.1.
For $v$ as in (3.13) we have
(3.14) $v(z,\beta)=\inf_{\tau}\mathbb{E}_{z,\beta}^{\mathbb{Q}}\bigg[\int_{0}^{\tau}e^{-rt}\,(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)\,dt\bigg],$
[…]
$d\hat{Z}_{t}$ $=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}\,dW^{\mathbb{Q}}_{t}+\hat{Z}_{t}\bigg(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}}\bigg)\,dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
$d\hat{\beta}_{t}$ $=\sigma_{\beta}\,dW^{\beta,\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
[…]
Proof.
See Appendix A.1. ∎
With reference to Proposition 3.1, we therefore now turn our attention to characterizing the solution to the optimal stopping problem
(3.15) $v(z,\beta):=\inf_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)dt\bigg],$
subject to
(3.16) $d\hat{Z}_{t}$ $=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}dW^{\mathbb{Q}}_{t}+\hat{Z}_{t}(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}})dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
(3.17) $d\hat{\beta}_{t}$ $=\sigma_{\beta}dW_{t}^{\beta,\mathbb{Q}}+(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta})dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
In the following, when needed, we stress the dependence of the unique strong solution to (3.16)-(3.17) on the initial data $(z,\beta)\in\mathcal{O}$ by writing $(\hat{Z}_{t}^{z,\beta})_{t}$ and $(\hat{\beta}_{t}^{\beta})_{t}$.
3.3. Preliminary Properties of the Optimal Stopping Value Function
In this subsection, we establish preliminary properties of the value function (3.15). For the proof of those, we make the following assumption on the model’s parameters. Such requirement in particular ensures well-posedness of $v$ as in (3.15) and, together with Assumption 2.1, it will be a standing assumption throughout the rest of the paper.
Assumption 3.2.
We assume that
$\gamma>\max\bigg\{1,\frac{|\rho|\sigma_{\beta}}{\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}\bigg\}.$
Notice that $\kappa+\frac{\rho\sigma_{\beta}}{\sigma}>0$ due to Assumption 2.1. We then have the following first preliminary finding.
Proposition 3.3.
It holds
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}|\tilde{u}^{\prime}(\hat{Z}_{t})+\ell|\,dt\bigg]<\infty.$
Moreover, one has
(3.18) $\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]\leq z^{-\frac{1}{\gamma}}\exp\bigg(-\tfrac{1}{\gamma}(\delta-r)t\bigg).$
Proof.
See Appendix A.2. ∎
The next result directly follows from the expression of $v$ as in (3.15) and the fact that $z\mapsto\hat{Z}^{z,\beta}_{t}$ is $\mathbb{Q}$-a.s. increasing for all $t\geq 0$.
Proposition 3.4.
One has that $z\mapsto v(z,\beta)$ is nondecreasing for all $\beta\in\mathbb{R}$.
Remark 3.5.
Note that the monotonicity of $\beta\mapsto v(z,\beta)$ is not clear, since the process $(\hat{\beta}_{t})_{t}$ also affects the volatility of $(\hat{Z}_{t})_{t}$, and therefore comparison theorems for solutions to SDEs do not apply.
The next results provide useful bounds and limit behavior of the value function $v$. Their proofs are given in the Appendix.
Proposition 3.6.
We have
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]\leq v(z,\beta)\leq 0.$
Proof.
See Appendix A.3. ∎
Proposition 3.7.
It holds that
$\lim_{z\to 0}v(z,\beta)=-\infty\quad\text{and}\quad\lim_{z\to\infty}v(z,\beta)=0.$
Proof.
See Appendix A.4. ∎
Proposition 3.8.
The value function $v(z,\beta)$ is continuous for all $(z,\beta)\in\mathcal{O}$.
Proof.
Step 1. We first show that $(z,\beta)\mapsto v(z,\beta)$ is upper semicontinuous for all $(z,\beta)\in\mathcal{O}$. Given (3.15), it suffices to show that
$\mathcal{J}(\tau;z,\beta):=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}\big(-\hat{Z}_{t}^{-\frac{1}{\gamma}}+\ell\big)dt\bigg]$
is continuous for all $(z,\beta)\in\mathcal{O}$ and fixed $\tau\geq 0$.
Fix $(z_{0},\beta_{0})\in\mathcal{O}$ and let $(z_{n},\beta_{n})\subseteq\mathcal{O}$ be a sequence converging to $(z_{0},\beta_{0})$. Then
(3.19) ${}\big|\mathcal{J}(\tau;z_{n},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{0})\big|$ $\leq\big|\mathcal{J}(\tau;z_{n},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{n})\big|$
$\quad+\big|\mathcal{J}(\tau;z_{0},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{0})\big|.$
For the first term on the right-hand side in (3.19), we have
$\big|\mathcal{J}(\tau;z_{n},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{n})\big|$ $\leq\mathbb{E}^{\mathbb{Q}}\bigg[\int_{0}^{\infty}e^{-rt}\big|(z_{0}\tilde{Z}_{t})^{-\frac{1}{\gamma}}-(z_{n}\tilde{Z}_{t})^{-\frac{1}{\gamma}}\big|dt\bigg]$
(3.20) $=|z_{0}^{-\frac{1}{\gamma}}-z_{n}^{-\frac{1}{\gamma}}|\,\mathbb{E}^{\mathbb{Q}}\bigg[\int_{0}^{\infty}e^{-rt}\tilde{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]$
$\leq|z_{0}^{-\frac{1}{\gamma}}-z_{n}^{-\frac{1}{\gamma}}|\int_{0}^{\infty}e^{-rt}e^{-\frac{1}{\gamma}(\delta-r)t}dt,$
where we have set
(3.21) $\tilde{Z}_{t}:=\exp\Big(\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds-\int_{0}^{t}\frac{\hat{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\Big),$
for $t\geq 0$, and we have used (3.18) upon noticing that $\tilde{Z}_{t}=\frac{\hat{Z}_{t}}{z}$. Given that the integral on the right-hand side of (3.3) is finite as $r+\frac{1}{\gamma}(\delta-r)>0$ because $\gamma>1$, we find
(3.22) $\lim_{(z_{n},\beta_{n})\to(z_{0},\beta_{0})}\big|\mathcal{J}(\tau;z_{n},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{n})\big|=0.$
For the second term in (3.19), Fubini-Tonelli Theorem yields
(3.23) $\big|\mathcal{J}(\tau;z_{0},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{0})\big|$ $\leq\int_{0}^{\infty}e^{-rt}\,\mathbb{E}^{\mathbb{Q}}\Big[\big|(\hat{Z}_{t}^{z_{0},\beta_{0}})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z_{0},\beta_{n}})^{-\frac{1}{\gamma}}\big|\Big]dt.$
By the same arguments used in the proof of Proposition 3.3 (see Appendix A.2), one can show that
$e^{-rt}\,\mathbb{E}^{\mathbb{Q}}\Big[\big|(\hat{Z}_{t}^{z_{0},\beta_{0}})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z_{0},\beta_{n}})^{-\frac{1}{\gamma}}\big|\Big]\leq 2z_{0}^{-\frac{1}{\gamma}}e^{-(r+\frac{1}{\gamma}(\delta-r))t},$
with
$\int_{0}^{\infty}2z_{0}^{-\frac{1}{\gamma}}e^{-(r+\frac{1}{\gamma}(\delta-r))t}dt<\infty,$
given that $r+\frac{1}{\gamma}(\delta-r)>0$ by $\gamma>1$. Hence, an application of the Dominated Convergence Theorem in (3.23) yields
(3.24) $\lim_{(z_{n},\beta_{n})\to(z_{0},\beta_{0})}\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}\Big[\big|(\hat{Z}_{t}^{z_{0},\beta_{0}})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z_{0},\beta_{n}})^{-\frac{1}{\gamma}}\big|\Big]dt$
(3.25) $=\int_{0}^{\infty}\lim_{(z_{n},\beta_{n})\to(z_{0},\beta_{0})}e^{-rt}\mathbb{E}^{\mathbb{Q}}\Big[\big|(\hat{Z}_{t}^{z_{0},\beta_{0}})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z_{0},\beta_{n}})^{-\frac{1}{\gamma}}\big|\Big]dt.$
Finally, by exploiting arguments as in the proof of Proposition 3.3 again, we have under Assumption 3.2 that
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\big[(\hat{Z}_{t}^{\beta})^{-p/\gamma}\big]\leq z_{0}^{-p/\gamma}e^{-\frac{p}{\gamma}(\delta-r)t}<\infty,$
where $p$ is chosen such that
$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\},$
upon noticing that, by Assumption 3.2, we have $\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)>|\rho|\sigma_{\beta}$. Therefore, by Vitali’s Convergence Theorem and continuity of $\beta\mapsto\hat{Z}_{t}^{z,\beta}$, we conclude that
$\lim_{(z_{n},\beta_{n})\to(z_{0},\beta_{0})}\mathbb{E}^{\mathbb{Q}}\Big[\big|(\hat{Z}_{t}^{z_{0},\beta_{0}})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z_{0},\beta_{n}})^{-\frac{1}{\gamma}}\big|\Big]=0,$
which implies due to (3.23) and (3.24) that
(3.26) $\lim_{(z_{n},\beta_{n})\to(z_{0},\beta_{0})}\big|\mathcal{J}(\tau;z_{0},\beta_{n})-\mathcal{J}(\tau;z_{0},\beta_{0})\big|=0.$
Finally, by combining (3.22) and (3.26), we obtain
$\lim_{(z_{n},\beta_{n})\to(z_{0},\beta_{0})}\mathcal{J}(\tau;z_{n},\beta_{n})=\mathcal{J}(\tau;z_{0},\beta_{0}).$
Step 2. Fix again $(z_{0},\beta_{0})\in\mathcal{O}$ and let $(z_{n},\beta_{n})\subseteq\mathcal{O}$ be a sequence converging to $(z_{0},\beta_{0})$ as $n\to\infty$. For each $(z_{n},\beta_{n})$, let $\tau_{n}$ be an $\varepsilon$-optimal stopping time for $(z_{n},\beta_{n})$, with $\varepsilon>0$; that is,
$v(z_{n},\beta_{n})\geq\mathcal{J}(\tau_{n};z_{n},\beta_{n})-\varepsilon.$
Since $\tau_{n}$ is suboptimal for $(z_{0},\beta_{0})$, we have $v(z_{0},\beta_{0})\leq\mathcal{J}(\tau_{n};z_{0},\beta_{0})$.
Defining
$\Delta_{n}:=\mathcal{J}(\tau_{n};z_{n},\beta_{n})-\mathcal{J}(\tau_{n};z_{0},\beta_{0}),$
it then holds
$|\Delta_{n}|\leq|\mathcal{J}(\tau_{n};z_{n},\beta_{n})-\mathcal{J}(\tau_{n};z_{0},\beta_{n})|+|\mathcal{J}(\tau_{n};z_{0},\beta_{n})-\mathcal{J}(\tau_{n};z_{0},\beta_{0})|.$
By arguments as in Step 1 above, we have $\Delta_{n}\to 0$ as $n\to\infty$. Hence,
$v(z_{n},\beta_{n})\geq\mathcal{J}(\tau_{n};z_{n},\beta_{n})-\varepsilon=\mathcal{J}(\tau_{n};z_{0},\beta_{0})+\Delta_{n}-\varepsilon\geq v(z_{0},\beta_{0})+\Delta_{n}-\varepsilon,$
which, by taking the limit as $n\to\infty$, yields
$\liminf_{n\to\infty}v(z_{n},\beta_{n})\geq v(z_{0},\beta_{0})-\varepsilon.$
Since $\varepsilon>0$ was arbitrary, we have that $v$ is lower-semicontinuous.
Step 3. Combining Step 1 and Step 2, we conclude that $v$ is continuous on $\mathcal{O}$. ∎
As it is customary in optimal stopping, we now define the continuation (waiting) and stopping regions as
(3.27) $\mathcal{W}:=\{(z,\beta)\in\mathcal{O}:v(z,\beta)<0\},\quad\mathcal{S}:=\{(z,\beta)\in\mathcal{O}:v(z,\beta)=0\}.$
By the continuity of $v$ (see Proposition 3.8), $\mathcal{W}$ is open and $\mathcal{S}$ is closed. Furthermore, the stopping time
(3.28) $\tau^{*}(z,\beta):=\inf\{s\geq 0:(\hat{Z}_{s},\hat{\beta}_{s})\in\mathcal{S}\}$
is optimal (see Corollary 2.9 in Chapter 1 of [35]).
Proposition 3.9.
The stopping region $\mathcal{S}$ is non-empty; that is, $\mathcal{S}\neq\emptyset$.
Proof.
Suppose $\mathcal{S}=\emptyset$. Then for all $(z,\beta)\in\mathcal{O}$ we have by (3.15), the fact that $\tilde{u}^{\prime}(\hat{Z}_{t})=-\hat{Z}_{t}^{-\frac{1}{\gamma}}$ (cf. (3.3)), and (3.18) that
$0>v(z,\beta)=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\big(-\hat{Z}_{t}^{-\frac{1}{\gamma}}+\ell\big)dt\bigg]$ $=\frac{\ell}{r}-\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]$
$\geq\frac{\ell}{r}-z^{-\frac{1}{\gamma}}\int_{0}^{\infty}e^{-(\frac{1}{\gamma}(\delta-r)+r)t}dt$
$=\frac{\ell}{r}-\frac{z^{-\frac{1}{\gamma}}}{\frac{1}{\gamma}(\delta-r)+r}.$
However, the last expression is strictly positive if
(3.29) $z>\Big(\frac{\ell}{r}\big(\frac{1}{\gamma}(\delta-r)+r\big)\Big)^{-\gamma}>0,$
which gives the desired contradiction. ∎
3.4. Optimal Stopping Boundary and Regularity of the Value Function
In this section, we establish the existence of a lower-semicontinuous optimal stopping boundary (free boundary) that separates continuation and stopping regions and prove further regularity of the value function of the optimal stopping problem.
We first show that the boundary $\partial\mathcal{W}$ can be represented by a function $z^{*}:\mathbb{R}\rightarrow[\ell^{-\gamma},\infty]$ and establish connectedness of $\mathcal{W}$ and $\mathcal{S}$ with respect to the $z$-variable.
Lemma 3.10.
There exists a free boundary $z^{*}:\mathbb{R}\rightarrow[0,\infty]$ such that
(3.30) $\mathcal{S}=\{(z,\beta)\in\mathcal{O}:z\geq z^{*}(\beta)\}.$
Moreover, we have
(3.31) $0<\ell^{-\gamma}\leq z^{*}(\beta)\quad\text{for all }\beta\in\mathbb{R},$
and $\beta\mapsto z^{*}(\beta)$ is lower-semicontinuous.
Proof.
By Proposition 3.4, we have that $z\mapsto v(z,\beta)$ is nondecreasing for all $\beta\in\mathbb{R}$. Hence, by defining
(3.32) $z^{*}(\beta):=\inf\{z>0:v(z,\beta)\geq 0\}$
(with the convention $\inf\emptyset=+\infty$), it follows from (3.27) that
$\mathcal{S}=\{(z,\beta)\in\mathcal{O}:z\geq z^{*}(\beta)\},\quad\text{and}\quad\mathcal{W}=\{(z,\beta)\in\mathcal{O}:z<z^{*}(\beta)\}.$
For the lower bound of $z^{*}$, we have from (3.15) and $\tilde{u}^{\prime}(z)=-z^{-\frac{1}{\gamma}}$ that, if $z^{-\frac{1}{\gamma}}+\ell<0$, it is optimal to continue, as stopping immediately yields 0 while continuing for a short time yields a negative contribution to the cost functional. Hence, we obtain
$\{(z,\beta)\in\mathcal{O}\mid-z^{-\frac{1}{\gamma}}+\ell<0\}\subseteq\mathcal{W}$ $\iff\{(z,\beta)\in\mathcal{O}\mid-z^{-\frac{1}{\gamma}}+\ell\geq 0\}\supseteq\mathcal{S}$
$\iff\{(z,\beta)\in\mathcal{O}\mid z\geq\ell^{-\gamma}\}\supseteq\mathcal{S}.$
It then follows from (3.32) that $z^{*}(\beta)\geq\ell^{-\gamma}>0,\;\text{for all}\;\beta\in\mathbb{R}.$
Finally, lower-semicontinuity of $z^{*}$ is due to the fact that (3.30) is closed thanks to (3.27) and continuity of $v$ (cf. Proposition 3.8). ∎
We continue by proving (local) Lipschitz continuity of $v$ and probabilistic representations of its weak derivatives.
Proposition 3.11.
The value function $v$ is (locally) Lipschitz continuous on $\mathcal{O}$. Moreover, its weak derivatives, denoted by $v_{z}$ and $v_{\beta}$, admit the following probabilistic representations:
(3.33) $v_{z}(z,\beta)=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\Big[\int_{0}^{\tau^{*}}e^{-rt}\frac{1}{\gamma}z^{-1}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\Big],$
and
(3.34) $v_{\beta}(z,\beta)=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right],$
where $a:=\kappa+\rho\frac{\sigma_{\beta}}{\sigma}>0$ by Assumption 2.1.
Proof.
See Appendix A.5. ∎
Standard results in optimal stopping theory (cf. Chapter 3 in [35]) together with the previous findings imply that the couple $(v,z^{*})$ satisfies the free boundary problem
(3.35) $\mathcal{L}v-rv+\tilde{u}^{\prime}(z)+\ell=0$ on $0<z<z^{*}(\beta)$
$v=0$ on $z\geq z^{*}(\beta)$,
where $\mathcal{L}$ is the infinitesimal generator of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ such that
(3.36) $(\mathcal{L}v)(z,\beta)=\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}(2zv_{z}+z^{2}v_{zz})+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+(\delta-r)zv_{z}+\kappa(\overline{\beta}-\beta)v_{\beta}-\rho\frac{\beta}{\sigma}\sigma_{\beta}\,(v_{\beta}+zv_{z\beta}),$
and the PDE above is intended in the sense of Schwartz distributions (see Corollary 5 in [36]). In Proposition 3.13 below we will show that $v$ actually solves (3.35) in the classical sense. In order to achieve this, we need the following result.
Lemma 3.12.
The process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$, given by (3.16) and (3.17), is strong Feller.
Proof.
Given that $Corr(W_{t}^{\mathbb{Q}},W_{t}^{\beta,\mathbb{Q}})=\rho t$, $\rho\in[-1,1]$, we can write
$W_{t}^{\beta,\mathbb{Q}}=\rho\,W_{t}^{\mathbb{Q}}+\sqrt{1-\rho^{2}}\,\widehat{W}_{t}^{2,\perp},\quad t\geq 0,$
where $\widehat{W}^{2,\perp}$ is a standard Brownian motion independent of $W^{\mathbb{Q}}$. Then (cf. (3.16) and (3.17)),
$d\hat{Z}_{t}=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}\,dW_{t}^{\mathbb{Q}}+\hat{Z}_{t}\bigg(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}}\bigg)dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
and
$d\hat{\beta}_{t}=\sigma_{\beta}\rho\,dW_{t}^{\mathbb{Q}}+\sigma_{\beta}\sqrt{1-\rho^{2}}\,d\widehat{W}_{t}^{2,\perp}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
Notice that (3.36) is not uniformly elliptic. As a matter of fact, denoting by $\Sigma(z,\beta)$ the diffusion matrix associated to (3.16)-(3.17) one has $\det(\Sigma\Sigma^{T}(z,\beta))=\frac{\beta^{2}z^{2}\sigma_{\beta}^{2}}{\sigma^{2}}(1-\rho^{2})$, which vanishes for $\beta=0$. We therefore now check that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the so-called Hörmander’s condition (see, e.g., condition (H) in Section 2.3.2 in [34]). This implies that the second-order infinitesimal generator $\mathcal{L}$ of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ as in (3.36) is hypoelliptic and therefore that $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is a strong Feller process (see Proposition 4 in [13]).
For $(z,\beta)\in\mathcal{O}$, arbitrary but fixed, we then define the Stratonovich-corrected drift vector field and the diffusion vector fields driven by $W^{\mathbb{Q}}$ and $\widehat{W}^{2,\perp}$ as follows:
$V^{(0)}(z,\beta)=\begin{pmatrix}\dfrac{z\big(\beta^{2}+\rho\sigma\sigma_{\beta}+2\sigma^{2}(\delta-r)\big)}{2\sigma^{2}}\\[8.0pt] \dfrac{-\beta\rho\sigma_{\beta}+\kappa\sigma(\overline{\beta}-\beta)}{\sigma}\end{pmatrix},\qquad V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{\beta}{\sigma}z\\[4.0pt] \sigma_{\beta}\rho\end{pmatrix}$,
$V^{(2)}(z,\beta)=\begin{pmatrix}0\\[4.0pt] \sigma_{\beta}\sqrt{1-\rho^{2}}\end{pmatrix}$.
First, consider the case $|\rho|<1$. We denote by $DV^{(i)}$ the Jacobian matrix of the vector $V^{(i)}$, $i=0,1,2$. The Lie-bracket between $V^{(1)}$ and $V^{(2)}$, denoted by $V^{(1)},V^{(2)}]$, is such that
$V^{(1)},V^{(2)}](z,\beta):=-DV^{(1)}V^{(2)}(z,\beta)=-\begin{pmatrix}-\frac{\beta}{\sigma}&-\frac{z}{\sigma}\\[6.0pt] 0&0\end{pmatrix}\begin{pmatrix}0\\[6.0pt] \sigma_{\beta}\sqrt{1-\rho^{2}}\end{pmatrix}=\begin{pmatrix}\dfrac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{\sigma}z\\[6.0pt] 0\end{pmatrix}$.
Since $z>0$, $\sigma_{\beta}>0$, and $|\rho|<1$, the vectors $V^{(2)}(z,\beta)$ and $V^{(1)},V^{(2)}](z,\beta)$ are linearly independent and therefore span $\mathcal{O}$.
Now consider the case $|\rho|=1$. The Lie bracket between $V^{(0)}$ and $V^{(1)}$ is
$V^{(0)},V^{(1)}](z,\beta):=DV^{(1)}V^{(0)}(z,\beta)-DV^{(0)}V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{z\kappa(\overline{\beta}-\beta)}{\sigma}\\[6.0pt] \dfrac{\rho\sigma_{\beta}}{\sigma}\big(\kappa\sigma+\rho\sigma_{\beta}\big)\end{pmatrix}$.
Since $\det(V^{(1)},[V^{(0)},V^{(1)}])$ is not necessarily non-zero, we proceed to the next bracket. Therefore, we compute
$V^{(1)},[V^{(0)},V^{(1)}]](z,\beta)=D[V^{(0)},V^{(1)}]V^{(1)}(z,\beta)-DV^{(1)}[V^{(0)},V^{(1)}]=\begin{pmatrix}\dfrac{\rho\sigma_{\beta}z}{\sigma^{2}}\big(2\kappa\sigma+\rho\sigma_{\beta}\big)\\[6.0pt] 0\end{pmatrix}$,
which implies
$\det(V^{(1)},[V^{(1)},[V^{(0)},V^{(1)}]])=-\dfrac{\rho^{2}\sigma_{\beta}^{2}z}{\sigma^{2}}\big(2\kappa\sigma+\rho\sigma_{\beta}\big)<0,$
where the last inequality is due to
$2\kappa\sigma+\rho\sigma_{\beta}>(2+\rho)\sigma_{\beta}>0,$
where Assumption 2.1 has been used. Hence, $V^{(1)}$ and $V^{(1)},[V^{(0)},V^{(1)}]]$ are linearly independent and thus span $\mathcal{O}$. Hörmander’s condition is therefore verified and we conclude that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is indeed strong Feller. ∎
Proposition 3.13.
$v\in C^{\infty}(\mathcal{W})$ and it solves in the classical sense
$\mathcal{L}v-rv+\tilde{u}^{\prime}(z)+\ell=0\quad\text{on}\quad\mathcal{W},$
where the second-order differential operator $\mathcal{L}$ is defined as in (3.36).
Proof.
Since the infinitesimal generator of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is hypoelliptic (satisfying Hörmander’s conditions; cf. the proof of Proposition 3.12), the drift and volatilities of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ belong to $C^{\infty}(\mathcal{W})$, and $\tilde{u}^{\prime}(z)+\ell\in C^{\infty}(\mathcal{W})$, Corollary 7 in [36] implies that $v$ is not just a solution to (3.35) in the sense of Schwartz distributions but $v\in C^{\infty}(\mathcal{W})$ and thus solves (3.35) in the classical sense. ∎
The next proposition states that the value function $v$ is not only locally Lipschitz continuous, but actually continuously differentiable. Its proof is based on an application of [23], upon noticing that the Hörmander’s condition, verified in the proof of Lemma 3.12, gives existence of a smooth transition density for the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$.
Proposition 3.14.
One has $v\in C^{1}(\mathcal{O})$.
Proof.
An application of strong Markov property allows to write
$v(z,\beta)=g(z,\beta)-f(z,\beta),\quad(z,\beta)\in\mathcal{O},$
where, for any $(z,\beta)\in\mathcal{O}$, we have set
(3.37) $g(z,\beta):=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)dt\bigg],$
and
(3.38) $f(z,\beta):=\sup_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\Big[e^{-r\tau}g(\hat{Z}_{\tau},\hat{\beta}_{\tau})\Big].$
Hence, the $C^{1}$-property of $v$ reduces to check that for $g$ and $f$.
By the proof of Lemma 3.12, we know that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the Hörmander’s condition. Given that (modulo a change of measure to remove the quadratic term in $\hat{\beta}$ appearing in the drift of the dynamics for $\hat{Z}$) the coefficients of the evolution of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ are linear and infinitely many times differentiable (see (3.16) and (3.17)), it thus follows from Theorem 9-(iii) and Remark 11 in [3] or Theorem 2.3.3 in [34], among others, that, for any $t>0$, $(\hat{Z}_{t},\beta_{t})_{t}$ admits a transition density that is infinitely many times differentiable in its arguments.
An application of the Dominated Convergence Theorem then shows that $g\in C^{1}(\mathcal{O})$. It thus remains to check the continuous differentiability of $f$. With reference to the notation in [23], for any $t\geq 0$, we set $\xi_{t}:=(\hat{Z}_{t},\hat{\beta}_{t})$,
$X_{t}:=e^{-rt}g(\xi_{t})=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{t}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\hat{\mathcal{F}}_{t}\bigg],$
and we can write $X_{t}=M_{t}+A_{t}$, where
$M_{t}:=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\hat{\mathcal{F}}_{t}\bigg],\quad\text{and}\quad A_{t}:=-\int_{0}^{t}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds.$
Notice that, since $\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}|\tilde{u}^{\prime}(\hat{Z}_{s})+\ell|\;ds\bigg]<\infty$ by Proposition 3.3, $(M_{t})_{t}$ is a uniformly integrable $\hat{\mathbb{F}}-$martingale and $dA_{t}=dA^{+}_{t}+dA^{-}_{t}$ with
$dA^{-}_{t}:=-e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)^{+}dt\quad\text{and}\quad dA^{+}_{t}:=e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)^{-}dt,$
which are clearly absolutely continuous with respect to the Lebesgue measure $dm_{2}:=dt$. Moreover, the set $\partial\mathcal{D}$ in [23] reads in our case $\{(z,\beta):\,z=z^{*}(\beta)\}$, which has zero measure with respect to $dm_{1}=dzd\beta$. Finally, as already noted, the process $(\xi_{t})_{t}=(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ admits a density with respect to $dm_{1}$ having spatial derivatives which are (locally) uniformly continuous in $\mathcal{O}\times[t_{0},t_{1}]$, for any $0<t_{0}<t_{1}<\infty$. Hence, Corollary 7 in [23] holds, $f\in C^{1}(\mathcal{O})$, and the proof is complete. ∎
An immediate consequence of Propositions 3.13, 3.14, (3.36), and the fact that $v=0$ in the interior of $\mathcal{S}$, denoted by $\mathring{\mathcal{S}}$, is the following Corollary.
Corollary 3.15.
One has $v\in C^{1}(\mathcal{O})\cap C^{\infty}(\mathcal{W}\cup\mathring{\mathcal{S}})$. Furthermore, $\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}z^{2}v_{zz}+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}-\frac{\rho\sigma_{\beta}}{\sigma}\,\beta zv_{z\beta}$ admits a continuous extension to $\overline{\mathcal{W}}$.
Remark 3.16.
The proof of Proposition 3.14 employs the probabilistic approach developed in [23] (and recently used in [16] and [28]) to establish continuous differentiability of the value function $v$ in the optimal stopping problem in the presence of a smooth transition density for the underlying state process.
An alternative approach to $C^{1}$-regularity in optimal stopping was developed in [8], where continuous differentiability of the value function is linked to the probabilistic regularity of points on the stopping boundary. In our setting, however, following the approach of [8] is challenging, as it is not clear how to prove the probabilistic regularity of the free boundary. Indeed, such a property is typically established in the literature when one can show either monotonicity or (local) Lipschitz continuity of the free boundary. However, proving these properties is highly non-trivial in our framework because the stochastic factor $(\hat{\beta}_{t})_{t}$ acts explicitly as stochastic volatility for the state process $(\hat{Z}_{t})_{t}$ (cf. (3.16)). This intrinsic stochastic volatility makes it difficult to exploit (3.33) and (3.34) to establish monotonicity of $z^{*}$ or to derive the uniform bounds required to prove that the free boundary is Lipschitz continuous (cf. [10]).
To conclude this section, we summarize the results obtained so far in the following theorem.
Theorem 3.17.
Recall the optimal stopping problem (cf. (3.15))
$v(z,\beta):=\inf_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)dt\bigg],$
where the procesess $(\hat{Z}_{t},\hat{\beta_{t}})_{t}$ are given by (3.16) and (3.17).
There exists a lower-semicontinuous free boundary
$z^{*}:\mathbb{R}\rightarrow[\ell^{-\gamma},\infty]$
such that the stopping and continuation regions (cf. (3.27)) are given by
$\mathcal{S}=\{(z,\beta)\in\mathcal{O}:z\geq z^{*}(\beta)\},\quad\text{and}\quad\mathcal{W}=\{(z,\beta)\in\mathcal{O}:z<z^{*}(\beta)\},$
and the optimal stopping time (cf. (3.28)) is given by
(3.39) $\tau^{*}(z,\beta)=\inf\{s\geq 0:\hat{Z}_{s}\geq z^{*}(\hat{\beta}_{s})\},\quad\mathbb{Q}_{z,\beta}\text{-a.s.}$
Additionally, $v\in C^{1}(\mathcal{O})\cap C^{\infty}(\mathcal{W}\cup\mathring{\mathcal{S}})$ and the couple $(v,z^{*})$ satisfies (in the classical sense) the free boundary problem:
$\begin{cases}\mathcal{L}v-rv+\tilde{u}^{\prime}(z)+\ell=0,&\text{on }\mathcal{W},\\ v=0,&\text{on }\mathcal{S},\\ v_{z}(z^{*}(\beta),\beta)=0=v_{\beta}(z^{*}(\beta),\beta),&\beta\in\mathbb{R},\end{cases}$
where the infinitesimal generator $\mathcal{L}$ is given by (3.36).
4. Back to the Primal Problem
In the previous section, we characterized the solution to the auxiliary optimal stopping problem given in (3.15), and hence also for the problem (3.13) due to Proposition 3.1. The following result establishes a connection between the singular control problem (3.6) and the auxiliary optimal stopping problem (3.13) using probabilistic arguments as in [2], [5] and [16].
Theorem 4.1.
It holds that
(4.1) $\tilde{V}(z,\beta)=\int_{0}^{z}v(y,\beta)\,dy,\quad(z,\beta)\in\mathcal{O},$
where $\tilde{V}$ and $v$ are given by (3.10) and (3.13) (equivalently, (3.14)). Furthermore, the optimal singular control for (3.10) is given by
(4.2) $D_{t}^{*}=\exp(-\xi_{t}^{*}),\quad t>0,\quad D_{0^{-}}^{*}=1,$
where, for $(z,\beta)\in\mathcal{O}$,
$\xi_{t}^{*}:=\sup\left\{y\geq 0\mid\tau^{*}(ze^{-y},\beta)<t\right\},\quad t>0,\quad\xi_{0^{-}}^{*}=0,$
and $\tau^{*}(z,\beta)$ is the optimal stopping time for (3.13) given in (3.39).
Proof.
Let us define the candidate value function $U(z,\beta):=\int_{0}^{z}v(y,\beta)\,dy$ with $v$ as in (3.13). We need to show $U(z,\beta)=\tilde{V}(z,\beta)$ for all $(z,\beta)\in\mathcal{O}$.
Step 1. In this step, we establish $U(z,\beta)\leq\tilde{V}(z,\beta)$ for all $(z,\beta)\in\mathcal{O}$. To that end, let $D\in\mathcal{D}$ be an arbitrary admissible singular control and we define its left-continuous inverse process $\tau^{D}$ as
(4.3) $\tau^{D}(\alpha):=\inf\{t\geq 0\mid D_{t}<\alpha\},\quad\alpha\in(0,1].$
The process $\tau^{D}:=\{\tau^{D}(\alpha)\mid\alpha\in(0,1]\}$ has nonincreasing, left-continuous sample paths and hence it admits right-limits $\tau^{D}_{+}(\alpha):=\inf\{t\geq 0\mid D_{t}\leq\alpha\}$. The set of points $\alpha\in(0,1]$ at which $\tau^{D}(\alpha)(\omega)\neq\tau^{D}_{+}(\alpha)(\omega)$ is countable for a.e. $\omega\in\Omega$. Since $(D_{t})_{t}$ is right-continuous and $\tau^{D}(\alpha)$ is the first entry time of an open set, it is an $\bar{{\mathcal{F}}}_{t+}$-stopping time for any given and fixed $\alpha\in(0,1]$. However, $(\bar{\mathcal{F}}_{t})_{t}$ is right-continuous, hence $\tau^{D}(\alpha)$ is an $\bar{\mathbb{F}}$-stopping time. For a fixed $z>0$ and $y\in(0,z]$, we consider $\alpha=\frac{y}{z}$ in the following. Upon using sub-optimality of $\tau^{D}(\frac{y}{z})$ for (3.13), we have
(4.4) $v(y,\beta)\leq\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau^{D}(\frac{y}{z})}e^{-rt}M_{t}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dt\bigg].$
Integrating (4.4) with respect to $y$ over $[0,z]$ then yields
$U(z,\beta)=\int_{0}^{z}v(y,\beta)\,dy\leq\int_{0}^{z}\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau^{D}(\frac{y}{z})}e^{-rt}M_{t}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dt\bigg]\,dy.$
By Fubini’s Theorem, the previous inequality is equivalent to
$U(z,\beta)\leq\mathbb{E}_{z,\beta}\bigg[\int_{0}^{z}\bigg(\int_{0}^{\infty}\mathbf{1}_{\{t<\tau^{D}(\frac{y}{z})\}}e^{-rt}M_{t}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dt\bigg)\,dy\bigg].$
Since
$t<\tau^{D}(\frac{y}{z})\iff D_{t}\geq\frac{y}{z}\iff y\leq zD_{t},$
we then have
(4.5) $U(z,\beta)\leq\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}M_{t}\bigg(\int_{0}^{zD_{t}}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dy\bigg)dt\bigg].$
For the inner integral on the right-hand side of (4.5), upon setting $w=\frac{y}{z}Z^{1}_{t}$ and noticing that, by (3.2), $Z^{D}_{t}=Z^{1}_{t}D_{t}$ with $Z_{t}^{1}=e^{\delta t}z\mathcal{H}_{t}$, we have
$\int_{0}^{zD_{t}}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dy=\frac{z}{Z^{1}_{t}}\int_{0}^{Z^{D}_{t}}(\tilde{u}^{\prime}(w)+\ell)\,dw=\frac{z}{Z^{1}_{t}}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big).$
Substituting this back into (4.5) and using the relation $e^{-rt}M_{t}\frac{z}{Z^{1}_{t}}=e^{-\delta t}$ (cf. (2.3) and (3.2)), we obtain
$U(z,\beta)\leq\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D}_{t})+\ell Z^{D}_{t}\big)\,dt\bigg].$
Since $D\in\mathcal{D}$ was arbitrary, we conclude $U(z,\beta)\leq\tilde{V}(z,\beta)$ for all $(z,\beta)\in\mathcal{O}$.
Step 2. Let $\tau^{*}(z,\beta)$ denote the optimal stopping time for the problem (3.13) with initial data $(z,\beta)$ which is given by (3.39). Since $z\mapsto v(z,\beta)$ is nondecreasing (cf. Proposition 3.4), $z\mapsto\tau^{*}(z,\beta)$ is nonincreasing, and therefore $\eta\mapsto\tau^{*}(ze^{-\eta},\beta)$ is nondecreasing on $(0,\infty)$, for all $(z,\beta)\in\mathcal{O}$. Hence, we can define the process $\xi^{*}$ as the generalized inverse
(4.6) $\xi^{*}_{t}:=\sup\left\{\eta\geq 0\mid\tau^{*}(ze^{-\eta},\beta)<t\right\},\quad t>0,\quad\xi_{0^{-}}^{*}=0.$
We now take $D^{*}_{t}=\exp(-\xi^{*}_{t})$, with $(\xi_{t}^{*})_{t}$ as in (4.6) (cf. (4.2) as well). Optimality of $\tau^{*}(y,\beta)$ for the optimal stopping problem (3.13) then implies
$v(y,\beta)=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau^{*}(y,\beta)}e^{-rt}M_{t}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dt\bigg].$
Since we have from (4.6)
$D_{t}^{*}<\frac{y}{z}\iff\xi_{t}^{*}>\ln(\frac{z}{y})\iff\tau^{*}(ze^{-\ln(\frac{z}{y})},\beta)<t\iff\tau^{*}(y,\beta)<t,$
recalling (4.3) for $D=D^{*}$, we obtain $\tau^{D^{*}}(\frac{y}{z})=\tau^{*}(y,\beta)$ for all $y\in(0,z]$. Hence, integrating over $[0,z]$ yields
$U(z,\beta)=\int_{0}^{z}\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau^{D^{*}}(\frac{y}{z})}e^{-rt}M_{t}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dt\bigg]\,dy.$
As before, an application of Fubini’s Theorem leads to
$U(z,\beta)=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{z}\bigg(\int_{0}^{\infty}\mathbf{1}_{\{t<\tau^{D^{*}}(\frac{y}{z})\}}e^{-rt}M_{t}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dt\bigg)\,dy\bigg].$
Observing that $t<\tau^{D^{*}}(\frac{y}{z})$ is equivalent to $y\leq zD^{*}_{t}$ then yields
(4.7) $U(z,\beta)=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}M_{t}\bigg(\int_{0}^{zD^{*}_{t}}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dy\bigg)\,dt\bigg].$
For the inner integral in (4.7) we use again the change of variables $w:=\frac{y}{z}Z^{1}_{t}$, which implies
$\int_{0}^{zD^{*}_{t}}\big(\tilde{u}^{\prime}(\tfrac{y}{z}Z^{1}_{t})+\ell\big)\,dy$ $=\frac{z}{Z^{1}_{t}}\int_{0}^{Z^{D^{*}}_{t}}\big(\tilde{u}^{\prime}(w)+\ell\big)\,dw$
(4.8) $=\frac{z}{Z^{1}_{t}}\big(\tilde{u}(Z^{D^{*}}_{t})+\ell Z^{D^{*}}_{t}\big),$
and, upon using $e^{-rt}M_{t}\frac{z}{Z^{1}_{t}}=e^{-\delta t}$ (cf. (2.3) and (3.2)), we obtain thanks to (4.7) and (4) that
$U(z,\beta)=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D^{*}}_{t})+\ell Z^{D^{*}}_{t}\big)\,dt\bigg]\geq\tilde{V}(z,\beta),$
where we used that $D^{*}\in\mathcal{D}$ since $(\xi_{t}^{*})_{t}$ is nondecreasing, càdlàg, and $\bar{\mathbb{F}}$-adapted.
Step 3. Combining Step 1 and Step 2, we finally have $\tilde{V}(z,\beta)=U(z,\beta)=\int_{0}^{z}v(y,\beta)dy$ for all $(z,\beta)\in\mathcal{O}$ and that $(D^{*}_{t})_{t}$ as in (4.2) is the optimal singular control. ∎
A direct consequence of Theorem 4.1 and Corollary 3.15 is the following Corollary.
Corollary 4.2.
For $\tilde{V}$ as in (3.10) and $v$ as in (3.13) (equivalently, (3.14)), we have
$\tilde{V}_{z}(z,\beta)=v(z,\beta),\quad(z,\beta)\in\mathcal{O}.$
Consequently, it holds that $\tilde{V}\in C^{1}(\mathcal{O})$ with $\tilde{V}_{z}\in C^{1}(\mathcal{O})\cap C^{\infty}(\mathcal{W}\cup\mathring{\mathcal{S}})$.
The next proposition further characterizes the optimal singular control of problem (3.6).
Proposition 4.3.
Recall (4.2). Then $D^{*}$ admits the representation
(4.9) $D_{t}^{*}=\inf_{0\leq s\leq t}\left(\frac{z^{*}(\beta_{s})}{Z_{s}^{1}}\wedge 1\right),\quad t\geq 0,\quad D_{0^{-}}^{*}=1.$
Proof.
By (4.2), $D^{*}_{0^{-}}=1$ and for any $t\geq 0$ one has
$D_{t}^{*}=\exp(-\xi_{t}^{*}),$
where $\xi_{t}^{*}=\sup\left\{y\geq 0\mid\tau^{*}(ze^{-y},\beta)<t\right\}$ and $\tau^{*}(z,\beta)=\inf\{s\geq 0:\hat{Z}^{z,\beta}_{s}\geq z^{*}(\hat{\beta}^{\beta}_{s})\}$ (cf. (3.39)). We then obtain the following chain of equivalences for $t\geq 0$:
$\tau^{*}(ze^{-y},\beta)<t$ $\iff\exists s\in[0,t]:e^{-y}Z_{s}^{1}\geq z^{*}(\beta_{s})\iff\exists s\in[0,t]:y\leq\ln(\frac{Z_{s}^{1}}{z^{*}(\beta_{s})})$
$\iff\xi_{t}^{*}=\sup_{0\leq s\leq t}\left[\ln\left(\frac{Z_{s}^{1}}{z^{*}(\beta_{s})}\right)\right]^{+}.$
Finally, substituting this into $D_{t}^{*}=\exp(-\xi_{t}^{*})$ and observing that $\exp(-(\ln x)^{+})=\min(1,\frac{1}{x})$, we obtain (4.9).
Clearly, $D^{*}$ as in (4.9) is nonincreasing and càdlàg given the lower-semicontinuity of $z^{*}$ (cf. Lemma 3.10). As a matter of fact, $t\mapsto D_{t}^{*}$ admits left-limits at any point since it is nonincreasing. To show that $D^{*}$ has right-continuous sample paths, we follow the proof of Proposition 5.8 in [9] and first notice that
(4.10) $\liminf_{s\downarrow t}\left(\frac{z^{*}(\beta_{s})}{Z_{s}^{1}}\wedge 1\right)\geq\frac{z^{*}(\beta_{t})}{Z_{t}^{1}}\wedge 1,$
by the lower-semicontinuity of $z^{*}$ (cf. Lemma 3.10) and the continuity of the state processes $(Z_{t}^{1},\beta_{t})_{t}$. Moreover, from (4.10) we obtain
$\lim_{s\downarrow t}D_{s}^{*}$ $=D_{t}^{*}\wedge\lim_{s\downarrow t}\inf_{t<u\leq s}\left(\frac{z^{*}(\beta_{u})}{Z_{u}^{1}}\wedge 1\right)$
(4.11) $=D_{t}^{*}\wedge\liminf_{s\downarrow t}\left(\frac{z^{*}(\beta_{s})}{Z_{s}^{1}}\wedge 1\right)$
$\geq D_{t}^{*}\wedge\left(\frac{z^{*}(\beta_{t})}{Z_{t}^{1}}\wedge 1\right)=D_{t}^{*}.$
Since $\lim_{s\downarrow t}D_{s}^{*}\leq D_{t}^{*}$ by the monotonicity of $t\mapsto D_{t}^{*}$, (4) implies right continuity. ∎
We now have all the necessary ingredients to derive the optimal controls for our primal optimization problem (cf. (3.11))
$V(x,\beta)=\max_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}_{x,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})\,dt\bigg].$
In the following, when needed, we stress the dependence of the unique strong solution to (3.8) on the initial data $(z,\beta)\in\mathcal{O}$ and on $D\in\mathcal{D}$ by writing $Z^{z,\beta,D}$.
Proposition 4.4.
Let $\tilde{V}$ as in (3.10) (see also (4.1)). We have
(4.12) $V(x,\beta)=\inf_{z>0}\big(\tilde{V}(z,\beta)+zx\big),\quad(x,\beta)\in\mathcal{O},$
and for all $(x,\beta)\in\mathcal{O}$, there exists $\hat{z}:=\hat{z}(x,\beta)>0$ such that $\tilde{V}_{z}(\hat{z},\beta)=-x$. Furthermore, the optimal primal controls for (3.11) are given by
$c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}\qquad\text{and}\qquad\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})-\frac{\rho\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
where $(D^{*}_{t})_{t}$ is the optimal singular control for problem (3.6); that is
$D_{t}^{*}=\inf_{0\leq s\leq t}\left(\frac{z^{*}(\beta_{s})}{Z_{s}^{1}}\right)\wedge 1,\quad t\geq 0,\quad D_{0^{-}}^{*}=1.$
Finally, the optimal wealth process is such that
$X_{t}^{*}=-\tilde{V}_{z}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
with $(X_{t}^{*})_{t\geq 0}:=(X_{t}^{\pi^{*},c^{*}})_{t\geq 0}$.
Proof.
Step 1. We first establish the existence of $\hat{z}$ such that $\tilde{V}_{z}(\hat{z},\beta)=-x$ for all $x>0$. Notice that the minimization problem $\inf_{z>0}(\tilde{V}(z,\beta)+zx)$ is equivalent to solving
(4.13) $v(z,\beta)=-x,$
upon using the identity $\tilde{V}_{z}=v$ given by Corollary 4.2. Since $v$ is continuous (see Proposition 3.8), nondecreasing in $z$ (see Proposition 3.4), by Proposition 3.7 satisfies
$\lim_{z\to 0}v(z,\beta)=-\infty,$
as well as $v(z,\beta)=0$ for any $z\geq z^{*}(\beta)$ and $z\mapsto v(z,\beta)$ is strictly increasing on $\mathcal{W}$, there exists a unique $0<\hat{z}<z^{*}(\beta)$ such that $\tilde{V}_{z}(\hat{z},\beta)=-x$.
Step 2. Next, we prove that the strong duality relation (4.12) indeed holds. Since we have already shown the weak-duality (cf. (3.5)), namely
(4.14) $V(x,\beta)\leq\inf_{z>0}\bigg(\tilde{V}(z,\beta)+zx\bigg),$
[…]
$\chi(z):=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}^{*}(c_{t}^{*}-\ell)\,dt\bigg],$
with the aim of showing that actually $\chi(z)=-\tilde{V}_{z}(z,\beta)$ for all $(z,\beta)\in\mathcal{O}$. To that end, we fix $(z,\beta)\in\mathcal{O}$ and $\varepsilon>0$ (small enough), and note that $D^{*}$ is independent of $\varepsilon$. Since $D^{*}$ is suboptimal for $\tilde{V}(z+\varepsilon,\beta)$, we find
$\tilde{V}(z+\varepsilon,\beta)-\tilde{V}(z,\beta)$ $\leq\mathbb{E}\Big[\int_{0}^{\infty}e^{-\delta t}\Big(\big(\tilde{u}(Z_{t}^{z+\varepsilon,\beta,D^{*}})-\tilde{u}(Z_{t}^{z,\beta,D^{*}})\big)+\ell\big(Z_{t}^{z+\varepsilon,\beta,D^{*}}-Z_{t}^{z,\beta,D^{*}}\big)\Big)dt\Big]$
$=\Big((z+\varepsilon)^{-\frac{1-\gamma}{\gamma}}-z^{-\frac{1-\gamma}{\gamma}}\Big)\mathbb{E}\Big[\int_{0}^{\infty}e^{-\delta t}\tilde{u}\Big(Z_{t}^{1,\beta,D^{*}}\Big)dt\Big]$
$\quad+\varepsilon\,\ell\,\mathbb{E}\Big[\int_{0}^{\infty}e^{-\delta t}Z_{t}^{1,\beta,D^{*}}dt\Big].$
Dividing by $\varepsilon$ and letting $\varepsilon\to 0$ gives
(4.15) $\limsup_{\varepsilon\to 0}\frac{\tilde{V}(z+\varepsilon,\beta)-\tilde{V}(z,\beta)}{\varepsilon}\leq\mathbb{E}\Big[\int_{0}^{\infty}e^{-\delta t}\Big(-(Z_{t}^{z,\beta,D^{*}})^{-\frac{1}{\gamma}}+\ell\Big)Z_{t}^{1,\beta,D^{*}}dt\Big].$
A symmetric argument applied to $\tilde{V}(z,\beta)-\tilde{V}(z-\varepsilon,\beta)$, by using now that $D^{*}$ is suboptimal for the problem starting at $(z-\varepsilon,\beta)$, gives
(4.16) $\liminf_{\varepsilon\to 0}\frac{\tilde{V}(z,\beta)-\tilde{V}(z-\varepsilon,\beta)}{\varepsilon}\geq\mathbb{E}\Big[\int_{0}^{\infty}e^{-\delta t}\Big(-(Z_{t}^{z,\beta,D^{*}})^{-\frac{1}{\gamma}}+\ell\Big)Z_{t}^{1,\beta,D^{*}}dt\Big].$
Upon using that $\tilde{V}_{z}$ exists and is continuous (as $\tilde{V}_{z}=v$), (4.15) and (4.16) imply
$\tilde{V}_{z}(z,\beta)=\mathbb{E}\Big[\int_{0}^{\infty}e^{-\delta t}\Big(-(Z_{t}^{z,\beta,D^{*}})^{-\frac{1}{\gamma}}+\ell\Big)Z_{t}^{1,\beta,D^{*}}dt\Big].$
Since now $(Z_{t}^{z,\beta,D^{*}})^{-\frac{1}{\gamma}}=c_{t}^{*}$ and $e^{-\delta t}Z_{t}^{1,\beta,D^{*}}=\mathcal{H}_{t}D_{t}^{*}$ (cf. (3.2)), we have
$\tilde{V}_{z}(z,\beta)=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}^{*}(c_{t}^{*}-\ell)\,dt\bigg]=\chi(z).$
In particular, for $\hat{z}$ as in Step 1, we have
[…]
The previous findings and (3.2) yield the following chain of equations
$\hat{z}\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}^{*}(c_{t}^{*}-\ell)\,dt\bigg]$ $=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}Z_{t}^{D^{*}}(c_{t}^{*}-\ell)\,dt\bigg]$
$=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(u(c_{t}^{*})-\tilde{u}(Z_{t}^{D^{*}})\big)\,dt\bigg]-\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\ell Z_{t}^{D^{*}}\,dt\bigg],$
and hence, by using (4.14), we obtain
$\tilde{V}(\hat{z},\beta)+\hat{z}x$ $=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t}^{*})\,dt\bigg]\leq\sup_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}_{x,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})\,dt\bigg]$
(4.18) $\leq\inf_{z>0}\big(\tilde{V}(z,\beta)+zx\big)\leq\tilde{V}(\hat{z},\beta)+\hat{z}x.$
This in turn yields the strong duality
$V(x,\beta)=\inf_{z>0}(\tilde{V}(z,\beta)+zx)=\tilde{V}(\hat{z},\beta)+\hat{z}x.$
Step 3. In this step, we derive the optimal primal controls associated to the stochastic control problem (3.11). It follows from (4) that
$V(x,\beta)=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t}^{*})\,dt\bigg],$
so that $c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}$ is optimal. By Proposition 2.6, there exists an investment strategy $\pi^{*}$ such that $(\pi^{*},c^{*})\in\mathcal{A}(x)$. Moreover, by the strong Markov property and $x=-\tilde{V}_{z}(\hat{z},\beta)$, we have
$X_{t}^{*}=-\tilde{V}_{z}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}).$
Recalling the regularity of $\tilde{V}_{z}=v$ as in Corollary 3.15, noticing that $\mathbb{P}((Z_{t}^{\hat{z},\beta,\mathcal{D}^{*}},\beta_{t})\in\mathcal{W})=1$ for all $t\geq 0$-, and that the support of the random measure $dD^{*}_{\cdot}$ is $\{t\geq 0:\,\tilde{V}_{zz}(z^{*}(\beta_{t}),\beta_{t})=0\}$ (due to $\tilde{V}_{zz}=v_{z}$ and Proposition 3.14), an application of Itô-Meyer’s formula leads to
(4.19) $dX_{t}^{*}=$ $-\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[(\delta-r)Z_{t}^{\hat{z},\beta,D^{*}}\,dt-\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\,dW_{t}\right]$
$-\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta}\right]$
$-\frac{1}{2}\frac{\beta_{t}^{2}}{\sigma^{2}}(Z_{t}^{\hat{z},\beta,D^{*}})^{2}\tilde{V}_{zzz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\;dt$
$-\frac{1}{2}\sigma_{\beta}^{2}\tilde{V}_{z\beta\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt+\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\sigma_{\beta}\rho\tilde{V}_{zz\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt,\quad t\geq 0.$
Recalling that $dW^{\beta}_{t}=\rho dW_{t}+\sqrt{1-\rho^{2}}dW^{2,\perp}_{t}$ due to $Corr(W_{t},W^{\beta}_{t})=\rho t$, comparing the $dW_{t}$ terms in (2.1) and (4.19) finally gives
$\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})-\frac{\rho\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
which completes the proof. ∎
5. Numerical Illustrations
In this section, we provide numerical illustrations of our results. In order to illustrate the singular control mechanism and how $(D^{*}_{t})_{t}$ ensures the nonnegativity of the wealth process, in Figure 1, we simulate the optimal primal state process $(X_{t}^{\pi^{*},c^{*}})_{t}$, the optimal dual state process $(Z^{\hat{z},D^{*}}_{t})_{t}$, and the optimal singular control $(D^{*}_{t})_{t}$. For this simulation, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\kappa=0.25$, $\bar{\beta}=0.05$, $\sigma_{\beta}=0.03$, $\sigma=0.18$, and $\rho=0.6$.
Figure 1(a) depicts a path of the controlled dual state process $(Z^{\hat{z},D^{*}}_{t})_{t}$ (black line) and the (moving stochastic) free boundary $(z^{*}(\beta_{t}))_{t}$ (red line). Observe that each time the dual state hits the free boundary, it is reflected and pushed downward. This action is driven by the singular control $(D^{*}_{t})_{t}$. Indeed, from Figure 1(b), we see that each time the free boundary is reached, the decreasing singular control jumps to keep the process below the boundary. Crucially, in Figure 1(c), we see that the points where the dual state touches the free boundary correspond exactly to the wealth process hitting zero. At these points, the wealth is reflected upward, ensuring that the no-borrowing constraint is satisfied.
(a) Simulation of dual state $(Z^{*}_{t})_{t}$.
(b) Simulation of optimal singular control $(D^{*}_{t})_{t}$.
(c) Simulation of optimal wealth $(X^{*}_{t})_{t}$.
Figure 1. Simulation of optimal state processes.
One of the main contributions of this paper is the inclusion of the stochastic factor $(\beta_{t})_{t}$, representing the expected excess return of the risky asset. In contrast to the standard Merton model, where $(\beta_{t})_{t}$ is constant, we model it as a stochastic process. Figure 2 compares the optimal wealth trajectories under these two frameworks.
For this comparison, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\bar{\beta}=0.05$, and $\sigma=0.18$. In the stochastic case, we additionally set $\kappa=0.25$ and $\sigma_{\beta}=0.03$, while in the constant case we set $\kappa=0=\sigma_{\beta}$ and $\beta_{0}=\bar{\beta}$ to ensure that $(\beta_{t})_{t}\equiv\bar{\beta}$. We perform the comparison for three distinct correlation parameters: $\rho\in\{-0.6,0,0.6\}$. To isolate the impact of the stochastic factor $(\beta_{t})_{t}$, we fix the same path for the Brownian motion $(W_{t})_{t}$ driving the risky asset price in both cases. For the agent who assumes a stochastic $(\beta_{t})_{t}$, we simulate $10,000$ paths of the Brownian motion $(W^{\beta}_{t})_{t}$ driving the stochastic factor, and plot the optimal wealth as the average over these paths (blue line). The red line depicts the wealth of the agent who assumes a constant $(\beta_{t})_{t}$. In the following, we denote by Stochastic Agent the agent with stochastic $(\beta_{t})_{t}$, and by Constant Agent the agent with constant $(\beta_{t})_{t}$.
In all scenarios, the Stochastic Agent systematically accumulates greater long-term wealth. This outcome is driven by a universal mechanism: The Stochastic Agent’s strategic asset allocation, namely increasing exposure to the risky asset when expected returns are high (that is, when $\beta_{t}$ is high), generates substantial excess returns. As a result, the Stochastic Agent’s wealth exceeds that of the Constant Agent in the long-run. In contrast, the Constant Agent lacks this flexibility and is confined to a rigid strategy that systematically fails to exploit time-varying risk premia, leading to inferior wealth accumulation.
(a) Case $\rho=-0.6$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
(b) Case $\rho=0$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
(c) Case $\rho=0.6$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
Figure 2. Comparison of optimal wealth trajectories between the Stochastic Agent and the Constant Agent ($\beta_{t}\equiv\bar{\beta}$) for different values of $\rho$. The Brownian path of $(W_{t})_{t}$ is fixed across both cases, while the Stochastic Agent’s wealth is averaged over 10,000 Brownian paths of $(W^{\beta}_{t})_{t}$.
While this long-run dominance is consistent across all regimes, short-run dynamics differ. For nonnegative correlations ($\rho=0$ and $\rho=0.6$), the advantage materializes immediately, and the Stochastic Agent maintains wealth consistently above that of the Constant Agent throughout the simulation. In contrast, the regime with $\rho=-0.6$ exhibits a distinct crossing pattern. In this case, the negative correlation provides a natural hedge, as market downturns signal improved future investment opportunities. This added safety reduces the agent’s precautionary savings motive and induces a positive income effect, effectively making the agent feel wealthier and optimally consume more in the present. The resulting higher initial consumption causes the Stochastic Agent to temporarily lag behind the Constant Agent. However, as the benefits of flexible investment decisions accumulate, the wealth paths eventually cross, confirming that the long-run advantage of adapting to $(\beta_{t})_{t}$ ultimately dominates.
Next, we illustrate the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ given by Proposition 4.4 for different parameter choices. To visualize the controls effectively, we plot the optimal strategies as functions of wealth $X_{t}$, while fixing $\beta_{t}$ at its equilibrium level $\bar{\beta}$. It is important to note that, while the stochastic factor is held constant in these figures, the policy functions and underlying value functions are derived from the full dynamic model. Thus, the agent explicitly accounts for the stochastic evolution of $(\beta_{t})_{t}$ throughout the optimization, and the plots represent a cross-section of this dynamic strategy.
In Figure 3, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of labor income $\ell\in\{0.2,0.6,1\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
We observe that both consumption and risky investment are increasing in wealth $X_{t}$ and labor income $\ell$. This is consistent with standard economic intuition: higher levels of wealth and labor income increase the agent’s total effective wealth, thereby relaxing the budget constraint. Consequently, the agent increases both consumption and their allocation to the risky asset.
(a) Optimal investment policy $\pi^{*}_{t}$ for different values of labor income $\ell$.
(b) Optimal consumption policy $c^{*}_{t}$ for different values of labor income $\ell$.
Figure 3. Optimal policies for different values of labor income $\ell$.
In Figure 4, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of risk aversion $\gamma\in\{1.2,1.5,2\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
As before, we observe that both optimal consumption $(c^{*}_{t})_{t}$ and risky investment $(\pi^{*}_{t})_{t}$ are strictly increasing in wealth $X_{t}$ since the agent has more money to allocate. Regarding the effect of risk preferences, the risky investment $(\pi^{*}_{t})_{t}$ is decreasing in the risk aversion parameter $\gamma$. This inverse relationship is consistent with Merton’s classic results, reflecting that more risk-averse agents fear market volatility more and consequently reduce their exposure to the risky asset.
Conversely, the consumption policy $(c^{*}_{t})_{t}$ roughly shifts upward as $\gamma$ increases. This behavior is driven by the Intertemporal Elasticity of Substitution (IES). Agents with lower risk aversion (higher IES) are more willing to postpone current consumption to capitalize on investment opportunities, whereas highly risk-averse agents (lower IES) invest less and choose to consume larger amounts of their current wealth.
(a) Optimal investment policy $\pi^{*}_{t}$ for different values of risk aversion $\gamma$.
(b) Optimal consumption policy $c^{*}_{t}$ for different values of risk aversion $\gamma$.
Figure 4. Optimal policies for different values of risk aversion $\gamma$.
In Figure 5, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of the correlation parameter $\rho\in\{-0.6,0,0.6\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
We observe that the consumption policy $(c_{t}^{*})_{t}$ is increasing in $\rho$. In particular, the agent with positive correlation ($\rho=0.6$) consumes the most, while the agent with negative correlation ($\rho=-0.6$) consumes the least. This pattern reflects the precautionary savings motive. When $\rho<0$, the risky asset acts as an intertemporal hedge, encouraging the agent to take a highly leveraged position. This leverage increases portfolio volatility, and to guard against risk and avoid breaching the borrowing constraint, the agent reduces current consumption. By contrast, when $\rho>0$, the asset offers weaker hedging benefits, leading to lower risky exposure, reduced volatility, and a weaker precautionary savings motive, allowing the agent to consume more.
The investment policy $(\pi_{t}^{*})_{t}$ is non-monotonic with respect to $\rho$. It rises with $\rho$ up to a wealth level of roughly $X_{t}\approx 6.5$, and then declines for higher wealth. In the high-wealth region, the policy follows standard Merton intuition: the negatively correlated agent ($\rho=-0.6$) invests more aggressively to exploit the asset’s hedging properties, while investment declines with $\rho$. At lower wealth levels ($X_{t}<6.5$), however, the pattern flips. The hedging agent’s highly leveraged strategy becomes too risky because of the threat of immediate bankruptcy, forcing sharp deleveraging. Meanwhile, the positively correlated agent ($\rho=0.6$), who naturally maintains a more conservative position, is less constrained and thus holds a higher allocation in the low-wealth region.
(a) Optimal investment policy $\pi^{*}_{t}$ for different values of correlation $\rho$.
(b) Optimal consumption policy $c^{*}_{t}$ for different values of correlation $\rho$.
Figure 5. Optimal policies for different values of correlation $\rho$.
Continuing with our cross-sectional approach, Figure 6 compares the optimal control policies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of expected excess returns $\beta_{t}\in\{0.02,0.05,0.12\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
We observe that the investment policy $(\pi_{t}^{*})_{t}$ is strictly increasing in the expected excess return $\beta_{t}$. When the risk premium is high, the agent aggressively leverages the portfolio to capitalize on favorable investment opportunities. Conversely, when the premium is low, the agent substantially reduces exposure to the risky asset.
The consumption policy $(c_{t}^{*})_{t}$ displays a non-monotonic relationship with $\beta_{t}$. At low wealth levels ($X_{t}<3$), consumption decreases as $\beta_{t}$ increases. In this region, the agent seeks to accumulate wealth more rapidly. Since a higher $\beta_{t}$ implies more profitable investment opportunities, the agent optimally cuts current consumption in order to finance larger risky positions. At higher wealth levels, however, the agent has sufficient capital to fully exploit the high expected returns. This relaxes the need for aggressive saving and allows consumption to rise, eventually exceeding the level observed in the low-$\beta_{t}$ state.
(a) Cross-sections of the optimal investment policy $\pi^{*}_{t}$ for different fixed realizations of the expected excess return $\beta_{t}$.
(b) Cross-sections of the optimal consumption policy $c^{*}_{t}$ for different fixed realizations of the expected excess return $\beta_{t}$.
Figure 6. Cross-sections of the optimal policies for different fixed realizations of the expected excess return $\beta_{t}$.
6. Conclusions
In this paper, we have studied the infinite-horizon consumption and portfolio problem of an investor subject to a no-borrowing constraint and labor income, in a market where expected excess returns follow a mean-reverting Ornstein-Uhlenbeck process. Using a Lagrange duality approach, we have formulated the dual problem as a two-dimensional singular control problem involving the marginal value of wealth and the stochastic factor. The solution is governed by an auxiliary optimal stopping problem, which features a free boundary separating the continuation and stopping regions. We have provided a detailed probabilistic analysis for this optimal stopping problem and established properties of the free boundary and of the value function. Finally, we have retrieved the solutions to the primal optimization problem via duality and provided numerical illustrations.
Acknowledgements. Financial support by the German Research Foundation (DFG) [RTG
2865/1 - 492988838] is gratefully acknowledged.
Appendix A Technical Proofs
A.1. Proof of Proposition 3.1
Proof.
Since $(M_{t})_{t}$ is a martingale, we can fix a finite time horizon $T\geq 0$ and define the new measure $\mathbb$$\bar${P}} on $(\Omega,\mathcal{F})$ via
[…]
$d\hat{Z}_{t}$ $=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}\,dW^{\mathbb{Q}}_{t}+\hat{Z}_{t}\bigg(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}}\bigg)\,dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
$d\hat{\beta}_{t}$ $=\sigma_{\beta}\,dW^{\beta,\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
We set $\hat{\mathbb{F}}:=(\hat{\mathcal{F}}_{t})_{t}$, where $\hat{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W,\mathbb{Q}}\vee\mathcal{F}_{t}^{W^{\beta},\mathbb{Q}}$, $t\geq 0$. We then define the following value functions (where the infimum is taken over all $\hat{\mathbb{F}}-$stopping times)
$v^{\mathbb{Q}}(z,\beta;T):=\inf_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau\wedge T}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)\,dt\right],$
and
$v^{\mathbb{Q}}(z,\beta):=\inf_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)\,dt\right].$
By using the Dominated Convergence Theorem (whose application is justified by using arguments as those in the proof of the upcoming Proposition 3.3), we obtain
(A.2) $\lim_{T\to\infty}v^{\mathbb{Q}}(z,\beta;T)=v^{\mathbb{Q}}(z,\beta),\quad\lim_{T\to\infty}v(z,\beta;T)=v(z,\beta),$
where we have set
$v(z,\beta;T):=\inf_{\tau}\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau\wedge T}e^{-rt}M_{t}\,(\tilde{u}^{\prime}(Z_{t}^{z,1})+\ell)\,dt\bigg].$
Since now
$\inf_{\tau}\mathbb{E}_{z,\beta}^{\bar{\mathbb{P}}}\bigg[\int_{0}^{\tau\wedge T}e^{-rt}\,(\tilde{u}^{\prime}(Z^{1}_{t})+\ell)\,dt\bigg]=\inf_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\tau\wedge T}e^{-rt}\,(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)\,dt\bigg],$
the equivalence in law of the processes $(Z^{1}_{t},\beta_{t},W_{t}^{\bar{\mathbb{P}}},W_{t}^{\beta,\bar{\mathbb{P}}})_{t}$ under $\bar{\mathbb{P}}$ and $(\hat{Z}_{t},\hat{\beta}_{t},W_{t}^{\mathbb{Q}},W_{t}^{\beta,\mathbb{Q}})_{t}$ under $\mathbb{Q}$ on $[0,T]$, and (A.1) allow us to write
$v^{\mathbb{Q}}(z,\beta)=\lim_{T\to\infty}v^{\mathbb{Q}}(z,\beta;T)=\lim_{T\to\infty}v(z,\beta;T)=v(z,\beta),$
which implies
$v(z,\beta)=\inf_{\tau}\mathbb{E}_{z,\beta}^{\mathbb{Q}}\bigg[\int_{0}^{\tau}e^{-rt}\,(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)\,dt\bigg].$
A.2. Proof of Proposition 3.3
Proof.
It follows from (3.3) that $\tilde{u}^{\prime}(\hat{Z}_{t})=-\hat{Z}^{-\frac{1}{\gamma}}_{t}$. Therefore,
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}|-\hat{Z}^{-\frac{1}{\gamma}}_{t}+\ell|dt\bigg]\leq\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]+\ell\int_{0}^{\infty}e^{-rt}dt.$
Hence, in order to prove that $v(z,\beta)\in\mathbb{R}$ it suffices to show that
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]<\infty.$
By Fubini-Tonelli Theorem, we may write
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]=\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]\,dt,$
and we recall that
$W_{t}^{\mathbb{Q}}=\rho\,W_{t}^{\beta,\mathbb{Q}}+\sqrt{1-\rho^{2}}\,\widehat{W}^{1,\perp}_{t},\quad t\geq 0,$
due to the correlation between $W^{\mathbb{Q}}$ and $W^{\beta,\mathbb{Q}}$, where $\widehat{W}^{1,\perp}$ is a Brownian motion independent of $W^{\beta,\mathbb{Q}}$. By using the explicit representation of the strong solution to (3.16), we find
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds+\frac{1}{\gamma}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\bigg)\bigg]$
[…]
$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\mathbb{E}^{\mathbb{Q}}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds$
$\quad+\frac{1}{\gamma}\rho\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW^{\beta,\mathbb{Q}}_{s}+\frac{1}{\gamma}\sqrt{1-\rho^{2}}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}d\widehat{W}^{1,\perp}_{s}\bigg)\bigg|\mathcal{F}_{t}^{W^{\beta},\mathbb{Q}}\bigg]\bigg]$
$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds+\frac{1}{\gamma}\rho\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW^{\beta,\mathbb{Q}}_{s}+\tfrac{1}{2}\frac{1}{\gamma^{2}}(1-\rho^{2})\int_{0}^{t}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}}ds\bigg)\bigg]$
[…]
$=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)$
$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\bigg(\tfrac{\frac{1}{\gamma^{2}}(1-\rho^{2})-\frac{1}{\gamma}}{2\sigma^{2}}\bigg)\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg]$
$=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)$
$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\tfrac{1-\gamma}{2\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\exp\bigg(\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg].$
The process
$N_{t}:=\exp\bigg(\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)$
defines a martingale under $\mathbb{Q}$ by Assumption 3.2. Hence, Girsanov’s Theorem allows us to define a new probability measure by
$\frac{d\mathbb{Q^{\prime}}}{d\mathbb{Q}}\bigg|_{\hat{\mathcal{F}}_{t}}=N_{t},$
and to obtain, as $\gamma>1$,
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]$ $=z^{-\frac{1}{\gamma}}\exp\big({-\frac{1}{\gamma}}(\delta-r)t\big)\mathbb{E}^{\mathbb{Q^{\prime}}}_{1,\beta}\bigg[\exp\bigg(\bigg(\tfrac{1-\gamma}{2\sigma^{2}\gamma^{2}}\bigg)\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg]$
$\leq z^{-\frac{1}{\gamma}}\exp\bigg(-\tfrac{1}{\gamma}(\delta-r)t\bigg),$
which proves (3.18). Overall,
(A.3) $\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{{-\frac{1}{\gamma}}}_{t}dt\bigg]=\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]dt\leq z^{-\frac{1}{\gamma}}\int_{0}^{\infty}e^{-(r+\frac{1}{\gamma}(\delta-r))t}\,dt<\infty,$
since $r+\frac{1}{\gamma}(\delta-r)>0$ as $\gamma>1$. ∎
A.3. Proof of Proposition 3.6
Proof.
Taking the suboptimal stopping time $\tau=0$ in (3.15) clearly yields $v(z,\beta)\leq 0$ for all $(z,\beta)$ on $\mathcal{O}$. As for the lower bound, using that $\tilde{u}^{\prime}(\hat{Z}_{t})=-\hat{Z}_{t}^{-\frac{1}{\gamma}}$ by (3.3), we have for any $(z,\beta)\in\mathcal{O}$ and any stopping time $\tau$
$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)dt\bigg]$ $\geq-\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]$
$\geq-\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]>-\infty,$
where the last inequality is due (3.18) and the fact that $r+\frac{1}{\gamma}(\delta-r)>0$ as $\gamma>1$. Arbitrariness of $\tau\geq 0$ then implies the result. ∎
A.4. Proof of Proposition 3.7
Proof.
We have
(A.4) $0\geq\lim_{z\to\infty}v(z,\beta)\geq\lim_{z\to\infty}\left(-z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]\right)=0,$
where we have used that the expectation in the right-hand side of (A.4) does not depend on $z$. Similarly, one also obtains
$\lim_{z\to 0}v(z,\beta)$ $\leq\lim_{z\to 0}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}(-\hat{Z}_{t}^{-\frac{1}{\gamma}}+\ell)dt\bigg]$
$=\lim_{z\to 0}\left(-z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}_{t}^{-\frac{1}{\gamma}}dt\bigg]+\frac{\ell}{r}\right)=-\infty.$
A.5. Proof of Proposition 3.11
Proof.
The proof is divided into two steps and borrows arguments from [10].
Step 1. We first show that the value function $v$ is (locally) Lipschitz continuous in the $z$-variable and derive the probabilistic representation for the weak derivative $v_{z}$. To that end, we fix $(z,\beta)\in\mathcal{O}$ and $\varepsilon>0$, and let $\tau^{*}$ be the optimal stopping time for the problem with initial data $(z,\beta)$ (independent of $\varepsilon$).
For the purpose of showing the Lipschitz property, we may restrict to $\varepsilon\leq\varepsilon_{0}$ with $\varepsilon_{0}\in(0,1)$ such that $z-\varepsilon>0$. Using $\tilde{u}^{\prime}(\hat{Z}_{t})=-\hat{Z}_{t}^{-\frac{1}{\gamma}}$ and the Mean-Value Theorem, we have
$|v(z+\varepsilon,\beta)-v(z,\beta)|$ $\leq\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}|z^{-\frac{1}{\gamma}}-(z+\varepsilon)^{-\frac{1}{\gamma}}|dt\Big]$
(A.5) $=|z^{-\frac{1}{\gamma}}-(z+\varepsilon)^{-\frac{1}{\gamma}}|\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big]$
$=\frac{1}{\gamma}\xi^{-\frac{1}{\gamma}-1}\,\varepsilon\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big],$
for some $\zeta\in[z,z+\varepsilon]$. Upon setting
$c(z,\beta):=\frac{1}{\gamma}z^{-\frac{1}{\gamma}-1}\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big],$
we thus obtain from (A.5)
$|v(z+\varepsilon,\beta)-v(z,\beta)|\leq c(z,\beta)\,\varepsilon,$
and by symmetry,
$|v(z,\beta)-v(z-\varepsilon,\beta)|\leq c(z,\beta)\,\varepsilon.$
Since $c(z,\beta)>0$ is finite by Proposition 3.3 and can be taken uniformly on compact sets, we conclude that $v(\cdot,\beta)$ is (locally) Lipschitz continuous in $z$.
In order to derive the probabilistic representation of the weak derivative $v_{z}$, we note that $\tau^{*}$ is suboptimal for $v(z+\varepsilon,\beta)$, and thus we have
$v(z+\varepsilon,\beta)-v(z,\beta)$ $\leq\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\tau^{*}}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}\big(z^{-\frac{1}{\gamma}}-(z+\varepsilon)^{-\frac{1}{\gamma}}\big)dt\Big]$
$=\big(z^{-\frac{1}{\gamma}}-(z+\varepsilon)^{-\frac{1}{\gamma}}\big)\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\tau^{*}}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big].$
Dividing by $\varepsilon$ and then letting $\varepsilon\to 0$ yields
(A.6) $\limsup_{\varepsilon\to 0}\frac{v(z+\varepsilon,\beta)-v(z,\beta)}{\varepsilon}\leq\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\tau^{*}}e^{-rt}\frac{1}{\gamma}z^{-1}(\hat{Z}^{z,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big].$
A symmetric argument applied to $v(z,\beta)-v(z-\varepsilon,\beta)$ gives the reverse inequality
(A.7) $\liminf_{\varepsilon\to 0}\frac{v(z,\beta)-v(z-\varepsilon,\beta)}{\varepsilon}\geq\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\tau^{*}}e^{-rt}\frac{1}{\gamma}z^{-1}(\hat{Z}^{z,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big].$
Hence, since $v$ is (locally) Lipschitz continuous in $z$, it follows from (A.6) and (A.7) that for any point $z$ of differentiability (belonging to a set of full measure) the weak derivative $v_{z}$ is given by
$v_{z}(z,\beta)=\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\tau^{*}}e^{-rt}\frac{1}{\gamma}z^{-1}(\hat{Z}^{z,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big].$
Step 2. In this step, we show that $v$ is (locally) Lipschitz continuous in the $\beta$-variable and derive the probabilistic representation for the weak derivative $v_{\beta}$. Again, let $(z,\beta)\in\mathcal{O}$, and to simplify notation, we set
[…]
so that the unique strong solution to (3.17) can be written as
(A.8) $\hat{\beta}^{\beta}_{t}=\beta e^{-at}+b(1-e^{-at})+\sigma_{\beta}\int_{0}^{t}e^{-a(t-s)}dW_{s}^{\beta,\mathbb{Q}},$
and it readily follows that $\frac{\partial\hat{\beta}^{\beta}_{t}}{\partial\beta}=e^{-at}$. Since
$\hat{Z}^{z,\beta}_{t}=z\exp\Bigg(\int_{0}^{t}\big(\delta-r+\frac{1}{2}\frac{(\hat{\beta}^{\beta}_{s})^{2}}{\sigma^{2}}\big)ds-\int_{0}^{t}\frac{\hat{\beta}^{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\Bigg),$
we have by Theorem $V.7.39$ in [38]
(A.9) $\frac{\partial\hat{Z}_{t}}{\partial\beta}=\hat{Z}_{t}\Bigg(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\Bigg),$
upon using that, due to (A.8), one has almost surely
$\partial_{\beta}\Big(\int_{0}^{t}\frac{\hat{\beta}^{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\Big):=\lim_{\varepsilon\to 0}\frac{\int_{0}^{t}\frac{\hat{\beta}_{s}^{\beta+\varepsilon}}{\sigma}dW_{s}^{\mathbb{Q}}-\int_{0}^{t}\frac{\hat{\beta}_{s}^{\beta}}{\sigma}dW_{s}^{\mathbb{Q}}}{\varepsilon}=\int_{0}^{t}\frac{e^{-as}}{\sigma}dW_{s}^{\mathbb{Q}}.$
As in Step 1, we may restrict to $\varepsilon\leq\varepsilon_{0}$ for some $\varepsilon_{0}\in(0,1)$, and by the Mean-Value Theorem and (A.9), we have
$\big|v(z,\beta+\varepsilon)-v(z,\beta)\big|$ $\leq\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}\Big[\big|(\hat{Z}_{t}^{z,\beta})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z,\beta+\varepsilon})^{-\frac{1}{\gamma}}\big|\Big]dt$
$=\frac{\varepsilon}{\gamma}\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}\Big[(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\Big|\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\Big|\Big]dt,$
where $\beta_{\varepsilon}\in(\beta,\beta+\varepsilon)$.
Let
$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\}$
and $q>1$ with $\frac{1}{p}+\frac{1}{q}=1$. Hölder’s inequality and (3.18) then yield
$\big|v(z,\beta+\varepsilon)-v(z,\beta)\big|$ $\leq\frac{\varepsilon}{\gamma}\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}[(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{p}{\gamma}}]^{1/p}\mathbb{E}^{\mathbb{Q}}\Big[\Big|\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\Big|^{q}\Big]^{1/q}dt$
(A.10) $\leq\frac{\varepsilon}{\gamma}z^{-\frac{1}{\gamma}}\int_{0}^{\infty}e^{(-\frac{1}{\gamma}(\delta-r)-r)t}\mathbb{E}^{\mathbb{Q}}\Big[\Big|\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\Big|^{q}\Big]^{1/q}dt.$
We define $\phi:=\frac{1}{\gamma}(\delta-r)+r>0$ and $A^{\varepsilon}_{t}:=\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}$, so that the inequality (A.5) then reads as
(A.11) $\big|v(z,\beta+\varepsilon)-v(z,\beta)\big|\leq\frac{\varepsilon}{\gamma}z^{-\frac{1}{\gamma}}\int_{0}^{\infty}e^{-\phi t}\left\lVert{A^{\varepsilon}_{t}}\right\rVert_{L^{q}}dt,$
where we have set $\left\lVert{\cdot}\right\rVert_{L^{q}}=\mathbb{E}[|\cdot|^{q}]^{\frac{1}{q}}$. We now decompose $A^{\varepsilon}_{t}$ into a deterministic part $G^{\varepsilon}_{t}$ and a stochastic part $B_{t}$. That is,
(A.12) $A^{\varepsilon}_{t}=G^{\varepsilon}_{t}+B_{t},$
with
(A.13) $G^{\varepsilon}_{t}$ $:=\frac{1}{\sigma^{2}}\int_{0}^{t}e^{-as}(\beta_{\varepsilon}e^{-as}+b(1-e^{-as}))ds,$
(A.14) $B_{t}$ $:=\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}e^{-as}\Big(\int_{0}^{s}e^{-a(s-u)}dW_{u}^{\beta,\mathbb{Q}}\Big)ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}.$
A direct calculation yields
$G^{\varepsilon}_{t}=\frac{1}{\sigma^{2}}\bigg(\frac{\beta_{\varepsilon}-b}{2a}(1-e^{-2at})+\frac{b}{a}(1-e^{-at})\bigg),$
which implies
(A.15) $|G^{\varepsilon}_{t}|\leq\frac{1}{\sigma^{2}a}\bigg(\frac{|\beta_{\varepsilon}-b|}{2}+|b|\bigg)\leq\frac{1}{\sigma^{2}a}\left(\frac{|\beta|+1+|b|}{2}+|b|\right)=:C_{G}(\beta),$
with $C_{G}(\beta)$ independent of $\varepsilon>0$. For $B_{t}$, an application of Stochastic Fubini Theorem (cf., e.g., Theorem $65$ in section IV of [38]) leads to
$B_{t}$ $=\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}\left(\int_{u}^{t}e^{-as}e^{-a(s-u)}ds\right)dW_{u}^{\beta,\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
(A.16) $=\frac{\sigma_{\beta}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\beta,\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
$=\frac{\sigma_{\beta}\rho}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\mathbb{Q}}+\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{2,\perp}$
$\quad-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
$=\int_{0}^{t}\underbrace{\left(\frac{\sigma_{\beta}\rho}{2a\sigma^{2}}(e^{-au}-e^{-2at+au})-\frac{1}{\sigma}e^{-au}\right)}_{=:f(u,t)}dW_{u}^{\mathbb{Q}}+\int_{0}^{t}\underbrace{\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}\left(e^{-au}-e^{-2at+au}\right)}_{=:g(u,t)}d\widehat{W}_{u}^{2,\perp},$
where we used $W_{u}^{\beta,\mathbb{Q}}=\rho W_{u}^{\mathbb{Q}}+\sqrt{1-\rho^{2}}\widehat{W}_{u}^{2,\perp}$ due to $Corr(W_{u}^{\mathbb{Q}},W_{u}^{\beta,\mathbb{Q}})=\rho u$. The deterministic functions $f(u,t)$ and $g(u,t)$ in (A.16) can be bounded as follows
$|f(u,t)|\leq e^{-au}\left(\frac{\sigma_{\beta}|\rho|}{2a\sigma^{2}}+\frac{1}{\sigma}\right)=:C_{f}e^{-au},\quad|g(u,t)|\leq\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}e^{-au}=:C_{g}e^{-au},$
and since $W^{\mathbb{Q}}$ is independent of $\widehat{W}^{2,\perp}$, we obtain
(A.17) $\mathrm{Var}(B_{t})$ $=\int_{0}^{t}f(u,t)^{2}du+\int_{0}^{t}g(u,t)^{2}du\leq\frac{C_{f}^{2}+C_{g}^{2}}{2a}=:K.$
Upon noticing that $B_{t}$ is Gaussian with mean 0 and variance $\mathrm{Var}(B_{t})\leq K$, we may write $B_{t}=\sqrt{\mathrm{Var}(B_{t})}Y$, where $Y\sim\mathcal{N}(0,1)$. Therefore, using (A.17), we obtain
(A.18) $\|B_{t}\|_{L^{q}}=\sqrt{\mathrm{Var}(B_{t})}\|Y\|_{L^{q}}\leq\sqrt{K}\|Y\|_{L^{q}}=:C_{q}.$
The fact that $Y\sim\mathcal{N}(0,1)$ implies $\|Y\|_{L^{q}}<\infty$ for all $q\geq 1$. Hence, (A.15) and (A.18) imply
(A.19) $\|A^{\varepsilon}_{t}\|_{L^{q}}\leq\|G^{\varepsilon}_{t}\|_{L^{q}}+\|B_{t}\|_{L^{q}}=|G^{\varepsilon}_{t}|+\|B_{t}\|_{L^{q}}\leq C_{G}(\beta)+C_{q}=:M_{q}(\beta),$
where $M_{q}(\beta)$ is independent of $\varepsilon>0$. Inequalities (A.11) and (A.19) in turn lead to
$\big|v(z,\beta+\varepsilon)-v(z,\beta)\big|\leq\frac{\varepsilon}{\gamma}z^{-\frac{1}{\gamma}}\int_{0}^{\infty}e^{-\phi t}\left\lVert{A^{\varepsilon}_{t}}\right\rVert_{L^{q}}dt$ $\leq\frac{\varepsilon}{\gamma}z^{-\frac{1}{\gamma}}M_{q}(\beta)\int_{0}^{\infty}e^{-\phi t}dt$
$=\frac{M_{q}(\beta)}{\phi\gamma}z^{-\frac{1}{\gamma}}\varepsilon=:c(z,\beta)\varepsilon,$
where we have $0<c(z,\beta)<\infty$. A symmetric argument also yields
$\big|v(z,\beta)-v(z,\beta-\varepsilon)\big|\leq c(z,\beta)\varepsilon.$
Therefore, the value function $v(z,\cdot)$ is (locally) Lipschitz continuous in $\beta$, with a constant that can be taken uniform over compact sets.
Next, we derive the probabilistic representation for the weak derivative $v_{\beta}$, which exists for almost every $\beta\in\mathbb{R}$. Suppose that $\beta$ is a point of differentiability and denote by $\tau^{*}$ the optimal stopping time for the problem with initial data $(z,\beta)$ (independent of $\varepsilon$). Since $\tau^{*}$ is suboptimal for $v(z,\beta+\varepsilon)$, we obtain
$v(z,\beta+\varepsilon)-v(z,\beta)\leq\mathbb{E}^{\mathbb{Q}}\!\left[\int_{0}^{\tau^{*}}e^{-rt}\Big((\hat{Z}_{t}^{z,\beta})^{-\frac{1}{\gamma}}-(\hat{Z}_{t}^{z,\beta+\varepsilon})^{-\frac{1}{\gamma}}\Big)\,dt\right].$
Dividing by $\varepsilon$ and using the Mean Value Theorem yields
(A.20) $\frac{v(z,\beta+\varepsilon)-v(z,\beta)}{\varepsilon}\leq\mathbb{E}^{\mathbb{Q}}\!\left[\int_{0}^{\tau^{*}}e^{-rt}\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}\,ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}\,dW_{s}^{\mathbb{Q}}\right)dt\right],$
where $\beta_{\varepsilon}\in(\beta,\beta+\varepsilon)$. Since
$\mathbb{E}^{\mathbb{Q}}\!\left[\int_{0}^{\infty}e^{-rt}\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left|\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}\,ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}\,dW_{s}^{\mathbb{Q}}\right|dt\right]<\infty,$
which follows from the same arguments used in the proof of (local) Lipschitz continuity of $v$ in the $\beta$-variable, using Fubini’s Theorem, the right-hand side of (A.20) becomes
(A.21) $\int_{0}^{\infty}\mathbb{E}^{\mathbb{Q}}\!\left[e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)\mathbf{1}_{\{t\leq\tau^{*}\}}\right]dt.$
As before, we may restrict to $\varepsilon\leq\varepsilon_{0}$ for some $\varepsilon_{0}\in(0,1)$. Again, exploiting the arguments from the proof of the Lipschitz continuity of $v$ in the $\beta$-variable, we can bound
$\left|\mathbb{E}^{\mathbb{Q}}\!\left[e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)\mathbf{1}_{\{t\leq\tau^{*}\}}\right]\right|$
by a function that is independent of $\varepsilon$ and Lebesgue-integrable over $(0,\infty)$. Therefore, Dominated Convergence Theorem implies
$\lim_{\varepsilon\to 0}\int_{0}^{\infty}\mathbb{E}^{\mathbb{Q}}\!\left[e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)\mathbf{1}_{\{t\leq\tau^{*}\}}\right]dt$
$=\int_{0}^{\infty}\lim_{\varepsilon\to 0}\mathbb{E}^{\mathbb{Q}}\!\left[e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)\mathbf{1}_{\{t\leq\tau^{*}\}}\right]dt.$
Next, we show that the family
(A.22) $\bigg\{\left|(\hat{Z}_{t}^{\beta_{\varepsilon}})^{-1/\gamma}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|\mathbf{1}_{\{t\leq\tau^{*}\}}\bigg\}_{\varepsilon\in(0,\varepsilon_{0})}$
is uniformly integrable. We let
$1<\tilde{p}<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\},$
[…]
$\mathbb{E}^{\mathbb{Q}}\!\left[(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{m}{\gamma}}\left|\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|^{m}\right]$
(A.23) $\qquad\leq\mathbb{E}^{\mathbb{Q}}\!\left[(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{\tilde{p}}{\gamma}}\right]^{\frac{1}{p}}\,\mathbb{E}^{\mathbb{Q}}\!\left[\left|\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right|^{mq}\right]^{\frac{1}{q}}.$
Now, recalling (A.12) and using the arguments used for the derivation of (A.19), we find
(A.24) $\|A_{t}\|_{L^{mq}}\leq M_{mq}(\beta),$
since $mq>1$, which is uniform in $\varepsilon$. Hence, combining (A.5), (A.24) and (3.18) gives
$\sup_{\varepsilon\leq\varepsilon_{0}}\mathbb{E}^{\mathbb{Q}}\!\left[\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|^{m}\mathbf{1}_{\{t\leq\tau^{*}\}}\right]\leq z^{-\frac{s}{\gamma}}\,e^{-\frac{s}{\gamma}(\delta-r)t}M_{mq}(\beta)^{m},$
and thus the uniform integrability of the family (A.22) follows. Vitali’s Convergence Theorem, together with the continuities of $\beta\mapsto\hat{Z}^{z,\beta}$ and $\beta\mapsto\hat{\beta}^{\beta}$, then imply
$\lim_{\varepsilon\to 0}\mathbb{E}^{\mathbb{Q}}\!\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right]$
(A.25) $\quad=\mathbb{E}^{\mathbb{Q}}\!\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right].$
Using (A.5) into (A.20) gives
$\limsup_{\varepsilon\to 0}\frac{v(z,\beta+\varepsilon)-v(z,\beta)}{\varepsilon}$
(A.26) $\leq\mathbb{E}_{z,\beta}^{\mathbb{Q}}\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right],$
and, arguing symmetrically, also
$\liminf_{\varepsilon\to 0}\frac{v(z,\beta)-v(z,\beta-\varepsilon)}{\varepsilon}$
(A.27) $\geq\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right].$
Hence, since $v$ is (locally) Lipschitz continuous in $\beta$, it follows from (A.5) and (A.5) that for almost every $\beta\in\mathbb{R}$ the weak derivative $v_{\beta}$ is given by
(A.28) $v_{\beta}(z,\beta)=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t}^{z,\beta})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right].$

### 11 Mar 2026 v3

@@
-Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model
+Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model: The Complete Market Case
 Abstract.
 In this paper, we study an intertemporal utility maximization problem in which an investor chooses consumption and portfolio strategies in the presence of a stochastic factor and a no-borrowing constraint.
-In the spirit of the Kim–Omberg model, the stochastic factor represents the excess return of the risky asset and follows an Ornstein–Uhlenbeck process, capturing the mean reversion of expected excess returns—a feature well supported by empirical evidence in financial markets.
+In the spirit of the Kim-Omberg model, the stochastic factor represents the expected excess return of the risky asset.
+It is perfectly negatively correlated with shocks to the risky asset, and follows an Ornstein-Uhlenbeck process, thereby capturing the mean reversion of expected excess returns–a feature well supported by empirical evidence in financial markets.
 The investor seeks to maximize expected utility from consumption, subject to the constraint that wealth remains nonnegative at all times.
@@
 Empirical evidence (see, e.g., [15] and [38]) suggests that expected excess returns are predictable and mean-reverting.
-Motivated by this finding, we model $(\beta_{t})_{t}$ as a mean-reverting Ornstein–Uhlenbeck process, following the so-called Kim–Omberg model introduced in [27].
-Furthermore, we allow the Brownian motion driving this stochastic factor to be correlated with the Brownian motion driving the risky asset’s returns.
-We derive optimal consumption and portfolio policies, as well as regularity results for the value function.
+Motivated by this finding, we model $(\beta_{t})_{t}$ as a mean-reverting Ornstein–Uhlenbeck process, following the so-called Kim-Omberg model introduced in [27].
+Furthermore, as in [41], we focus on the case in which the Brownian motion driving $(\beta_{t})_{t}$ is perfectly negatively correlated with the Brownian motion driving the risky asset’s returns, a specification that is empirically motivated by the return-predictability literature (see, e.g., [5], [15] and [38]), where variables proxying time-varying expected excess returns are found to be strongly negatively correlated with contemporaneous stock returns and positively correlated with future returns.
+In such a complete market setting, we derive optimal consumption and portfolio policies, as well as regularity results for the value function.
 As discussed in more detail below, this is achieved by means of a Lagrange duality approach, which allows us to connect the (primal) dynamic optimization problem with a no-borrowing constraint to a singular stochastic control problem.
@@
 Crucially, because the stochastic factor $(\beta_{t})_{t}$ directly drives the diffusion of the dual state, the system inherently features stochastic volatility.
-It is worth noticing that directly approaching the dynamic programming principle (Hamilton-Jacobi-Bellmann, HJB) equation related to the primal problem is particularly challenging in our setting, because of a non-geometric wealth dynamics due to the presence of labor income.
+It is worth noticing that directly approaching the dynamic programming principle (Hamilton-Jacobi-Bellman, HJB) equation related to the primal problem is particularly challenging in our setting, because of a non-geometric wealth dynamics due to the presence of labor income.
 As a matter of fact, in many classical settings, the wealth process is geometric, and the positivity of wealth is automatically guaranteed.
@@
 This critical relationship allows us to uniquely characterize the optimal singular control $(D^{*}_{t})_{t}$, which is the minimal process (à la Skorokhod) that keeps the dual state process $(Z_{t},\beta_{t})_{t}$ within the region $\{(z,\beta):z<z^{*}(\beta)\}$.
-Finally, we prove that strong duality holds, demonstrating that the primal value function can be recovered from the dual value function, and vice versa.
-Using these results, we are then able to retrieve the optimal consumption plan $(c^{*}_{t})_{t}$, the optimal portfolio strategy $(\pi^{*}_{t})_{t}$, and the optimal wealth process $(X^{*}_{t})_{t}$, which, as required, adheres to the dynamic no-borrowing constraint.
+Finally, in our complete market setting, we prove that strong duality holds, demonstrating that the primal value function can be recovered from the dual value function, and vice versa.
+Using these results, we are then able to retrieve the optimal consumption plan $(c^{*}_{t})_{t}$, the optimal portfolio strategy $(\pi^{*}_{t})_{t}$, and the optimal wealth process $(X^{*}_{t})_{t}$, which, as required, satisfies the dynamic no-borrowing constraint.
 Economically, these mathematical results offer a clear interpretation of the system’s dynamics.
@@
 Closely related to our work is the Kim-Omberg model [27], which studies the dynamic non-myopic portfolio behavior of an investor trading a risk-free and a risky asset, with expected excess returns following a mean-reverting Ornstein-Uhlenbeck process.
-Extending this framework to include intermediate consumption, [40] derives exact optimal portfolio and consumption rules.
-However, unlike our model, [40] assumes complete markets by driving both the asset return and the stochastic factor with the same Brownian motion, does not incorporate labor income, and enforces the non-negativity constraint on wealth only at the terminal time $T$.
+Extending this framework to include intermediate consumption, [41] appears to be the first to derive exact optimal portfolio and consumption rules in this setting.
+However, unlike our model, [41] does not incorporate labor income, works with a finite-time horizon and enforces the non-negativity constraint on wealth only at the terminal time $T$.
 Furthermore, [18] derive optimal consumption and investment policies in a complete market featuring a stochastic factor, modeled as a general scalar diffusion that drives investment opportunities.
@@
 The agent can invest in a risk-free asset with a risk-free rate $r>0$ as well as a risky asset $(S_{t})_{t}$ whose dynamics are given by
-$dS_{t}=(r+\beta_{t})S_{t}\,dt+\sigma S_{t}\,dW_{t},\quad t>0,\quad S_{0}=s>0,$
+(2.1) $dS_{t}=(r+\beta_{t})S_{t}\,dt+\sigma S_{t}\,dW_{t},\quad t>0,\quad S_{0}=s>0,$
 where $(W_{t})_{t}$ is a standard Brownian motion on $(\Omega,\mathcal{F},\mathbb{P})$ generating the filtration (completed by -$\mathbb{P}\text{-null sets of}\;\mathcal{F}$) $\mathbb{F}^{W}:=(\mathcal{F}_{t}^{W})_{t}$, and $\sigma>0$ denotes the volatility of the risky asset.
@@
 To capture this mean-reversion, we define $(\beta_{t})_{t}$ as
-$d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
+(2.2) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt-\sigma_{\beta}\,dW_{t},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
 where $\kappa>0$ is the speed of mean reversion, $\overline{\beta}>0$ is the equilibrium level, and $\sigma_{\beta}>0$ is the volatility of $(\beta_{t})_{t}$.
-The standard Brownian motion $(W^{\beta}_{t})_{t}$ on $(\Omega,\mathcal{F},\mathbb{P})$ is correlated with $(W_{t})_{t}$ with correlation parameter $\rho\in[-1,1]$, and generates the filtration (completed by -$\mathbb{P}\text{-null sets of}\;\mathcal{F}$) $\mathbb{F}^{W^{\beta}}:=(\mathcal{F}_{t}^{W^{\beta}})_{t}$.
-Due to correlation, we may write, for $t\geq 0$,
-$W_{t}=\rho W^{\beta}_{t}+\sqrt{1-\rho^{2}}W^{1,\perp}_{t},$
-and
-$W^{\beta}_{t}=\rho W_{t}+\sqrt{1-\rho^{2}}W^{2,\perp}_{t},$
-where the independent standard Brownian motions $(W^{1,\perp}_{t})_{t}$ and $(W^{2,\perp}_{t})_{t}$ are such that $W$ is independent of $W^{2,\perp}$, and $W^{\beta}$ is independent of $W^{1,\perp}$.
-By definition, the excess return $(\beta_{t})_{t}$ reverts to its long-run average $\overline{\beta}$.
+By definition, the expected excess return $(\beta_{t})_{t}$ reverts to its long-run average $\overline{\beta}$.
+As one can see from (2.1) and (2.2), $(S_{t})_{t}$ and $(\beta_{t})_{t}$ are perfectly negatively correlated.
+This complete-market specification is in the spirit of [41], where perfect negative correlation is also assumed in a Kim-Omberg setting, and it plays a fundamental role in the subsequent mathematical analysis.
+Empirically, this assumption is motivated by the evidence that variables proxying time-varying expected excess returns, such as the dividend-price ratio, are strongly negatively correlated with contemporaneous stock returns and positively correlated with future returns (cf.
+[5], [15] and [38]).
 The agent chooses a consumption plan $(c_{t})_{t}$, with $c_{t}\geq 0$, and an investment strategy for the risky asset, denoted by $(\pi_{t})_{t}$.
@@
 (2.4) $d\mathcal{H}_{t}=-r\mathcal{H}_{t}\,dt-\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}\,dW_{t},\quad t>0,\quad\mathcal{H}_{0}=1.$
-We denote by $\bar{\mathbb{F}}:=(\bar{\mathcal{F}}_{t})_{t}$ the completed filtration jointly generated by $W$ and $W^{\beta}$, that is, $\bar{\mathbb{F}}:=\mathbb{F}^{W}\vee\mathbb{F}^{\beta}$, and make the following standing assumption.
+We make the following standing assumption.
 Assumption 2.1.
@@
 (2.5) $M_{t}:=\exp\bigg(-\int_{0}^{t}\frac{1}{2}\frac{\beta_{s}^{2}}{\sigma^{2}}ds-\int_{0}^{t}\frac{\beta_{s}}{\sigma}\,dW_{s}\bigg)=e^{rt}\mathcal{H}_{t}$
-is an $\bar{\mathbb{F}}$- martingale under $\mathbb{P}$, with $(\mathcal{H}_{t})_{t}$ as in (2.2).
+is an $\mathbb{F}^{W}$- martingale under $\mathbb{P}$, with $(\mathcal{H}_{t})_{t}$ as in (2.4).
 The agent faces a no-borrowing constraint; that is, $X_{t}\geq 0$ $\mathbb{P}$-a.s.
@@
   1. (1)
-    $(c_{t})_{t}$ and $(\pi_{t})_{t}$ are $\bar{\mathbb{F}}$-progressively measurable, and are such that $c_{t}\geq 0$ $\mathbb{P}$-a.s.
+    $(c_{t})_{t}$ and $(\pi_{t})_{t}$ are $\mathbb{F}^{W}$-progressively measurable, and are such that $c_{t}\geq 0$ $\mathbb{P}$-a.s.
 for all $t\geq 0$, $\int_{0}^{T}c_{s}\,ds<\infty$ and $\int_{0}^{T}\pi_{s}^{2}\,ds<\infty$ $\mathbb{P}$-a.s.
@@
 To that end, we define
-(2.5) $\mathcal{D}:=\Big\{(D_{t})_{t\geq 0}:D\text{ is }\bar{\mathbb{F}}\text{-adapted, nonnegative, nonincreasing, càdlàg, and }D_{0^{-}}=1\;\mathbb{P}\text{-a.s.}\Big\},$
+(2.7) $\mathcal{D}:=\Big\{(D_{t})_{t\geq 0}:D\text{ is }\mathbb{F}^{W}\text{-adapted, nonnegative, nonincreasing, càdlàg, and }D_{0^{-}}=1\;\mathbb{P}\text{-a.s.}\Big\},$
 and we then have the following result.
@@
   2. (2)
-    For any nonnegative $\bar{\mathbb{F}}$-progressively measurable $(c_{t})_{t}$ with $\int_{0}^{T}c_{s}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$, and such that
+    For any nonnegative $\mathbb{F}^{W}$-progressively measurable $(c_{t})_{t}$ with $\int_{0}^{T}c_{s}\,ds<\infty$ $\mathbb{P}$-a.s. for all $T>0$, and such that
     (2.9) $\sup_{D\in\mathcal{D}}\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right]=x,$
@@
 Proof.
-The proof borrows arguments from [24].
-Proof of (1).
-Let $(c_{t})_{t}$ be a consumption plan such that $(\pi,c)\in\mathcal{A}(x)$ and let $D\in\mathcal{D}$.
-An application of Itô’s formula for semimartingales (see Theorem 32 on p.
-78 in [38]) yields
+Proof of (1). The proof borrows arguments from [24].
+Let $(c_{t})_{t}$ be a consumption plan such that $(\pi,c)\in\mathcal{A}(x)$ and let $D\in\mathcal{D}$. An application of Itô’s formula for semimartingales (see Theorem 32 on p. 78 in [39]) yields
 (2.10) $d(\mathcal{H}_{t}X_{t}D_{t-})=\mathcal{H}_{t}D_{t}(\ell-c_{t})dt+\mathcal{H}_{t}D_{t}(\sigma\pi_{t}-\frac{\beta_{t}}{\sigma}X_{t})dW_{t}+\mathcal{H}_{t}X_{t}dD_{t-}.$
@@
 (2.12) $\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}|c_{t}-\ell|\,dt\right]\leq x+\frac{2\ell}{r}<\infty.$
-Proof of (2). For $T>0$ arbitrary but fixed, consider now the reflected backward stochastic differential equation (RBSDE):
-(2.11) $\begin{cases}\tilde{X}_{s}=\int_{s}^{T}\mathcal{H}_{t}(c_{t}-\ell)\,dt-\int_{s}^{T}\tilde{Z}_{t}\,dW_{t}+\int_{s}^{T}d\tilde{K}_{t},\quad\forall s\geq 0,\\ \tilde{X}_{t}\geq 0,\quad\int_{0}^{T}\tilde{X}_{t}\,d\tilde{K}_{t}=0.\end{cases}$
+Proof of (2). We set
+(2.13) $A_{t}:=\int_{0}^{t}\mathcal{H}_{s}(c_{s}-\ell)ds,\quad t\geq 0,$
+and have for any $\mathbb{F}^{W}$-stopping time $\eta$
+$|A_{\eta}|\leq\int_{0}^{\infty}\mathcal{H}_{s}|c_{s}-\ell|ds.$
+From (2.12), which holds thanks to (2.9) and can be shown as in the proof of (1), it then follows that the process $(A_{t})_{t}$ is continuous and belongs to Class D [^1] A process $(\zeta_{t})_{t}$ is of Class (D) if the family of random variables $\left\{\zeta_{\tau}:\tau<\infty\;\mathbb{P}\text{-a.s.\ is a stopping time}\right\}$ is uniformly integrable..
+Now fix an almost surely finite $\mathbb{F}^{W}$-stopping time $\eta$ and define
+$D_{t}^{\eta}:=\mathbf{1}_{\{t<\eta\}},\qquad t\geq 0,$
+so that $D^{\eta}\in\mathcal{D}$, which implies by (2.9)
+$\mathbb{E}\left[\int_{0}^{\eta}\mathcal{H}_{t}(c_{t}-\ell)\,dt\right]\leq x,$
+or equivalently
+(2.14) $\mathbb{E}[A_{\eta}]\leq x,$
+upon using (2.13). We then define the Snell envelope (cf., e.g., page 8 in [36]) of $(A_{t})_{t}$ by
+$Y_{t}:=\operatorname*{ess\,sup}_{\eta\geq t}\mathbb{E}[A_{\eta}\mid\mathcal{F}^{W}_{t}],\qquad t\geq 0.$
+Since $(A_{t})_{t}$ is continuous and belongs to Class D, it thus follows that $(Y_{t})_{t\geq 0}$ is a continuous supermartingale belonging to Class D, and it is the smallest supermartingale dominating $(A_{t})_{t}$; that is,
+(2.15) $Y_{t}\geq A_{t},\qquad\text{for all }t\geq 0.$
+Hence, (2.14) implies
+(2.16) $Y_{0}=\sup_{\eta\geq 0}\mathbb{E}[A_{\eta}]\leq x.$
+Next, we define the continuous supermartingale
+$\bar{Y}_{t}:=Y_{t}+(x-Y_{0}),\qquad t\geq 0,$
+which, thanks to (2.15) and (2.16), satisfies
+$\bar{Y}_{0}=x\qquad\text{and}\qquad\bar{Y}_{t}\geq A_{t}\quad\text{for all }t\geq 0.$
+By the Doob-Meyer decomposition (cf., e.g., Theorem 3.1 on page 56 in [36]), there exists a uniformly integrable continuous martingale $(\bar{N}_{t})_{t\geq 0}$ and a continuous predictable increasing process $(K_{t})_{t\geq 0}$ with $K_{0}=0$ such that
+(2.17) $\bar{Y}_{t}=x+\bar{N}_{t}-K_{t},\qquad t\geq 0.$
+Moreover, the martingale representation theorem (cf., e.g., Theorem 4.3.4 on page 53 in [35]) yields a progressively measurable process $(\tilde{Z}_{t})_{t\geq 0}$ such that
+(2.18) $\bar{N}_{t}=\int_{0}^{t}\tilde{Z}_{s}\,dW_{s},\qquad t\geq 0,$
+and
 […]
-$\mathbb{E}\left[\sup_{t\in[0,T]}|\tilde{X}_{t}|^{\nu}\right]<\infty\quad\text{and}\quad\mathbb{E}\left[\left(\int_{0}^{T}|\tilde{Z}_{t}|^{2}\,dt\right)^{\nu/2}\right]<\infty.$
-Note that the driver of the RBSDE (2.11), namely $\mathcal{H}_{t}(c_{t}-\ell)$, does not depend on $\tilde{Z}$ and $\tilde{X}$. Furthermore, due to (2.10), we have for all $T>0$
-$\mathbb{E}\left[\int_{0}^{T}\mathcal{H}_{t}|c_{t}-\ell|\,dt\right]<\infty.$
+We now define
+(2.19) $\widehat{X}_{t}:=\bar{Y}_{t}-A_{t}+K_{t},\qquad t\geq 0.$
+Since $\bar{Y}_{t}\geq A_{t}$ and $K_{t}\geq 0$, we have
+$\widehat{X}_{t}\geq 0,\qquad t\geq 0,$
+and, upon using $K_{0}=A_{0}=0$ and $\bar{Y}_{0}=x$, also $\widehat{X}_{0}=x$. Plugging (2.13) and (2.17) into (2.19) yields
+$\widehat{X}_{t}=x+\bar{N}_{t}-K_{t}-\int_{0}^{t}\mathcal{H}_{s}(c_{s}-\ell)\,ds+K_{t}=x+\int_{0}^{t}\mathcal{H}_{s}(\ell-c_{s})\,ds+\int_{0}^{t}\tilde{Z}_{s}\,dW_{s},$
+where we have used (2.18) in the final display equation above, as well. Therefore,
+(2.20) $d\widehat{X}_{t}=\mathcal{H}_{t}(\ell-c_{t})\,dt+\tilde{Z}_{t}\,dW_{t},\qquad\widehat{X}_{0}=x.$
+Let us now construct the candidate investment policy $(\pi_{t})_{t}$ as follows
+(2.21) $\pi_{t}:=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}\widehat{X}_{t}\right),\qquad t\geq 0,$
+which is clearly $\mathbb{F}^{W}$-progressively measurable.
+In order to show
 […]
-Let then $D\in\mathcal{D}$ and define the sequence of localizing stopping times
-$\vartheta_{n}:=\inf\left\{s\geq 0:\int_{0}^{s}\tilde{Z}_{u}^{2}D_{u}^{2}\,du\geq n\right\},\quad n\in\mathbb{N}.$
-Note that $\vartheta_{n}\to\infty$ $\mathbb{P}$-a.s. as $n\to\infty$. Using Itô’s formula on the interval $0,T\wedge\vartheta_{n}]$ and taking expectations, we obtain
-(2.12) $\tilde{X}_{0}$ $=\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}D_{T\wedge\vartheta_{n}}+\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right]$
-$\quad+\mathbb{E}\left[\int_{0}^{T\wedge\vartheta_{n}}D_{t-}\,d\tilde{K}_{t}-\int_{0}^{T\wedge\vartheta_{n}}\tilde{X}_{t}\,dD_{t-}\right].$
-Since $D$ is decreasing and $\tilde{K}$ is increasing, we have from (2.12)
-$\tilde{X}_{0}\geq\mathbb{E}\left[\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right],$
-and letting $n\to\infty$ and $T\to\infty$, (2.10) allows us to invoke the Dominated Convergence Theorem to obtain
-$\tilde{X}_{0}\geq\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}(c_{t}-\ell)\,dt\right].$
-Since $D\in\mathcal{D}$ was arbitrary, it follows from (2.7) that
-(2.13) $\tilde{X}_{0}\geq x.$
-On the other hand, let
-$\tilde{D}_{s}:=\mathbf{1}_{\{s<\tilde{\eta}\}},\;s\geq 0,\quad\tilde{\eta}:=\inf\left\{s\geq 0:\tilde{X}_{s}\leq 0\right\}\wedge(T\wedge\vartheta_{n}).$
-Then,
-$\tilde{X}_{\tilde{\eta}}\mathbf{1}_{\{\tilde{\eta}<T\wedge\vartheta_{n}\}}=0,\quad\tilde{K}_{t}=0\quad\text{on}\quad 0\leq t<\tilde{\eta}$,
-and
-$\int_{0}^{T\wedge\vartheta_{n}}\tilde{D}_{t-}\,d\tilde{K}_{t}=\int_{0}^{T\wedge\vartheta_{n}}\tilde{X}_{t}\,d\tilde{D}_{t-}=0,$
-so that (2.12) implies
-(2.14) $\tilde{X}_{0}=\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}\tilde{D}_{T\wedge\vartheta_{n}}+\int_{0}^{T\wedge\vartheta_{n}}\mathcal{H}_{t}\tilde{D}_{t}(c_{t}-\ell)\,dt\right].$
-Since $\tilde{X}$ belongs to the Class D, we have $\lim_{n\to\infty}\mathbb{E}\left[\tilde{X}_{T\wedge\vartheta_{n}}\right]=\mathbb{E}\left[\tilde{X}_{T}\right]=0$.
-Hence, using the Dominated Convergence Theorem again to take limits as $n\rightarrow\infty$ and $T\rightarrow\infty$ in (2.14), we obtain
-(2.15) $\tilde{X}_{0}\leq\mathbb{E}\left[\int_{0}^{\infty}\mathcal{H}_{t}\tilde{D}_{t}(c_{t}-\ell)\,dt\right]\leq x,$
-where we used the fact that $\tilde{D}\in\mathcal{D}$. Therefore, combining (2.13) and (2.15), we conclude that $\tilde{X}_{0}=x$.
-From the RBSDE (2.11), we know that $\tilde{X}$ satisfies
-$0\leq\tilde{X}_{s}+\int_{0}^{s}d\tilde{K}_{t}=x+\int_{0}^{s}\mathcal{H}_{t}(\ell-c_{t})\,dt+\int_{0}^{s}\tilde{Z}_{t}\,dW_{t},\quad\forall s\geq 0.$
-If we now choose $\pi$ defined as
-(2.16) $\pi_{t}:=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}X^{\pi,c}_{t}\mathcal{H}_{t}\right),\quad t\geq 0,$
-where $(X^{\pi,c}_{t})_{t}$ is the wealth process controlled through $(\pi,c)$, we obtain for any $s\geq 0$
-$0\leq\tilde{X}_{s}+\int_{0}^{s}d\tilde{K}_{t}=x+\int_{0}^{s}\mathcal{H}_{t}(\ell-c_{t})\,dt+\int_{0}^{s}\mathcal{H}_{t}\left(\sigma\pi_{t}-\frac{\beta_{t}}{\sigma}X^{\pi,c}_{t}\right)\,dW_{t}.$
-We need to check that such a $\pi$ is indeed such that $(\pi,c)\in\mathcal{A}(x)$.
-Observe that the expression on the right-hand side of the last display equation is exactly the discounted wealth $X^{\pi,c}_{s}\mathcal{H}_{s}$.
-Since $\mathcal{H}_{s}$ is strictly positive for all $s\geq 0$, we thus get $X^{\pi,c}_{s}\geq 0$ for all $s\geq 0$.
-Let us now check that
-$\int_{0}^{T}\pi_{t}^{2}\,dt<\infty\quad\mathbb{P}-a.s.\ \text{for all}\;T>0.$
-Since $X_{t}^{\pi,c}=\mathcal{H}_{t}^{-1}(\tilde{X}_{t}+\tilde{K}_{t})$, by substituting this expression into (2.16), we find
-$\pi_{t}=\frac{1}{\sigma\mathcal{H}_{t}}\left(\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}(\tilde{X}_{t}+\tilde{K}_{t})\right).$
-As $(\mathcal{H}_{t})_{t}$, $(\beta_{t})_{t}$, $(\tilde{X}_{t})_{t}$ and $(\tilde{K}_{t})_{t}$ are continuous, these processes are $\mathbb{P}$-a.s.
-locally bounded.
-Furthermore, $\int_{0}^{T}\tilde{Z}_{t}^{2}dt<\infty$ $\mathbb{P}$-a.s.
-for all $T>0$, by definition of the $L^{1}$-solution to (2.11).
-Thus, $\int_{0}^{T}\pi_{t}^{2}dt<\infty$ $\mathbb{P}$-a.s.
-for all $T>0$, which implies $(\pi,c)\in\mathcal{A}(x)$.
-∎
+we note that the processes $(\mathcal{H}_{t})_{t}$, $(\mathcal{H}^{-1}_{t})_{t}$, $(\beta_{t})_{t}$, and $(\widehat{X}_{t})_{t}$ are continuous, and hence pathwise bounded.
+Since we also have $\int_{0}^{T}\tilde{Z}_{s}^{2}\,ds<\infty$ $\mathbb{P}$-a.s.
+for all $T>0$, we obtain (2.22) from (2.21).
+It remains to verify that $(X^{\pi,c}_{t})_{t\geq 0}$ is such that $X^{\pi,c}_{t}\geq 0$ for all $t\geq 0$, where $(X^{\pi,c}_{t})_{t\geq 0}$ denotes the wealth controlled through $(\pi,c)$.
+As the process $(\mathcal{H}_{t})_{t}$ (cf.
+(2.4)) is such that
+$d(\mathcal{H}_{t}^{-1})=\mathcal{H}_{t}^{-1}\left(r+\frac{\beta_{t}^{2}}{\sigma^{2}}\right)dt+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\,dW_{t},$
+applying Itô’s formula to the process $(\mathcal{H}_{t}^{-1}\widehat{X}_{t})_{t}$ yields
+$d(\mathcal{H}_{t}^{-1}\widehat{X}_{t})$ $=\mathcal{H}_{t}^{-1}d\widehat{X}_{t}+\widehat{X}_{t}\,d(\mathcal{H}_{t}^{-1})+d[\widehat{X},\mathcal{H}^{-1}]_{t}$
+$=(\ell-c_{t})\,dt+\mathcal{H}_{t}^{-1}\tilde{Z}_{t}\,dW_{t}+\mathcal{H}_{t}^{-1}\widehat{X}_{t}\left(r+\frac{\beta_{t}^{2}}{\sigma^{2}}\right)dt+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\widehat{X}_{t}\,dW_{t}+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\tilde{Z}_{t}\,dt,$
+where (2.20) has also been used. Rearranging terms gives
+$d(\mathcal{H}_{t}^{-1}\widehat{X}_{t})=\left(r\mathcal{H}_{t}^{-1}\widehat{X}_{t}+\ell-c_{t}+\frac{\beta_{t}}{\sigma}\left(\mathcal{H}_{t}^{-1}\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\widehat{X}_{t}\right)\right)dt+\left(\mathcal{H}_{t}^{-1}\tilde{Z}_{t}+\frac{\beta_{t}}{\sigma}\mathcal{H}_{t}^{-1}\widehat{X}_{t}\right)dW_{t},$
+or equivalently, upon using (2.21),
+$d(\mathcal{H}_{t}^{-1}\widehat{X}_{t})=(r\mathcal{H}_{t}^{-1}\widehat{X}_{t}+\beta_{t}\pi_{t}-c_{t}+\ell)\,dt+\sigma\pi_{t}\,dW_{t},\qquad\mathcal{H}_{0}^{-1}\widehat{X}_{0}=x.$
+[…]
+(2.23) $X_{t}^{\pi,c}=\mathcal{H}_{t}^{-1}\widehat{X}_{t},\qquad t\geq 0,$
+by the uniqueness of the solution to the SDE. Finally, since $\widehat{X}_{t}\geq 0$ and $\mathcal{H}_{t}>0$ for all $t\geq 0$, we have from (2.23)
+$X_{t}^{\pi,c}\geq 0\qquad\text{for all }t\geq 0,$
+and therefore, $(\pi,c)\in\mathcal{A}(x)$. ∎
 Remark 2.5.
@@
 (3.8) $dZ^{D}_{t}=(\delta-r)Z^{D}_{t}\,dt-\frac{\beta_{t}}{\sigma}Z^{D}_{t}\,dW_{t}+Z^{D}_{t}\frac{dD_{t}}{D_{t}},\quad t>0,\quad Z_{0^{-}}^{D}=z>0,$
-(3.9) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
+(3.9) $d\beta_{t}=\kappa(\overline{\beta}-\beta_{t})\,dt-\sigma_{\beta}\,dW_{t},\quad t>0,\quad\beta_{0}=\beta\in\mathbb{R},$
 with $D\in\mathcal{D}$. Notice that the dynamics of $(Z_{t}^{D})_{t}$ is easily obtained from (3.2) via Itô’s formula.
@@
 Such an optimal stopping problem is introduced and studied in the following sections.
+Remark 3.1.
+We emphasize that while the complete-market assumption is needed in order to establish strong duality and recover the solution to the primal problem from the dual formulation, the singular stochastic control problem (3.6) and the associated auxiliary optimal stopping problem introduced and studied below can be formulated and analyzed as in the subsequent sections more generally when the factor process $(\beta_{t})_{t}$ is driven by a Brownian motion $(W_{t}^{\beta})_{t}$ satisfying
+$W_{t}^{\beta}=\rho W_{t}+\sqrt{1-\rho^{2}}\,W_{t}^{\perp},\qquad t\geq 0,$
+where $(W^{\perp}_{t})_{t}$ is a standard Brownian motion independent of $(W_{t})_{t}$ and $\rho\in[-1,1]$.
 3.2. Derivation of the Auxiliary Optimal Stopping Problem
@@
 (3.13) $v(z,\beta):=\inf_{\tau}\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\tau}e^{-rt}M_{t}\,(\tilde{u}^{\prime}(Z_{t}^{1})+\ell)\,dt\bigg],$
-where $\tilde{u}^{\prime}(z)=-z^{1/\gamma}$, and where we take the infimum over $\bar{\mathbb{F}}$-stopping times $\tau\geq 0$, $Z^{1}$ evolves as in (3.12), and $\beta$ as in (3.9).
+where $\tilde{u}^{\prime}(z)=-z^{1/\gamma}$, and where we take the infimum over $\mathbb{F}^{W}$-stopping times $\tau\geq 0$, $Z^{1}$ evolves as in (3.12), and $\beta$ as in (3.9).
 We expect $v$ to be such that $\tilde{V}_{z}=v$ on $\mathcal{O}$.
@@
 $d\hat{Z}_{t}$ $=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}\,dW^{\mathbb{Q}}_{t}+\hat{Z}_{t}\bigg(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}}\bigg)\,dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
-$d\hat{\beta}_{t}$ $=\sigma_{\beta}\,dW^{\beta,\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
+$d\hat{\beta}_{t}$ $=-\sigma_{\beta}\,dW^{\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})+\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
 […]
@@
 (3.16) $d\hat{Z}_{t}$ $=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}dW^{\mathbb{Q}}_{t}+\hat{Z}_{t}(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}})dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
-(3.17) $d\hat{\beta}_{t}$ $=\sigma_{\beta}dW_{t}^{\beta,\mathbb{Q}}+(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta})dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
+(3.17) $d\hat{\beta}_{t}$ $=-\sigma_{\beta}dW_{t}^{\mathbb{Q}}+(\kappa(\overline{\beta}-\hat{\beta}_{t})+\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta})dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
 In the following, when needed, we stress the dependence of the unique strong solution to (3.16)-(3.17) on the initial data $(z,\beta)\in\mathcal{O}$ by writing $(\hat{Z}_{t}^{z,\beta})_{t}$ and $(\hat{\beta}_{t}^{\beta})_{t}$.
@@
 We assume that
-$\gamma>\max\bigg\{1,\frac{|\rho|\sigma_{\beta}}{\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}\bigg\}.$
-Notice that $\kappa+\frac{\rho\sigma_{\beta}}{\sigma}>0$ due to Assumption 2.1. We then have the following first preliminary finding.
+$\gamma>\max\bigg\{1,\frac{\sigma_{\beta}}{\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}\bigg\}.$
+Notice that $\kappa-\frac{\sigma_{\beta}}{\sigma}>0$ due to Assumption 2.1. We then have the following first preliminary finding.
 Proposition 3.4.
@@
 where $p$ is chosen such that
-$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\},$
-upon noticing that, by Assumption 3.2, we have $\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)>|\rho|\sigma_{\beta}$.
+$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}{\sigma_{\beta}}\bigg\},$
+upon noticing that, by Assumption 3.3, we have $\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)>\sigma_{\beta}$.
 Therefore, by Vitali’s Convergence Theorem and continuity of $\beta\mapsto\hat{Z}_{t}^{z,\beta}$, we conclude that
@@
 (3.34) $v_{\beta}(z,\beta)=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau^{*}}e^{-rt}\left(\frac{1}{\gamma}(\hat{Z}_{t})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right)dt\right],$
-where $a:=\kappa+\rho\frac{\sigma_{\beta}}{\sigma}>0$ by Assumption 2.1.
+where $a:=\kappa-\frac{\sigma_{\beta}}{\sigma}>0$ by Assumption 2.1.
 Proof.
@@
 where $\mathcal{L}$ is the infinitesimal generator of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ such that
-(3.36) $(\mathcal{L}v)(z,\beta)=\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}(2zv_{z}+z^{2}v_{zz})+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+(\delta-r)zv_{z}+\kappa(\overline{\beta}-\beta)v_{\beta}-\rho\frac{\beta}{\sigma}\sigma_{\beta}\,(v_{\beta}+zv_{z\beta}),$
+(3.36) $(\mathcal{L}v)(z,\beta)=\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}(2zv_{z}+z^{2}v_{zz})+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+(\delta-r)zv_{z}+\kappa(\overline{\beta}-\beta)v_{\beta}+\frac{\beta}{\sigma}\sigma_{\beta}\,(v_{\beta}+zv_{z\beta}),$
 and the PDE above is intended in the sense of Schwartz distributions (see Corollary 5 in [37]).
@@
 Proof.
-Given that $Corr(W_{t}^{\mathbb{Q}},W_{t}^{\beta,\mathbb{Q}})=\rho t$, $\rho\in[-1,1]$, we can write
-$W_{t}^{\beta,\mathbb{Q}}=\rho\,W_{t}^{\mathbb{Q}}+\sqrt{1-\rho^{2}}\,\widehat{W}_{t}^{2,\perp},\quad t\geq 0,$
-where $\widehat{W}^{2,\perp}$ is a standard Brownian motion independent of $W^{\mathbb{Q}}$. Then (cf. (3.16) and (3.17)),
+Recall the dynamics of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ (cf. (3.16) and (3.17)),
 $d\hat{Z}_{t}=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}\,dW_{t}^{\mathbb{Q}}+\hat{Z}_{t}\bigg(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}}\bigg)dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
 and
-$d\hat{\beta}_{t}=\sigma_{\beta}\rho\,dW_{t}^{\mathbb{Q}}+\sigma_{\beta}\sqrt{1-\rho^{2}}\,d\widehat{W}_{t}^{2,\perp}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
+$d\hat{\beta}_{t}=-\sigma_{\beta}\,dW_{t}^{\mathbb{Q}}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})+\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
 Notice that (3.36) is not uniformly elliptic.
-As a matter of fact, denoting by $\Sigma(z,\beta)$ the diffusion matrix associated to (3.16)-(3.17) one has $\det(\Sigma\Sigma^{T}(z,\beta))=\frac{\beta^{2}z^{2}\sigma_{\beta}^{2}}{\sigma^{2}}(1-\rho^{2})$, which vanishes for $\beta=0$.
+As a matter of fact, denoting by $\Sigma(z,\beta)$ the diffusion matrix associated to (3.16)-(3.17) one has $\det(\Sigma\Sigma^{T}(z,\beta))=0$.
 We therefore now check that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the so-called Hörmander’s condition (see, e.g., condition (H) in Section 2.3.2 in [34]).
 This implies that the second-order infinitesimal generator $\mathcal{L}$ of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ as in (3.36) is hypoelliptic and therefore that $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is a strong Feller process (see Proposition 4 in [14]).
-For $(z,\beta)\in\mathcal{O}$, arbitrary but fixed, we then define the Stratonovich-corrected drift vector field and the diffusion vector fields driven by $W^{\mathbb{Q}}$ and $\widehat{W}^{2,\perp}$ as follows:
-$V^{(0)}(z,\beta)=\begin{pmatrix}\dfrac{z\big(\beta^{2}+\rho\sigma\sigma_{\beta}+2\sigma^{2}(\delta-r)\big)}{2\sigma^{2}}\\[8.0pt] \dfrac{-\beta\rho\sigma_{\beta}+\kappa\sigma(\overline{\beta}-\beta)}{\sigma}\end{pmatrix},\qquad V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{\beta}{\sigma}z\\[4.0pt] \sigma_{\beta}\rho\end{pmatrix}$,
-$V^{(2)}(z,\beta)=\begin{pmatrix}0\\[4.0pt] \sigma_{\beta}\sqrt{1-\rho^{2}}\end{pmatrix}$.
-First, consider the case $|\rho|<1$.
-We denote by $DV^{(i)}$ the Jacobian matrix of the vector $V^{(i)}$, $i=0,1,2$.
-The Lie-bracket between $V^{(1)}$ and $V^{(2)}$, denoted by $V^{(1)},V^{(2)}]$, is such that
-$V^{(1)},V^{(2)}](z,\beta):=-DV^{(1)}V^{(2)}(z,\beta)=-\begin{pmatrix}-\frac{\beta}{\sigma}&-\frac{z}{\sigma}\\[6.0pt] 0&0\end{pmatrix}\begin{pmatrix}0\\[6.0pt] \sigma_{\beta}\sqrt{1-\rho^{2}}\end{pmatrix}=\begin{pmatrix}\dfrac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{\sigma}z\\[6.0pt] 0\end{pmatrix}$.
-Since $z>0$, $\sigma_{\beta}>0$, and $|\rho|<1$, the vectors $V^{(2)}(z,\beta)$ and $V^{(1)},V^{(2)}](z,\beta)$ are linearly independent and therefore span $\mathcal{O}$.
-Now consider the case $|\rho|=1$. The Lie bracket between $V^{(0)}$ and $V^{(1)}$ is
-$V^{(0)},V^{(1)}](z,\beta):=DV^{(1)}V^{(0)}(z,\beta)-DV^{(0)}V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{z\kappa(\overline{\beta}-\beta)}{\sigma}\\[6.0pt] \dfrac{\rho\sigma_{\beta}}{\sigma}\big(\kappa\sigma+\rho\sigma_{\beta}\big)\end{pmatrix}$.
+For $(z,\beta)\in\mathcal{O}$, arbitrary but fixed, we then define the Stratonovich-corrected drift vector field and the diffusion vector field driven by $W^{\mathbb{Q}}$ as follows:
+$V^{(0)}(z,\beta)=\begin{pmatrix}\dfrac{z\big(\beta^{2}-\sigma\sigma_{\beta}+2\sigma^{2}(\delta-r)\big)}{2\sigma^{2}}\\[8.0pt] \dfrac{\beta\sigma_{\beta}+\kappa\sigma(\overline{\beta}-\beta)}{\sigma}\end{pmatrix},\qquad V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{\beta}{\sigma}z\\[4.0pt] -\sigma_{\beta}\end{pmatrix}$.
+We denote by $DV^{(i)}$ the Jacobian matrix of the vector $V^{(i)}$, $i=0,1$. The Lie bracket between $V^{(0)}$ and $V^{(1)}$, denoted by $V^{(0)},V^{(1)}]$, is such that
+$V^{(0)},V^{(1)}](z,\beta):=DV^{(1)}V^{(0)}(z,\beta)-DV^{(0)}V^{(1)}(z,\beta)=\begin{pmatrix}-\dfrac{z\kappa(\overline{\beta}-\beta)}{\sigma}\\[6.0pt] -\dfrac{\sigma_{\beta}}{\sigma}\big(\kappa\sigma-\sigma_{\beta}\big)\end{pmatrix}$.
 Since $\det(V^{(1)},[V^{(0)},V^{(1)}])$ is not necessarily non-zero, we proceed to the next bracket. Therefore, we compute
-$V^{(1)},[V^{(0)},V^{(1)}]](z,\beta)=D[V^{(0)},V^{(1)}]V^{(1)}(z,\beta)-DV^{(1)}[V^{(0)},V^{(1)}]=\begin{pmatrix}\dfrac{\rho\sigma_{\beta}z}{\sigma^{2}}\big(2\kappa\sigma+\rho\sigma_{\beta}\big)\\[6.0pt] 0\end{pmatrix}$,
+$V^{(1)},[V^{(0)},V^{(1)}]](z,\beta)=D[V^{(0)},V^{(1)}]V^{(1)}(z,\beta)-DV^{(1)}[V^{(0)},V^{(1)}]=\begin{pmatrix}\dfrac{-\sigma_{\beta}z}{\sigma^{2}}\big(2\kappa\sigma-\sigma_{\beta}\big)\\[6.0pt] 0\end{pmatrix}$,
 which implies
-$\det(V^{(1)},[V^{(1)},[V^{(0)},V^{(1)}]])=-\dfrac{\rho^{2}\sigma_{\beta}^{2}z}{\sigma^{2}}\big(2\kappa\sigma+\rho\sigma_{\beta}\big)<0,$
+$\det(V^{(1)},[V^{(1)},[V^{(0)},V^{(1)}]])=-\dfrac{\sigma_{\beta}^{2}z}{\sigma^{2}}\big(2\kappa\sigma-\sigma_{\beta}\big)<0,$
 where the last inequality is due to
-$2\kappa\sigma+\rho\sigma_{\beta}>(2+\rho)\sigma_{\beta}>0,$
-where Assumption 2.1 has been used.
+$2\kappa\sigma-\sigma_{\beta}>0,$
+in which Assumption 2.1 has been used.
 Hence, $V^{(1)}$ and $V^{(1)},[V^{(0)},V^{(1)}]]$ are linearly independent and thus span $\mathcal{O}$.
@@
 Since the infinitesimal generator of the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ is hypoelliptic (satisfying Hörmander’s conditions; cf.
-the proof of Proposition 3.12), the drift and volatilities of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ belong to $C^{\infty}(\mathcal{W})$, and $\tilde{u}^{\prime}(z)+\ell\in C^{\infty}(\mathcal{W})$, Corollary 7 in [36] implies that $v$ is not just a solution to (3.35) in the sense of Schwartz distributions but $v\in C^{\infty}(\mathcal{W})$ and thus solves (3.35) in the classical sense.
+the proof of Lemma 3.13), the drift and volatilities of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ belong to $C^{\infty}(\mathcal{W})$, and $\tilde{u}^{\prime}(z)+\ell\in C^{\infty}(\mathcal{W})$, Corollary 7 in [37] implies that $v$ is not just a solution to (3.35) in the sense of Schwartz distributions but $v\in C^{\infty}(\mathcal{W})$ and thus solves (3.35) in the classical sense.
 ∎
@@
 By the proof of Lemma 3.13, we know that the process $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ satisfies the Hörmander’s condition.
-Given that (modulo a change of measure to remove the quadratic term in $\hat{\beta}$ appearing in the drift of the dynamics for $\hat{Z}$) the coefficients of the evolution of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ are linear and infinitely many times differentiable (see (3.16) and (3.17)), it thus follows from Theorem 9-(iii) and Remark 11 in [3] or Theorem 2.3.3 in [34], among others, that, for any $t>0$, $(\hat{Z}_{t},\beta_{t})_{t}$ admits a transition density that is infinitely many times differentiable in its arguments.
+Given that (modulo a change of measure to remove the quadratic term in $\hat{\beta}$ appearing in the drift of the dynamics for $\hat{Z}$) the coefficients of the evolution of $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ are linear and infinitely many times differentiable (see (3.16) and (3.17)), it thus follows from Theorem 9-(iii) and Remark 11 in [3] or Theorem 2.3.3 in [34], among others, that, for any $t>0$, $(\hat{Z}_{t},\hat{\beta}_{t})_{t}$ admits a transition density that is infinitely many times differentiable in its arguments.
 An application of the Dominated Convergence Theorem then shows that $g\in C^{1}(\mathcal{O})$.
@@
 With reference to the notation in [23], for any $t\geq 0$, we set $\xi_{t}:=(\hat{Z}_{t},\hat{\beta}_{t})$,
-$X_{t}:=e^{-rt}g(\xi_{t})=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{t}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\hat{\mathcal{F}}_{t}\bigg],$
+$X_{t}:=e^{-rt}g(\xi_{t})=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{t}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\mathcal{F}_{t}^{W,\mathbb{Q}}\bigg],$
 and we can write $X_{t}=M_{t}+A_{t}$, where
-$M_{t}:=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\hat{\mathcal{F}}_{t}\bigg],\quad\text{and}\quad A_{t}:=-\int_{0}^{t}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds.$
-Notice that, since $\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}|\tilde{u}^{\prime}(\hat{Z}_{s})+\ell|\;ds\bigg]<\infty$ by Proposition 3.3, $(M_{t})_{t}$ is a uniformly integrable $\hat{\mathbb{F}}-$martingale and $dA_{t}=dA^{+}_{t}+dA^{-}_{t}$ with
+$M_{t}:=\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds\,\Big|\,\mathcal{F}_{t}^{W,\mathbb{Q}}\bigg],\quad\text{and}\quad A_{t}:=-\int_{0}^{t}e^{-rs}(\tilde{u}^{\prime}(\hat{Z}_{s})+\ell)ds.$
+Notice that, since $\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rs}|\tilde{u}^{\prime}(\hat{Z}_{s})+\ell|\;ds\bigg]<\infty$ by Proposition 3.4, $(M_{t})_{t}$ is a uniformly integrable $\mathbb{F}^{W,\mathbb{Q}}$-martingale and $dA_{t}=dA^{+}_{t}+dA^{-}_{t}$ with
 $dA^{-}_{t}:=-e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)^{+}dt\quad\text{and}\quad dA^{+}_{t}:=e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)^{-}dt,$
@@
 ∎
-An immediate consequence of Propositions 3.13, 3.14, (3.36), and the fact that $v=0$ in the interior of $\mathcal{S}$, denoted by $\mathring{\mathcal{S}}$, is the following Corollary.
+An immediate consequence of Propositions 3.14, 3.15, (3.36), and the fact that $v=0$ in the interior of $\mathcal{S}$, denoted by $\mathring{\mathcal{S}}$, is the following Corollary.
 Corollary 3.16.
 One has $v\in C^{1}(\mathcal{O})\cap C^{\infty}(\mathcal{W}\cup\mathring{\mathcal{S}})$.
-Furthermore, $\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}z^{2}v_{zz}+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}-\frac{\rho\sigma_{\beta}}{\sigma}\,\beta zv_{z\beta}$ admits a continuous extension to $\overline{\mathcal{W}}$.
+Furthermore, $\frac{1}{2}\frac{\beta^{2}}{\sigma^{2}}z^{2}v_{zz}+\frac{1}{2}\sigma_{\beta}^{2}v_{\beta\beta}+\frac{\sigma_{\beta}}{\sigma}\,\beta zv_{z\beta}$ admits a continuous extension to $\overline{\mathcal{W}}$.
 Remark 3.17.
@@
 $\omega\in\Omega$.
-Since $(D_{t})_{t}$ is right-continuous and $\tau^{D}(\alpha)$ is the first entry time of an open set, it is an $\bar{{\mathcal{F}}}_{t+}$-stopping time for any given and fixed $\alpha\in(0,1]$.
-However, $(\bar{\mathcal{F}}_{t})_{t}$ is right-continuous, hence $\tau^{D}(\alpha)$ is an $\bar{\mathbb{F}}$-stopping time.
+Since $(D_{t})_{t}$ is right-continuous and $\tau^{D}(\alpha)$ is the first entry time of an open set, it is an $\mathcal{F}^{W}_{t+}$-stopping time for any given and fixed $\alpha\in(0,1]$.
+However, $(\mathcal{F}^{W}_{t})_{t}$ is right-continuous, hence $\tau^{D}(\alpha)$ is an $\mathbb{F}^{W}$-stopping time.
 For a fixed $z>0$ and $y\in(0,z]$, we consider $\alpha=\frac{y}{z}$ in the following.
@@
 $U(z,\beta)=\mathbb{E}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(\tilde{u}(Z^{D^{*}}_{t})+\ell Z^{D^{*}}_{t}\big)\,dt\bigg]\geq\tilde{V}(z,\beta),$
-where we used that $D^{*}\in\mathcal{D}$ since $(\xi_{t}^{*})_{t}$ is nondecreasing, càdlàg, and $\bar{\mathbb{F}}$-adapted.
+where we used that $D^{*}\in\mathcal{D}$ since $(\xi_{t}^{*})_{t}$ is nondecreasing, càdlàg, and $\mathbb{F}^{W}$-adapted.
 Step 3.
@@
 and for all $(x,\beta)\in\mathcal{O}$, there exists $\hat{z}:=\hat{z}(x,\beta)>0$ such that $\tilde{V}_{z}(\hat{z},\beta)=-x$. Furthermore, the optimal primal controls for (3.11) are given by
-$c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}\qquad\text{and}\qquad\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})-\frac{\rho\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
+$c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}\qquad\text{and}\qquad\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})+\frac{\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
 where $(D^{*}_{t})_{t}$ is the optimal singular control for problem (3.6); that is
@@
 $\hat{z}\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}\mathcal{H}_{t}D_{t}^{*}(c_{t}^{*}-\ell)\,dt\bigg]$ $=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}Z_{t}^{D^{*}}(c_{t}^{*}-\ell)\,dt\bigg]$
-$=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(u(c_{t}^{*})-\tilde{u}(Z_{t}^{D^{*}})\big)\,dt\bigg]-\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\ell Z_{t}^{D^{*}}\,dt\bigg],$
-and hence, by using (4.14), we obtain
+$=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(u(c_{t}^{*})-\tilde{u}(Z_{t}^{D^{*}})\big)\,dt\bigg]-\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\ell Z_{t}^{D^{*}}\,dt\bigg].$
+By Proposition 2.4-(2), there exists an investment strategy $\pi^{*}$ such that $(\pi^{*},c^{*})\in\mathcal{A}(x)$ thanks to $\chi(\hat{z})=x$. Hence, by using (4.14), we obtain
 $\tilde{V}(\hat{z},\beta)+\hat{z}x$ $=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t}^{*})\,dt\bigg]\leq\sup_{(\pi,c)\in\mathcal{A}(x)}\mathbb{E}_{x,\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t})\,dt\bigg]$
@@
 $V(x,\beta)=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}u(c_{t}^{*})\,dt\bigg],$
-so that $c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}$ is optimal.
-By Proposition 2.6, there exists an investment strategy $\pi^{*}$ such that $(\pi^{*},c^{*})\in\mathcal{A}(x)$.
-Moreover, by the strong Markov property and $x=-\tilde{V}_{z}(\hat{z},\beta)$, we have
+so that $c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}$ is optimal. Moreover, by the strong Markov property and $x=-\tilde{V}_{z}(\hat{z},\beta)$, we have
 $X_{t}^{*}=-\tilde{V}_{z}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}).$
@@
 (4.19) $dX_{t}^{*}=$ $-\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[(\delta-r)Z_{t}^{\hat{z},\beta,D^{*}}\,dt-\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\,dW_{t}\right]$
-$-\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta}\right]$
+$-\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[\kappa(\overline{\beta}-\beta_{t})\,dt-\sigma_{\beta}\,dW_{t}\right]$
 $-\frac{1}{2}\frac{\beta_{t}^{2}}{\sigma^{2}}(Z_{t}^{\hat{z},\beta,D^{*}})^{2}\tilde{V}_{zzz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\;dt$
-$-\frac{1}{2}\sigma_{\beta}^{2}\tilde{V}_{z\beta\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt+\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\sigma_{\beta}\rho\tilde{V}_{zz\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt,\quad t\geq 0.$
-Recalling that $dW^{\beta}_{t}=\rho dW_{t}+\sqrt{1-\rho^{2}}dW^{2,\perp}_{t}$ due to $Corr(W_{t},W^{\beta}_{t})=\rho t$, comparing the $dW_{t}$ terms in (2.1) and (4.19) finally gives
-$\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})-\frac{\rho\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
+$-\frac{1}{2}\sigma_{\beta}^{2}\tilde{V}_{z\beta\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt-\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\sigma_{\beta}\tilde{V}_{zz\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt,\quad t\geq 0.$
+Comparing the $dW_{t}$ terms in (2.3) and (4.19) finally gives
+$\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})+\frac{\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
 which completes the proof. ∎
@@
 In order to illustrate the singular control mechanism and how $(D^{*}_{t})_{t}$ ensures the nonnegativity of the wealth process, in Figure 1, we simulate the optimal primal state process $(X_{t}^{\pi^{*},c^{*}})_{t}$, the optimal dual state process $(Z^{\hat{z},D^{*}}_{t})_{t}$, and the optimal singular control $(D^{*}_{t})_{t}$.
-For this simulation, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\kappa=0.25$, $\bar{\beta}=0.05$, $\sigma_{\beta}=0.03$, $\sigma=0.18$, and $\rho=0.6$.
+For this simulation, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\kappa=0.25$, $\bar{\beta}=0.05$, $\sigma_{\beta}=0.03$, and $\sigma=0.18$.
 Figure 1(a) depicts a path of the controlled dual state process $(Z^{\hat{z},D^{*}}_{t})_{t}$ (black line) and the (moving stochastic) free boundary $(z^{*}(\beta_{t}))_{t}$ (red line).
@@
 In the stochastic case, we additionally set $\kappa=0.25$ and $\sigma_{\beta}=0.03$, while in the constant case we set $\kappa=0=\sigma_{\beta}$ and $\beta_{0}=\bar{\beta}$ to ensure that $(\beta_{t})_{t}\equiv\bar{\beta}$.
-We perform the comparison for three distinct correlation parameters: $\rho\in\{-0.6,0,0.6\}$.
-To isolate the impact of the stochastic factor $(\beta_{t})_{t}$, we fix the same path for the Brownian motion $(W_{t})_{t}$ driving the risky asset price in both cases.
-For the agent who assumes a stochastic $(\beta_{t})_{t}$, we simulate $10,000$ paths of the Brownian motion $(W^{\beta}_{t})_{t}$ driving the stochastic factor, and plot the optimal wealth as the average over these paths (blue line).
-The red line depicts the wealth of the agent who assumes a constant $(\beta_{t})_{t}$.
+In both cases, we simulate $10,000$ paths of the Brownian motion $(W_{t})_{t}$, and plot the optimal wealth as the average over these paths.
+The blue line depicts the wealth of the agent who assumes a stochastic $(\beta_{t})_{t}$, while the red line represents the wealth of the agent who assumes a constant $(\beta_{t})_{t}$.
 In the following, we denote by Stochastic Agent the agent with stochastic $(\beta_{t})_{t}$, and by Constant Agent the agent with constant $(\beta_{t})_{t}$.
-In all scenarios, the Stochastic Agent systematically accumulates greater long-term wealth.
-This outcome is driven by a universal mechanism: The Stochastic Agent’s strategic asset allocation, namely increasing exposure to the risky asset when expected returns are high (that is, when $\beta_{t}$ is high), generates substantial excess returns.
-As a result, the Stochastic Agent’s wealth exceeds that of the Constant Agent in the long-run.
+We observe that the Stochastic Agent systematically accumulates greater wealth.
+This outcome is driven by a robust economic mechanism: The Stochastic Agent’s strategic asset allocation, namely increasing exposure to the risky asset when expected returns are high (that is, when $\beta_{t}$ is high), generates substantial excess returns.
+As a result, the Stochastic Agent’s wealth exceeds that of the Constant Agent.
 In contrast, the Constant Agent lacks this flexibility and is confined to a rigid strategy that systematically fails to exploit time-varying risk premia, leading to inferior wealth accumulation.
-(a) Case $\rho=-0.6$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
-(b) Case $\rho=0$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
-(c) Case $\rho=0.6$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
-Figure 2.
-Comparison of optimal wealth trajectories between the Stochastic Agent and the Constant Agent ($\beta_{t}\equiv\bar{\beta}$) for different values of $\rho$.
-The Brownian path of $(W_{t})_{t}$ is fixed across both cases, while the Stochastic Agent’s wealth is averaged over 10,000 Brownian paths of $(W^{\beta}_{t})_{t}$.
-While this long-run dominance is consistent across all regimes, short-run dynamics differ.
-For nonnegative correlations ($\rho=0$ and $\rho=0.6$), the advantage materializes immediately, and the Stochastic Agent maintains wealth consistently above that of the Constant Agent throughout the simulation.
-In contrast, the regime with $\rho=-0.6$ exhibits a distinct crossing pattern.
-In this case, the negative correlation provides a natural hedge, as market downturns signal improved future investment opportunities.
-This added safety reduces the agent’s precautionary savings motive and induces a positive income effect, effectively making the agent feel wealthier and optimally consume more in the present.
-The resulting higher initial consumption causes the Stochastic Agent to temporarily lag behind the Constant Agent.
-However, as the benefits of flexible investment decisions accumulate, the wealth paths eventually cross, confirming that the long-run advantage of adapting to $(\beta_{t})_{t}$ ultimately dominates.
+Figure 2. Comparison of the average optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
+At the same time, the initial advantage builds up only gradually, so that the two wealth paths remain relatively close at the beginning of the horizon.
+Economically, under perfect negative correlation, adverse contemporaneous shocks to the risky asset are accompanied by improvements in future investment opportunities.
+This hedging effect dampens the short-run urgency of aggressive portfolio adjustment and smooths the initial wealth dynamics.
+Hence, although the Stochastic Agent is still able to exploit the time-varying investment opportunity set and eventually achieves substantially higher wealth, the momentum of this outperformance is weaker at the start.
+Over time, however, the cumulative benefits of adapting consumption and portfolio decisions to the stochastic factor $(\beta_{t})_{t}$ become increasingly important, leading to a persistent and widening wealth advantage over the Constant Agent.
 Next, we illustrate the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ given by Proposition 4.4 for different parameter choices.
@@
 Thus, the agent explicitly accounts for the stochastic evolution of $(\beta_{t})_{t}$ throughout the optimization, and the plots represent a cross-section of this dynamic strategy.
-In Figure 3, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of labor income $\ell\in\{0.2,0.6,1\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
+In Figure 3, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of labor income $\ell\in\{0.2,0.6,1\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
 We observe that both consumption and risky investment are increasing in wealth $X_{t}$ and labor income $\ell$.
@@
 Figure 3. Optimal policies for different values of labor income $\ell$.
-In Figure 4, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of risk aversion $\gamma\in\{1.2,1.5,2\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
+In Figure 4, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of risk aversion $\gamma\in\{1.2,1.5,2\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
 As before, we observe that both optimal consumption $(c^{*}_{t})_{t}$ and risky investment $(\pi^{*}_{t})_{t}$ are strictly increasing in wealth $X_{t}$ since the agent has more money to allocate.
@@
 This inverse relationship is consistent with Merton’s classic results, reflecting that more risk-averse agents fear market volatility more and consequently reduce their exposure to the risky asset.
-Conversely, the consumption policy $(c^{*}_{t})_{t}$ roughly shifts upward as $\gamma$ increases.
+Conversely, the consumption policy $(c^{*}_{t})_{t}$ shifts upward as $\gamma$ increases.
 This behavior is driven by the Intertemporal Elasticity of Substitution (IES).
@@
 Figure 4. Optimal policies for different values of risk aversion $\gamma$.
-In Figure 5, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of the correlation parameter $\rho\in\{-0.6,0,0.6\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
-We observe that the consumption policy $(c_{t}^{*})_{t}$ is increasing in $\rho$.
-In particular, the agent with positive correlation ($\rho=0.6$) consumes the most, while the agent with negative correlation ($\rho=-0.6$) consumes the least.
-This pattern reflects the precautionary savings motive.
-When $\rho<0$, the risky asset acts as an intertemporal hedge, encouraging the agent to take a highly leveraged position.
-This leverage increases portfolio volatility, and to guard against risk and avoid breaching the borrowing constraint, the agent reduces current consumption.
-By contrast, when $\rho>0$, the asset offers weaker hedging benefits, leading to lower risky exposure, reduced volatility, and a weaker precautionary savings motive, allowing the agent to consume more.
-The investment policy $(\pi_{t}^{*})_{t}$ is non-monotonic with respect to $\rho$.
-It rises with $\rho$ up to a wealth level of roughly $X_{t}\approx 6.5$, and then declines for higher wealth.
-In the high-wealth region, the policy follows standard Merton intuition: the negatively correlated agent ($\rho=-0.6$) invests more aggressively to exploit the asset’s hedging properties, while investment declines with $\rho$.
-At lower wealth levels ($X_{t}<6.5$), however, the pattern flips.
-The hedging agent’s highly leveraged strategy becomes too risky because of the threat of immediate bankruptcy, forcing sharp deleveraging.
-Meanwhile, the positively correlated agent ($\rho=0.6$), who naturally maintains a more conservative position, is less constrained and thus holds a higher allocation in the low-wealth region.
-(a) Optimal investment policy $\pi^{*}_{t}$ for different values of correlation $\rho$.
-(b) Optimal consumption policy $c^{*}_{t}$ for different values of correlation $\rho$.
-Figure 5. Optimal policies for different values of correlation $\rho$.
-Continuing with our cross-sectional approach, Figure 6 compares the optimal control policies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of expected excess returns $\beta_{t}\in\{0.02,0.05,0.12\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
+Continuing with our cross-sectional approach, Figure 5 compares the optimal control policies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of expected excess returns $\beta_{t}\in\{0.02,0.05,0.12\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
 We observe that the investment policy $(\pi_{t}^{*})_{t}$ is strictly increasing in the expected excess return $\beta_{t}$.
@@
 The consumption policy $(c_{t}^{*})_{t}$ displays a non-monotonic relationship with $\beta_{t}$.
-At low wealth levels ($X_{t}<3$), consumption decreases as $\beta_{t}$ increases.
+At low wealth levels ($X_{t}<3.5$), consumption decreases as $\beta_{t}$ increases.
 In this region, the agent seeks to accumulate wealth more rapidly.
@@
 6. Conclusions
-In this paper, we have studied the infinite-horizon consumption and portfolio problem of an investor subject to a no-borrowing constraint and labor income, in a market where expected excess returns follow a mean-reverting Ornstein-Uhlenbeck process.
+In this paper, we have studied the infinite-horizon consumption and portfolio problem of an investor subject to a no-borrowing constraint and labor income in the complete-market case of the Kim–Omberg model, where expected excess returns follow a mean-reverting Ornstein–Uhlenbeck process and are perfectly negatively correlated with the risky asset return shocks.
 Using a Lagrange duality approach, we have formulated the dual problem as a two-dimensional singular control problem involving the marginal value of wealth and the stochastic factor.
@@
 Finally, we have retrieved the solutions to the primal optimization problem via duality and provided numerical illustrations.
+An interesting direction is the extension to the incomplete-market case.
+In that setting, the duality approach developed here is no longer directly applicable, as our proof of strong duality would break down, and a different mathematical approach, likely involving the associated HJB equation and verification arguments, would be needed.
+Such a challenging extension is left for future research.
 Acknowledgements. Financial support by the German Research Foundation (DFG) [RTG
@@
 […]
-[…]
+Under $\mathbb$$\bar${P}}, we then have for any $\mathbb{F}^{W}$-stopping time $\tau$:
 […]
@@
 $d\hat{Z}_{t}$ $=-\frac{\hat{\beta}_{t}}{\sigma}\hat{Z}_{t}\,dW^{\mathbb{Q}}_{t}+\hat{Z}_{t}\bigg(\delta-r+\frac{\hat{\beta}_{t}^{2}}{\sigma^{2}}\bigg)\,dt,\quad t>0,\quad\hat{Z}_{0}=z>0,$
-$d\hat{\beta}_{t}$ $=\sigma_{\beta}\,dW^{\beta,\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})-\rho\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
-We set $\hat{\mathbb{F}}:=(\hat{\mathcal{F}}_{t})_{t}$, where $\hat{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W,\mathbb{Q}}\vee\mathcal{F}_{t}^{W^{\beta},\mathbb{Q}}$, $t\geq 0$.
-We then define the following value functions (where the infimum is taken over all $\hat{\mathbb{F}}-$stopping times)
+$d\hat{\beta}_{t}$ $=-\sigma_{\beta}\,dW^{\mathbb{Q}}_{t}+\bigg(\kappa(\overline{\beta}-\hat{\beta}_{t})+\,\frac{\hat{\beta}_{t}}{\sigma}\sigma_{\beta}\bigg)\,dt,\quad t>0,\quad\hat{\beta}_{0}=\beta\in\mathbb{R}.$
+We then define the following value functions (where the infimum is taken over all $\mathbb{F}^{W,\mathbb{Q}}$-stopping times)
 $v^{\mathbb{Q}}(z,\beta;T):=\inf_{\tau}\mathbb{E}^{\mathbb{Q}}_{z,\beta}\left[\int_{0}^{\tau\wedge T}e^{-rt}(\tilde{u}^{\prime}(\hat{Z}_{t})+\ell)\,dt\right],$
@@
 By Fubini-Tonelli Theorem, we may write
-$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]=\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]\,dt,$
-and we recall that
-$W_{t}^{\mathbb{Q}}=\rho\,W_{t}^{\beta,\mathbb{Q}}+\sqrt{1-\rho^{2}}\,\widehat{W}^{1,\perp}_{t},\quad t\geq 0,$
-due to the correlation between $W^{\mathbb{Q}}$ and $W^{\beta,\mathbb{Q}}$, where $\widehat{W}^{1,\perp}$ is a Brownian motion independent of $W^{\beta,\mathbb{Q}}$.
+$\mathbb{E}^{\mathbb{Q}}_{z,\beta}\bigg[\int_{0}^{\infty}e^{-rt}\hat{Z}^{-\frac{1}{\gamma}}_{t}dt\bigg]=\int_{0}^{\infty}e^{-rt}\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]\,dt.$
 By using the explicit representation of the strong solution to (3.16), we find
 $\mathbb{E}^{\mathbb{Q}}_{z,\beta}[\hat{Z}_{t}^{-\frac{1}{\gamma}}]=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds+\frac{1}{\gamma}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\bigg)\bigg]$
-[…]
-$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\mathbb{E}^{\mathbb{Q}}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds$
-$\quad+\frac{1}{\gamma}\rho\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW^{\beta,\mathbb{Q}}_{s}+\frac{1}{\gamma}\sqrt{1-\rho^{2}}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}d\widehat{W}^{1,\perp}_{s}\bigg)\bigg|\mathcal{F}_{t}^{W^{\beta},\mathbb{Q}}\bigg]\bigg]$
-$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}\int_{0}^{t}(\delta-r+\tfrac{1}{2}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}})ds+\frac{1}{\gamma}\rho\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW^{\beta,\mathbb{Q}}_{s}+\tfrac{1}{2}\frac{1}{\gamma^{2}}(1-\rho^{2})\int_{0}^{t}\tfrac{\hat{\beta}_{s}^{2}}{\sigma^{2}}ds\bigg)\bigg]$
-[…]
+$=z^{-\frac{1}{\gamma}}\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{\gamma}(\delta-r)t-\tfrac{1}{2\sigma^{2}\gamma}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\frac{1}{\gamma}\int_{0}^{t}\tfrac{\hat{\beta}_{s}}{\sigma}dW_{s}^{\mathbb{Q}}\bigg)\bigg]$
+$=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)\cdot$
+$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(-\frac{1}{2\sigma^{2}\gamma}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg]$
 $=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)$
-$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\bigg(\tfrac{\frac{1}{\gamma^{2}}(1-\rho^{2})-\frac{1}{\gamma}}{2\sigma^{2}}\bigg)\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds+\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg]$
-$=z^{-\frac{1}{\gamma}}\exp\big(-\frac{1}{\gamma}(\delta-r)t\big)$
-$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\tfrac{1-\gamma}{2\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\exp\bigg(\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg].$
+$\quad\cdot\mathbb{E}^{\mathbb{Q}}_{1,\beta}\bigg[\exp\bigg(\tfrac{1-\gamma}{2\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\exp\bigg(\tfrac{1}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)\bigg].$
 The process
-$N_{t}:=\exp\bigg(\tfrac{\rho}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\beta,\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{\rho^{2}}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)$
+$N_{t}:=\exp\bigg(\tfrac{1}{\sigma\gamma}\int_{0}^{t}\hat{\beta}_{s}dW^{\mathbb{Q}}_{s}-\tfrac{1}{2}\tfrac{1}{\sigma^{2}\gamma^{2}}\int_{0}^{t}\hat{\beta}_{s}^{2}ds\bigg)$
 defines a martingale under $\mathbb{Q}$ by Assumption 3.3. Hence, Girsanov’s Theorem allows us to define a new probability measure by
-$\frac{d\mathbb{Q^{\prime}}}{d\mathbb{Q}}\bigg|_{\hat{\mathcal{F}}_{t}}=N_{t},$
+$\frac{d\mathbb{Q^{\prime}}}{d\mathbb{Q}}\bigg|_{\mathcal{F}^{W,\mathbb{Q}}_{t}}=N_{t},$
 and to obtain, as $\gamma>1$,
@@
 (A.5) $=|z^{-\frac{1}{\gamma}}-(z+\varepsilon)^{-\frac{1}{\gamma}}|\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big]$
-$=\frac{1}{\gamma}\xi^{-\frac{1}{\gamma}-1}\,\varepsilon\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big],$
+$=\frac{1}{\gamma}\zeta^{-\frac{1}{\gamma}-1}\,\varepsilon\,\mathbb{E}^{\mathbb{Q}}\Big[\int_{0}^{\infty}e^{-rt}(\hat{Z}^{1,\beta}_{t})^{-\frac{1}{\gamma}}dt\Big],$
 for some $\zeta\in[z,z+\varepsilon]$. Upon setting
@@
 so that the unique strong solution to (3.17) can be written as
-(A.8) $\hat{\beta}^{\beta}_{t}=\beta e^{-at}+b(1-e^{-at})+\sigma_{\beta}\int_{0}^{t}e^{-a(t-s)}dW_{s}^{\beta,\mathbb{Q}},$
+(A.8) $\hat{\beta}^{\beta}_{t}=\beta e^{-at}+b(1-e^{-at})-\sigma_{\beta}\int_{0}^{t}e^{-a(t-s)}dW_{s}^{\mathbb{Q}},$
 and it readily follows that $\frac{\partial\hat{\beta}^{\beta}_{t}}{\partial\beta}=e^{-at}$. Since
@@
 Let
-$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\}$
+$1<p<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}{\sigma_{\beta}}\bigg\}$
 and $q>1$ with $\frac{1}{p}+\frac{1}{q}=1$. Hölder’s inequality and (3.18) then yield
@@
 (A.13) $G^{\varepsilon}_{t}$ $:=\frac{1}{\sigma^{2}}\int_{0}^{t}e^{-as}(\beta_{\varepsilon}e^{-as}+b(1-e^{-as}))ds,$
-(A.14) $B_{t}$ $:=\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}e^{-as}\Big(\int_{0}^{s}e^{-a(s-u)}dW_{u}^{\beta,\mathbb{Q}}\Big)ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}.$
+(A.14) $B_{t}$ $:=-\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}e^{-as}\Big(\int_{0}^{s}e^{-a(s-u)}dW_{u}^{\mathbb{Q}}\Big)ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}.$
 A direct calculation yields
@@
 with $C_{G}(\beta)$ independent of $\varepsilon>0$. For $B_{t}$, an application of Stochastic Fubini Theorem (cf., e.g., Theorem $65$ in section IV of [39]) leads to
-$B_{t}$ $=\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}\left(\int_{u}^{t}e^{-as}e^{-a(s-u)}ds\right)dW_{u}^{\beta,\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
-(A.16) $=\frac{\sigma_{\beta}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\beta,\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
-$=\frac{\sigma_{\beta}\rho}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\mathbb{Q}}+\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{2,\perp}$
-$\quad-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
-$=\int_{0}^{t}\underbrace{\left(\frac{\sigma_{\beta}\rho}{2a\sigma^{2}}(e^{-au}-e^{-2at+au})-\frac{1}{\sigma}e^{-au}\right)}_{=:f(u,t)}dW_{u}^{\mathbb{Q}}+\int_{0}^{t}\underbrace{\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}\left(e^{-au}-e^{-2at+au}\right)}_{=:g(u,t)}d\widehat{W}_{u}^{2,\perp},$
-where we used $W_{u}^{\beta,\mathbb{Q}}=\rho W_{u}^{\mathbb{Q}}+\sqrt{1-\rho^{2}}\widehat{W}_{u}^{2,\perp}$ due to $Corr(W_{u}^{\mathbb{Q}},W_{u}^{\beta,\mathbb{Q}})=\rho u$.
-The deterministic functions $f(u,t)$ and $g(u,t)$ in (A.16) can be bounded as follows
-$|f(u,t)|\leq e^{-au}\left(\frac{\sigma_{\beta}|\rho|}{2a\sigma^{2}}+\frac{1}{\sigma}\right)=:C_{f}e^{-au},\quad|g(u,t)|\leq\frac{\sigma_{\beta}\sqrt{1-\rho^{2}}}{2a\sigma^{2}}e^{-au}=:C_{g}e^{-au},$
-and since $W^{\mathbb{Q}}$ is independent of $\widehat{W}^{2,\perp}$, we obtain
-(A.17) $\mathrm{Var}(B_{t})$ $=\int_{0}^{t}f(u,t)^{2}du+\int_{0}^{t}g(u,t)^{2}du\leq\frac{C_{f}^{2}+C_{g}^{2}}{2a}=:K.$
+$B_{t}$ $=-\frac{\sigma_{\beta}}{\sigma^{2}}\int_{0}^{t}\left(\int_{u}^{t}e^{-as}e^{-a(s-u)}ds\right)dW_{u}^{\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
+(A.16) $=-\frac{\sigma_{\beta}}{2a\sigma^{2}}\int_{0}^{t}\left(e^{-au}-e^{-2at+au}\right)dW_{u}^{\mathbb{Q}}-\frac{1}{\sigma}\int_{0}^{t}e^{-au}dW_{u}^{\mathbb{Q}}$
+$=\int_{0}^{t}\underbrace{\left(-\frac{\sigma_{\beta}}{2a\sigma^{2}}(e^{-au}-e^{-2at+au})-\frac{1}{\sigma}e^{-au}\right)}_{=:f(u,t)}dW_{u}^{\mathbb{Q}}.$
+The deterministic function $f(u,t)$ in (A.16) can be bounded as follows
+$|f(u,t)|\leq e^{-au}\left(\frac{\sigma_{\beta}}{2a\sigma^{2}}+\frac{1}{\sigma}\right)=:C_{f}e^{-au},$
+and hence, we obtain
+(A.17) $\mathrm{Var}(B_{t})$ $=\int_{0}^{t}f(u,t)^{2}du\leq\frac{C_{f}^{2}}{2a}=:K.$
 Upon noticing that $B_{t}$ is Gaussian with mean 0 and variance $\mathrm{Var}(B_{t})\leq K$, we may write $B_{t}=\sqrt{\mathrm{Var}(B_{t})}Y$, where $Y\sim\mathcal{N}(0,1)$.
@@
 Next, we show that the family
-(A.22) $\bigg\{\left|(\hat{Z}_{t}^{\beta_{\varepsilon}})^{-1/\gamma}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|\mathbf{1}_{\{t\leq\tau^{*}\}}\bigg\}_{\varepsilon\in(0,\varepsilon_{0})}$
+(A.22) $\bigg\{\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-1/\gamma}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|\mathbf{1}_{\{t\leq\tau^{*}\}}\bigg\}_{\varepsilon\in(0,\varepsilon_{0})}$
 is uniformly integrable. We let
-$1<\tilde{p}<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa+\frac{\rho\sigma_{\beta}}{\sigma}\bigg)}{|\rho|\sigma_{\beta}}\bigg\},$
+$1<\tilde{p}<\min\bigg\{\gamma,\frac{\gamma\sigma\bigg(\kappa-\frac{\sigma_{\beta}}{\sigma}\bigg)}{\sigma_{\beta}}\bigg\},$
 […]
@@
 Now, recalling (A.12) and using the arguments used for the derivation of (A.19), we find
-(A.24) $\|A_{t}\|_{L^{mq}}\leq M_{mq}(\beta),$
+(A.24) $\|A^{\varepsilon}_{t}\|_{L^{mq}}\leq M_{mq}(\beta),$
 since $mq>1$, which is uniform in $\varepsilon$. Hence, combining (A.5), (A.24) and (3.18) gives
-$\sup_{\varepsilon\leq\varepsilon_{0}}\mathbb{E}^{\mathbb{Q}}\!\left[\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|^{m}\mathbf{1}_{\{t\leq\tau^{*}\}}\right]\leq z^{-\frac{s}{\gamma}}\,e^{-\frac{s}{\gamma}(\delta-r)t}M_{mq}(\beta)^{m},$
+$\sup_{\varepsilon\leq\varepsilon_{0}}\mathbb{E}^{\mathbb{Q}}\!\left[\left|(\hat{Z}_{t}^{z,\beta_{\varepsilon}})^{-\frac{1}{\gamma}}\left(\int_{0}^{t}e^{-as}\frac{\hat{\beta}_{s}^{\beta_{\varepsilon}}}{\sigma^{2}}ds-\frac{1}{\sigma}\int_{0}^{t}e^{-as}dW_{s}^{\mathbb{Q}}\right)\right|^{m}\mathbf{1}_{\{t\leq\tau^{*}\}}\right]\leq z^{-\frac{m}{\gamma}}\,e^{-\frac{m}{\gamma}(\delta-r)t}M_{mq}(\beta)^{m},$
 and thus the uniform integrability of the family (A.22) follows.